### v1

# Interpolation theorems for conjugations and applications

Abstract.
Let $H$ be a separable complex Hilbert space. A conjugate-linear map $C:H\to H$ is a conjugation if it is an involutive isometry. In this paper, we consider the following interpolation problems: Let $\{x_{i}\}_{i\in I}$ and $\{y_{i}\}_{i\in I}$ be two orthogonal sets of vectors in $H$, and let $N$ and $\{N_{k}\}_{k\in K}$ be normal operators such that the $N_{k}$’s mutually commute. Then, under which conditions does there exist a conjugation $C$ on $H$ such that
  1. (a)
    $Cx_{i}=y_{i}$ and $CN_{k}C=N_{k}^{*}$ for all $i\in I$ and $k\in K$; or
  2. (b)
    $Cx_{i}=y_{i}$, for every $i\in I$, and $CNC=-N^{*}$.
We provide complete answers to problems (a) and (b). As a consequence of our results, we give necessary and sufficient conditions for the existence of solutions of some equations in $L^{\infty}(\mu)$.
Key words and phrases:
Conjugations, interpolation, complex-symmetric operators, skew-symmetric operators
2020 Mathematics Subject Classification:
47B15, 47B38, 47A15.
1. Introduction
Throughout this paper, $H$ and $K$ are separable complex Hilbert spaces. We use the notation .,$\langle$.,$\rangle$ . , . ⟩ for the inner product of both $H$ and $K$. The algebra of all bounded linear operators on $H$ is denoted by ${\mathcal{B}}(H)$, and the identity operator on $H$ is denoted by $I_{H}$; if there is no ambiguity, we simply write $I$. The range and the null space of an operator $T$ are denoted by ${\rm ran}(T)$ and ${\rm ker}(T)$, respectively. All direct sums in this paper are orthogonal.
A conjugation $C$ on $H$ is a conjugate-linear map $C:H\to H$ that satisfies
  1. (i)
    $C^{2}=I$;
  2. (ii)
    $\langle Ch,Ck\rangle=\langle k,h\rangle$ for all $h,k\in H$.
The simplest example of a conjugation on $H$ is the map given by $C\left(\sum_{i}\lambda_{i}e_{i}\right)=\overline{\lambda_{i}}e_{i}$, where $\{e_{i}\}_{i}$ is an orthonormal basis of $H$. In fact, according to [13, Lemma 1], any conjugation can be represented in this manner with respect to some orthonormal basis.
Conjugations have been the subject of intensive study in recent years [5, 6, 11, 12, 20, 21], with origins in physics. The composition $\mathcal{P}\mathcal{T}$, of the parity operator $[\mathcal{P}f](x)=f(-x)$ and the time-reversal operator $[\mathcal{T}f](x)=\overline{f(x)}$ defines a conjugation on $L^{2}(\mathbb{R}^{n})$ which holds significant importance in the theory of $\mathcal{P}\mathcal{T}$-symmetric quantum theory (see, for instance, [3, 4]). For more historical comments about conjugations, we refer the interested reader to [5, 26] and references therein. Recent investigations into conjugations have largely stemmed from their connections to complex-symmetric and skew-symmetric operators.
An operator $T\in{\mathcal{B}}(H)$ is called complex-symmetric (resp. skew-symmetric) if there exists a conjugation $C$ on $H$ such that $CTC=T^{*}$ (resp. $CTC=-T^{*}$); in this case, we say more precisely that $T$ is $C$-symmetric (resp. $C$-skew symmetric). An equivalent definition is that $T$ has a symmetric (resp. skew-symmetric) matrix $M$ with respect to some orthonormal basis; i.e., $M^{\rm tr}=M$ (resp. $M^{\rm tr}=-M$) where ${\rm tr}$ stands for the transpose ([14, 17]). These classes of linear operators have received significant attention in the last two decades. Applications of complex symmetric operators and skew-symmetric matrices are found in various areas (see, for instance, [13, 14, 15, 24] and [7, 22, 23]).
In [18, 19, 26, 29], the authors investigated various interpolation problems for conjugations. In [18], Liu et al. established that for two orthogonal projections $P$ and $Q$ in ${\mathcal{B}}(H)$, a conjugation $C$ on $H$ exists with $CPC=Q$ if and only if ${\rm ran}(P)$ and ${\rm ran}(Q)$ are in symmetric position; i.e.,
$\dim\left({\rm ran}(P)^{\perp}\cap{\rm ran}(Q)\right)=\dim\left({\rm ran}(P)\cap{\rm ran}(Q)^{\perp}\right).$
In [26], Wang et al. proved that for two commuting self-adjoint operators $P,Q\in{\mathcal{B}}(H)$, there exists a conjugation $C$ on $H$ satisfying $CPC=Q$ if and only if there exists an isometric isomorphism $U$ between Hilbert spaces such that
$UPU^{-1}=B_{1}\oplus B_{2}\oplus A\quad\mbox{and}\quad UQU^{-1}=B_{2}\oplus B_{1}\oplus A$
where $B_{1}$, $B_{2}$ and $A$ are self-adjoint operators such that $B_{1}$ and $B_{2}$ have the same underlying Hilbert space and commute.
It is well-known that for every normal operator $N$, there exists a conjugation $C$ such that $CNC=N^{*}$ ([13]). However, this is not always the case if it is also required that $Cx_{i}=y_{i}$ for some families of vectors $\{x_{i}\}_{i\in I}$ and $\{y_{i}\}_{i\in I}$. It is the purpose of this paper to investigate such a problem. The paper is structured as follows.
In Section 2, we present our primary results regarding the interpolation problem: when does there exist a conjugation $C$ on $H$ satisfying $Cx_{i}=y_{i}$ and $CN_{k}C=N_{k}^{*}$ for all $i\in I$ and $k\in K$. Here, $\{x_{i}\}_{i\in I}$ and $\{y_{i}\}_{i\in I}$ are orthogonal subsets of $H$, and $\{N_{k}\}_{k\in K}$ is a collection of mutually commuting normal operators on $H$. We establish geometric conditions to answer this question, along with several equivalent conditions. As a consequence of these results, we extend a recent result by Mashreghi et al. [20] concerning hyperinvariant subspaces of unitary operators (see Proposition 2.5).
In Section 3, we investigate a “skew-symmetric” version of the earlier interpolation problem. We aim to identify the conditions under which there exists a conjugation $C$ on $H$ that can fulfill the conditions $Cx_{i}=y_{i}$ and $CNC=-N^{*}$ for all $i\in I$, with $N$ being a normal operator. The simplest answer is obtained when $I$ is finite.
In the last section, we apply our main results to give necessary and sufficient conditions for the existence of solutions of certain equations in $L^{\infty}(\mu)$ with coefficients in $L^{2}(\mu)$.
2. Interpolation theorems for conjugations : C-symmetric versions
Recall that a linear (or conjugate-linear) operator $V$ on $H$ is called partial isometry if the restriction of $V$ to ${\rm ker}(V)^{\perp}$ is an isometry.
2.1. Main results
Theorem 2.1.
Let $\mathcal{M}$ be an abelian von Neumann algebra of operators in ${\mathcal{B}}(H)$, and let $\{x_{i}\}_{i\in I}$ and $\{y_{i}\}_{i\in I}$ be two orthogonal sets of non-zero vectors in $H$. The following statements are equivalents:
  1. (i)
    There exists a conjugation $C$ on $H$ such that
    $CTC=T^{*}\;\;\mbox{and}\;\;Cx_{i}=y_{i}\;\;\;\mbox{for all $T\in\mathcal{M}$ and $i\in I$}.$
  2. (ii)
    There exists a conjugate-linear partial isometry $V$ on $H$ such that
    $VT=T^{*}V,\;\;x_{i}\in{\rm ker}(V)^{\perp},\;\;Vx_{i}=y_{i}\;\;\mbox{and}\;\;Vy_{i}=x_{i}\;\;\;\mbox{for all $T\in\mathcal{M}$ and $i\in I$}.$
  3. (iii)
    For every orthogonal projection $P\in\mathcal{M}$ and for all $i,j\in I$, we have
    $\langle Px_{i},x_{j}\rangle=\langle Py_{j},y_{i}\rangle\;\;\mbox{and}\;\;\langle Px_{i},y_{j}\rangle=\langle Px_{j},y_{i}\rangle.$
Remark 2.2.
Noting that $C(\lambda I)C=(\lambda I)^{*}$ for all conjugations $C$ and $\lambda\in\mathbb{C}$, one can see that when =$\mathcal{M}=\{\lambda I:\lambda\in\mathbb{C}\}$, we recapture [29, Theorem 2.1]. This theorem states that for two orthonormal sets $\{x_{i}\}_{i\in I}$ and $\{y_{i}\}_{i\in I}$ in $H$, there exists a conjugation $C$ on $H$ satisfying $Cx_{i}=y_{i}$ for every $i\in I$ if and only if $\langle x_{i},y_{j}\rangle=\langle x_{j},y_{i}\rangle$ for all $i,j\in I$. It should be noted that their result is important in proving our main results.
For a normal operator $N$, we denote by $E_{N}$ the associated spectral measure defined on Borel subsets of $\mathbb{C}$.
If $\mathcal{M}=W^{*}(N_{k}:k\in K)$, meaning that $\mathcal{M}$ is the smallest von Neumann algebra containing a family of mutually commuting normal operators $\{N_{k}\}_{k\in K}$, then the conditions outlined in Theorem 2.1 can be refined as follows.
Theorem 2.3.
Let $\{N_{k}\}_{k\in K}\subset{\mathcal{B}}(H)$ be a family of mutually commuting normal operators, and let $\{x_{i}\}_{i\in I}$ and $\{y_{i}\}_{i\in I}$ be two orthogonal sets of non-zero vectors in $H$. The following statements are equivalent:
  1. (i)
    There exists a conjugation $C$ on $H$ such that $CTC=T^{*}$ and $Cx_{i}=y_{i}$ for all $T\in W^{*}(N_{k}:k\in K)$ and $i\in I$.
  2. (ii)
    There exists a conjugation $C$ on $H$ such that $CN_{k}C=N_{k}^{*}$ and $Cx_{i}=y_{i}$ for all $k\in K$ and $i\in I$.
  3. (iii)
    There exists a conjugate-linear partial isometry $V$ on $H$ such that $VN_{k}=N_{k}^{*}V$, $x_{i}\in{\rm ker}(V)^{\perp}$, $Vx_{i}=y_{i}$ and $Vy_{i}=x_{i}$ for all $k\in K$ and $i\in I$.
  4. (iv)
    For all finite subsets $K_{0}\subseteq K$, Borel subsets $\{\Delta_{k}\}_{k\in K_{0}}$ of $\mathbb{C}$, and $i,j\in I$ we have
    $\left\langle\left(\prod_{k\in K_{0}}E_{N_{k}}(\Delta_{k})\right)x_{i},x_{j}\right\rangle=\left\langle\left(\prod_{k\in K_{0}}E_{N_{k}}(\Delta_{k})\right)y_{j},y_{i}\right\rangle$
    and
    $\left\langle\left(\prod_{k\in K_{0}}E_{N_{k}}(\Delta_{k})\right)x_{i},y_{j}\right\rangle=\left\langle\left(\prod_{k\in K_{0}}E_{N_{k}}(\Delta_{k})\right)x_{j},y_{i}\right\rangle.$
Given non-zero vectors $x,y\in H$, we denote by $x\otimes y$ the rank-one operator defined by $(x\otimes y)h=\langle h,y\rangle x$ for every $h\in H$.
If $I$ is of cardinal one, then Theorem 2.3 reduces to:
Theorem 2.4.
Let $\{N_{k}\}_{k\in K}\subset{\mathcal{B}}(H)$ be a family of mutually commuting normal operators, and let $x,y\in H$ be non-zero vectors. The following statements are equivalent:
  1. (i)
    There exists a conjugation $C$ on $H$ satisfying $CN_{k}C=N_{k}^{*}$ and $Cx=y$ for every $k\in K$.
  2. (ii)
    There exists a unitary operator $U\in{\mathcal{B}}(H)$ satisfying $UN_{k}=N_{k}U$ and $Ux=y$ for every $k\in K$.
  3. (iii)
    There exists a partial isometry $V\in{\mathcal{B}}(H)$ satisfying $x\in{\rm ker}(V)^{\perp}$, $VN_{k}=N_{k}V$ and $Vx=y$ for every $k\in K$.
  4. (iv)
    For all finite subsets $K_{0}\subseteq K$ and Borel subsets $\{\Delta_{k}\}_{k\in K_{0}}$ of $\mathbb{C}$, we have
    $\left\|\left(\prod_{k\in K_{0}}E_{N_{k}}(\Delta_{k})\right)x\right\|=\left\|\left(\prod_{k\in K_{0}}E_{N_{k}}(\Delta_{k})\right)y\right\|.$
  5. (v)
    $\|x\|=\|y\|$ and there exists a conjugation $C$ on $H$ such that $N_{k}+\lambda x\otimes y$ is $C$-symmetric for all $k\in K$ and $\lambda\in\mathbb{C}$.
The proofs of Theorems 2.1, 2.3, and 2.4 will be given later in this section, following the necessary preliminary preparations.
A slight reformulation of [1, Theorem 2.1] shows that if $N\in{\mathcal{B}}(H)$ is a normal operator, $U\in{\mathcal{B}}(H)$ is a unitary operator commuting with $N$, and $x\in H$, then for every $\lambda\in\mathbb{C}$, there exists a conjugation $C$ on $H$ such that $N+\lambda x\otimes Ux$ is $C$-symmetric. It can be seen that the implication (ii)$\implies$(v) in Theorem 2.4 generalizes this result, as the conjugation $C$ is independent of $\lambda$.
A subspace $M$ of $H$ is said to be hyperinvariant for $T\in{\mathcal{B}}(H)$ if $M$ is invariant under every operator that commutes with $T$. Recently, in [20], Mashreghi et all. proved that $M$ is hyperinvariant for a unitary operator $U\in{\mathcal{B}}(H)$ if and only if $CM\subseteq M$ for every conjugation $C$ on $H$ satisfying $CUC=U^{*}$. As a consequence of Theorem 2.4, this result is extended to all normal operators.
Proposition 2.5.
Let $N\in{\mathcal{B}}(H)$ be normal and $M$ a subspace of $H$. Then, $M$ is hyperinvariant for $N$ if and only if $CM\subseteq M$ for every conjugation $C$ on $H$ satisfying $CNC=N^{*}$.
The proof of the necessity is similar to the one in [20]. We include it here for the sake of completeness.
Proof of Proposition 2.5.
Assume that $M$ is hyperinvariant for $N$ and denote by $P\in{\mathcal{B}}(H)$ the orthogonal projection onto $M$. Let $C$ be a conjugation on $N$ satisfying $CNC=N^{*}$; equivalently, $CN=N^{*}C$. Clearly, we need to show that $CP=PC$. According to [25, Proposition 6.9], there exists a Borel subset $\Delta\subseteq\mathbb{C}$ such that $P=E_{N}(\Delta)$. Consequently, Lemma 2.11 implies that $CP=CE_{N}(\Delta)=E_{N}(\Delta)C=PC$.
For the sufficiency, we need to show that for each $T\in\{N\}^{\prime}$, the commutant of $N$, $TM\subseteq M$. Additionally, since $\{N\}^{\prime}$ is a von Neumann algebra, it is spanned by its unitary operators ([9, Page 61, Proposition 13.3 (b)]). Consequently, we can assume without loss of generality that $T$ is unitary. Let $x\in M$ be non-zero, and put $y=Tx$. Then, according to Theorem 2.4, there exists a conjugation $C$ on $H$ satisfying $CNC=N^{*}$ and $Cx=y$. Hence, by hypothesis, $Tx=y=Cx\in M$. Since $x$ was arbitrary, we conclude that $TM\subseteq M$. ∎
The following example shows that the result in Proposition 2.5 cannot be extended to an arbitrary operator in ${\mathcal{B}}(H)$.
Example 2.6.
Let $x$ and $y$ be orthonormal vectors in $H$, and put $T=A+{\rm i}B$ where $A=x\otimes x$ and $B=(x+y)\otimes(x+y)$. Clearly, $A$ and $B$ are self-adjoint operators; moreover, $A=\frac{1}{2}(T+T^{*})$. It follows that if $C$ is a conjugation on $H$ such that $CTC=T^{*}$, then $CAC=A$. Hence, if $M={\rm span}\{x\}$, then $CM=M$. Therefore, the subspace $M$ satisfies the necessary condition in Proposition 2.5; however, $T\in\{T\}^{\prime}$ and $TM\nsubset M$.
Given a conjugation $C$ on $H$, an operator $T\in{\mathcal{B}}(H)$ is said to be $C$-normal if $CTT^{*}C=T^{*}T$. When $T$ is $C$-normal for some conjugation $C$, we say that $T$ is conjugate-normal. It is easy to verify that if $T$ is $C$-symmetric or $C$-skew symmetric, then it is also $C$-normal. Consequently, conjugate-normality is considered a more general type of symmetry. For further details and properties of such operators, the interested reader is referred to [27].
Another consequence of Theorem 2.4 is the following result, which characterizes normal operators in terms of complex symmetric and conjugate-normal ones.
Corollary 2.7.
Let $T\in{\mathcal{B}}(H)$. The following statements are equivalent:
  1. (i)
    $T$ is normal.
  2. (ii)
    For every non-zero $x\in H$, there exists a conjugation $C$ on $H$ such that $T$ is $C$-symmetric and $Cx=x$;
  3. (iii)
    For every non-zero $x\in H$, there exists a conjugation $C$ on $H$ such that $T$ is $C$-normal and $Cx=x$.
Proof.
The implication (i)$\implies$(ii) follows directly from Theorem 2.4 applied to the normal operator $T$ and the vectors $x=y$.
(ii)$\implies$(iii) is obvious.
(iii)$\implies$(ii). Let $x\in H$ be non-zero. Then, by hypothesis, there exists a conjugation $C$ on $H$ such that $Cx=x$ and $CTT^{*}C=T^{*}T$. It follows that
$\langle T^{*}Tx,x\rangle=\langle CTT^{*}Cx,x\rangle=\langle Cx,CCTT^{*}Cx\rangle=\langle x,TT^{*}x\rangle=\langle TT^{*}x,x\rangle.$
Since $x$ was arbitrary, we infer that $T^{*}T=TT^{*}$, the desired equality. ∎
2.2. Proofs of Theorems 2.1, 2.3 and 2.4
The proofs require some preparation and will be provided at the end of this section. In what follows, we denote the set of all complex polynomials in two variables by $\mathbb{C}[x,y]$. We begin with the following technical lemma.
Lemma 2.8.
Let $N\in{\mathcal{B}}(H)$ and $M\in\mathcal{B}(K)$ be normal operators, and let $x,y\in H$ and $h,k\in K$ be vectors such that
$\langle E_{N}(\Delta)x,y\rangle=\langle E_{M}(\Delta)h,k\rangle\quad\mbox{for every Borel subset $\Delta\subseteq\mathbb{C}$}.$
Then, for every $p\in\mathbb{C}[x,y]$, we have $\langle p(N,N^{*})x,y\rangle=\langle p(M,M^{*})h,k\rangle.$
Proof.
Let $p\in\mathbb{C}[x,y]$, and put $f(z)=p(z,\overline{z})\chi_{K}(z)$ where $K=\sigma(N)\cup\sigma(M)$ and $\chi_{K}$ is the characteristic function of $K$. Then, $f$ is a bounded Borel function on $\mathbb{C}$; furthermore, $f(N)=p(N,N^{*})$ and $f(M)=p(M,M^{*})$. This implies that
$\langle p(N,N^{*})x,y\rangle=\langle f(N)x,y\rangle=\int f{\,\rm d}\langle E_{N}(.)x,y\rangle$
and
$\langle p(M,M^{*})h,k\rangle=\langle f(M)h,k\rangle=\int f{\,\rm d}\langle E_{M}(.)h,k\rangle.$
As the complex-valued measures $\langle E_{N}(.)x,y\rangle$ and $\langle E_{M}(.)h,k\rangle$ are presumed equal by hypothesis, we deduce $\int f{\,\rm d}\langle E_{N}(.)x,y\rangle=\int f{\,\rm d}\langle E_{M}(.)h,k\rangle$. Consequently, $\langle p(N,N^{*})x,y\rangle=\langle p(M,M^{*})x,y\rangle.$
Given a normal operator $N\in{\mathcal{B}}(H)$ and a family of vectors $X\subset H$, we denote
$L_{N,X}=\bigvee\big{\{}p(N,N^{*})h:p\in\mathbb{C}[x,y]\mbox{ and }h\in X\big{\}}$
where $\vee$ denotes the closed linear span. It is easy to see that $L_{N,X}$ is a reducing subspace of $N$. The restriction of $N$ to $L_{N,X}$ will be denoted by $N_{X}$.
Proposition 2.9.
Let $N\in{\mathcal{B}}(H)$ and $M\in\mathcal{B}(K)$ be normal operators, and let $X=\{x_{i}\}_{i\in I}\subseteq H$ and $Y=\{y_{i}\}_{i\in I}\subseteq K$ be two families of vectors such that
$\langle E_{N}(\Delta)x_{i},x_{j}\rangle=\langle E_{M}(\Delta)y_{j},y_{i}\rangle\quad\mbox{for all Borel subsets $\Delta\subseteq\mathbb{C}$ and $i,j\in I$}.$
Then, there exists a conjugate-linear surjective isometry $W:L_{N,X}\to L_{M,Y}$ such that
$Wx_{i}=y_{i}\quad\mbox{and}\quad Wp(N_{X},N_{X}^{*})=p(M_{Y},M_{Y}^{*})^{*}W\quad\mbox{for all $i\in I$ and $p\in\mathbb{C}[x,y]$}.$
Proof.
According to Lemma 2.8, for every polynomial $p\in\mathbb{C}[x,y]$ and all $i,j\in I$, we have
(2.1) $\left\langle p(N,N^{*})x_{i},x_{j}\right\rangle=\left\langle p(M,M^{*})y_{j},y_{i}\right\rangle.$
For convenience, put
$H_{0}={\rm span}\big{\{}p(N,N^{*})x_{i}:p\in\mathbb{C}[x,y],i\in I\big{\}}$
and
$K_{0}={\rm span}\big{\{}p(M,M^{*})y_{i}:p\in\mathbb{C}[x,y],i\in I\big{\}},$
and let $W:H_{0}\to K_{0}$ be the conjugate-linear transformation given by
$W(p(N,N^{*})x_{i})=p(M,M^{*})^{*}y_{i}\quad\mbox{for all $p\in\mathbb{C}[x,y]$and $i\in I$.}$
We will first show that $W$ is well-defined. Arbitrarily choose $h\in H_{0}$; then, there exists a finite subset $J$ of $I$ and polynomials $p_{i}\in\mathbb{C}[x,y]$, for $i\in J$, such that $h=\sum_{i\in J}p_{i}(N,N^{*})x_{i}$. Then,
$\|h\|^{2}$ $=\left\langle\sum_{i\in J}p_{i}(N,N^{*})x_{i},\sum_{j\in J}p_{j}(N,N^{*})x_{j}\right\rangle$
$=\sum_{i\in J}\sum_{j\in J}\left\langle p_{i}(N,N^{*})x_{i},p_{j}(N,N^{*})x_{j}\right\rangle$
$=\sum_{i\in J}\sum_{j\in J}\left\langle p_{j}(N,N^{*})^{*}p_{i}(N,N^{*})x_{i},x_{j}\right\rangle$
[…]
$=\sum_{i\in J}\sum_{j\in J}\left\langle p_{j}(M,M^{*})^{*}y_{j},p_{i}(M,M^{*})^{*}y_{i}\right\rangle$
$=\left\langle\sum_{j\in J}p_{j}(M,M^{*})^{*}y_{j},\sum_{i\in J}p_{i}(M,M^{*})^{*}y_{i}\right\rangle=\|Wh\|^{2}.$
Hence, if $h,k\in H_{0}$ are such that $\|h-k\|=0$, we get $\|Wh-Wk\|=\|W(h-k)\|=\|h-k\|=0$. Therefore, $W$ is a well-defined conjugate-linear isometry. Moreover, by its construction, $W$ is surjective. By density of $H_{0}$ and $K_{0}$ in $L_{N,X}$ and $L_{M,Y}$, respectively, it follows that $W$ can be extended to a conjugate-linear surjective isometry from $L_{N,X}$ to $L_{M,Y}$, which we also denote by $W$. Furthermore, it can be seen that $Wx_{i}=y_{i}$ for every $i\in I$.
Now, fix $p\in\mathbb{C}[x,y]$ and let us check that $Wp(N_{X},N_{X}^{*})=p(M_{Y},M_{Y}^{*})^{*}W$. Let $q\in\mathbb{C}[x,y]$, $i\in I$, and put $h=q(N,N^{*})x_{i}$. Then,
$Wp(N_{X},N_{X}^{*})h=Wp(N,N^{*})h$ $=Wp(N,N^{*})q(N,N^{*})x_{i}=p(M,M^{*})^{*}q(M,M^{*})^{*}y_{i}$
$=p(M,M^{*})^{*}Wq(N,N^{*})x_{i}=p(M_{Y},M_{Y}^{*})^{*}Wh.$
Since $q$ and $i$ were arbitrary, and $Wp(N_{X},N_{X}^{*})$ and $p(M_{Y},M_{Y}^{*})^{*}W$ are both conjugate-linear, we obtain $Wp(N_{X},N_{X}^{*})h=p(M_{Y},M_{Y}^{*})^{*}Wh$ for every $h\in H_{0}$. Consequently, by density of $H_{0}$ in $L_{N,X}$, we conclude that $Wp(N_{X},N_{X}^{*})=p(M_{Y},M_{Y}^{*})^{*}W$. This completes the proof. ∎
Theorem 2.10.
Let $N\in{\mathcal{B}}(H)$ be normal, and let $X=\{x_{i}\}_{i\in I}$ and $Y=\{y_{i}\}_{i\in I}$ be two orthogonal sets of non-zero vectors in $H$. Let $M\in\{N,-N\}$ and assume that for every Borel subset $\Delta\subseteq\mathbb{C}$ and all $i,j\in I$, we have
  1. (a)
    $\langle E_{N}(\Delta)x_{i},x_{j}\rangle=\langle E_{M}(\Delta)y_{j},y_{i}\rangle$;
  2. (b)
    $\langle E_{N}(\Delta)x_{i},y_{j}\rangle=\langle E_{M}(\Delta)x_{j},y_{i}\rangle$.
Then, there exists a conjugation $C$ on $L_{N,X\cup Y}$ such that
$CN_{X\cup Y}C=M_{X\cup Y}^{*}\quad\mbox{and}\quad Cx_{i}=y_{i}\;\;\mbox{for every }i\in I.$
Proof.
First, note that $L_{N,X\cup Y}=L_{M,X\cup Y}$; hence, $N_{X\cup Y}$ and $M_{X\cup Y}$ are both operators on $L_{N,X\cup Y}$, and so the statement $CN_{X\cup Y}C=M_{X\cup Y}^{*}$ makes sense. Additionally, without loss of generality, we may assume that $X$ and $Y$ are orthonormal subsets of $H$. In fact, by condition (a),
$r_{i}:=\|x_{i}\|=\sqrt{\langle E_{N}(\mathbb{C})x_{i},x_{i}\rangle}=\sqrt{\langle E_{M}(\mathbb{C})y_{i},y_{i}\rangle}=\|y_{i}\|$
for every $i\in I$; thus, the orthonormal subsets $\{r_{i}^{-1}x_{i}\}_{i\in I}$ and $\{r_{i}^{-1}y_{i}\}_{i\in I}$ satisfy conditions (a) and (b).
Using condition (a), we infer from Proposition 2.9 that there exists a conjugate-linear surjective isometry $W:L_{N,X}\to L_{M,Y}$ such that
(2.2) $Wx_{i}=y_{i}\;\mbox{and}\;Wp(N_{X},N_{X}^{*})=p(M_{Y},M_{Y}^{*})^{*}W\;\mbox{ for all $i\in I$ and $p\in\mathbb{C}[x,y]$.}$
In particular,
(2.3) $WN_{X}=M_{Y}^{*}W\quad\mbox{and}\quad WN_{X}^{*}=M_{Y}W.$
Choose an arbitrary orthonormal basis $\{e_{j}\}_{j\in J}$ of $L_{N,X}$ that contains $X$ and, for each $j\in J$, put $f_{j}=We_{j}$. Then, since $W$ is a conjugate-linear surjective isometry, the set $\{f_{j}\}_{j\in J}$ must be an orthonormal basis of $L_{M,Y}$.
Note that, for each $j\in J$, there exist sequences of finite subsets $I_{j,n}\subseteq I$ and polynomials $p_{j,n,k}\in\mathbb{C}[x,y]$, for $n\in\mathbb{Z}^{+}$ and $k\in I_{j,n}$, such that
(2.4) $e_{j}=\lim_{n}\sum_{k\in I_{j,n}}p_{j,n,k}(N,N^{*})x_{k};$
hence, by (2.2),
(2.5) $f_{j}=We_{j}=\lim_{n}\sum_{k\in I_{j,n}}p_{j,n,k}(M,M^{*})^{*}y_{k}.$
Here, $\mathbb{Z}^{+}$ stands for the set of positive integers.
Claim: For all $i,j\in J$, we have $\langle e_{i},f_{j}\rangle=\langle e_{j},f_{i}\rangle$. First, note that in the light of condition (b), Lemma 2.8 gives
(2.6) $\langle p(N,N^{*})x_{i},y_{j}\rangle=\langle p(M,M^{*})x_{j},y_{i}\rangle\quad\mbox{for all $i,j\in I$ and $p\in\mathbb{C}[x,y]$}.$
Now, fix $i,j\in J$ and adopt the notations in (2.4) and (2.5); then
$\left\langle e_{i},f_{j}\right\rangle$ $=\lim_{n}\left\langle\sum_{k\in I_{i,n}}p_{i,n,k}(N,N^{*})x_{k},\sum_{l\in I_{j,n}}p_{j,n,l}(M,M^{*})^{*}y_{l}\right\rangle$
$=\lim_{n}\sum_{k\in I_{i,n}}\sum_{l\in I_{j,n}}\left\langle p_{i,n,k}(N,N^{*})x_{k},p_{j,n,l}(M,M^{*})^{*}y_{l}\right\rangle$
$=\lim_{n}\sum_{k\in I_{i,n}}\sum_{l\in I_{j,n}}\left\langle p_{j,n,l}(M,M^{*})p_{i,n,k}(N,N^{*})x_{k},y_{l}\right\rangle.$
Hence, using (2.6), if $M=N$, we can see that
$\left\langle e_{i},f_{j}\right\rangle=\lim_{n}\sum_{k\in I_{i,n}}\sum_{l\in I_{j,n}}\left\langle p_{j,n,l}(N,N^{*})p_{i,n,k}(N,N^{*})x_{l},y_{k}\right\rangle,$
and if $M=-N$, we obtain
$\left\langle e_{i},f_{j}\right\rangle=\lim_{n}\sum_{k\in I_{i,n}}\sum_{l\in I_{j,n}}\left\langle p_{j,n,l}(N,N^{*})p_{i,n,k}(-N,-N^{*})x_{l},y_{k}\right\rangle.$
In both cases, we have
$\left\langle e_{i},f_{j}\right\rangle=\lim_{n}\sum_{k\in I_{i,n}}\sum_{l\in I_{j,n}}\left\langle p_{j,n,l}(N,N^{*})p_{i,n,k}(M,M^{*})x_{l},y_{k}\right\rangle.$
Consequently, we obtain
$\left\langle e_{i},f_{j}\right\rangle$ $=\lim_{n}\sum_{k\in I_{i,n}}\sum_{l\in I_{j,n}}\left\langle p_{j,n,l}(N,N^{*})x_{l},p_{i,n,k}(M,M^{*})^{*}y_{k}\right\rangle$
$=\lim_{n}\left\langle\sum_{l\in I_{j,n}}p_{j,n,l}(N,N^{*})x_{l},\sum_{k\in I_{i,n}}p_{i,n,k}(M,M^{*})^{*}y_{k}\right\rangle$
[…]
This proves the claim.
Note that $L_{N,X\cup Y}=\vee\left(L_{N,X}+L_{M,Y}\right)$; hence, $\{e_{j}\}_{j\in J}$ and $\{f_{j}\}_{j\in J}$ are both orthonormal subsets of $L_{N,X\cup Y}$. By [29, Theorem 2.1], there exists a conjugation $C$ on $L_{N,X\cup Y}$ such that $Ce_{j}=f_{j}$ for every $j\in J$. In particular, since $\{e_{j}\}_{j\in J}$ contains $X$, we get from (2.2) that
$Cx_{i}=Wx_{i}=y_{i}\quad\mbox{for every $i\in I$.}$
It remains to check that $CN_{X\cup Y}C=M_{X\cup Y}^{*}$. Noting that, for every $j\in J$, $Ce_{j}=f_{j}=We_{j}$ and $Cf_{j}=C^{2}e_{j}=e_{j}=W^{-1}f_{j}$, and taking into account that $C$, $W$ and $W^{-1}$ are conjugate-linear, it follows that $W=C_{1}$ and $W^{-1}=C_{2}$ where $C_{1}$ and $C_{2}$ are defined as
$\begin{array}[]{rcl}C_{1}:L_{N,X}&\to&L_{M,Y}\\ h&\mapsto&Ch\end{array}\quad\mbox{and}\quad\begin{array}[]{rcl}C_{2}:L_{M,Y}&\to&L_{N,X}\\ h&\mapsto&Ch\end{array}.$
Combining these facts with (2.3), we get that for every $j\in J$,
$CM_{X\cup Y}Ce_{j}=CM_{X\cup Y}We_{j}=CM_{Y}We_{j}=CWN_{X}^{*}e_{j}=C_{2}WN_{X}^{*}e_{j}=N_{X}^{*}e_{j}=N_{X\cup Y}^{*}e_{j}$
and
$CN_{X\cup Y}Cf_{j}=CN_{X\cup Y}e_{j}=CN_{X}e_{j}=C_{1}N_{X}e_{j}=WN_{X}e_{j}=M_{Y}^{*}We_{j}=M_{Y}^{*}f_{j}=M_{X\cup Y}^{*}f_{j}.$
Suppose that $M=N$; it follows that $CN_{X\cup Y}Ce_{j}=N_{X\cup Y}^{*}e_{j}$ and $CN_{X\cup Y}Cf_{j}=N_{X\cup Y}^{*}f_{j}$ for every $j\in J$. Hence, since $L_{N,X\cup Y}=\vee\left(L_{N,X}+L_{M,Y}\right)=\vee\left(\{e_{j},f_{j}\}_{j\in J}\right)$, we obtain $CN_{X\cup Y}C=N_{X\cup Y}^{*}$. On the other hand, if $M=-N$, we get $CN_{X\cup Y}Ce_{j}=-N_{X\cup Y}^{*}e_{j}$ and $CN_{X\cup Y}Cf_{j}=-N_{X\cup Y}^{*}f_{j}$ for every $j\in J$. Therefore, $CN_{X\cup Y}C=-N_{X\cup Y}^{*}$. In both cases, we have the desired result. This completes the proof of the theorem.
The following lemma will be very useful throughout the paper.
Lemma 2.11.
Let $N\in\mathcal{B}(H)$ and $M\in\mathcal{B}(K)$ be normal operators, and let $T:H\to K$ be a bounded linear (or conjugate-linear) operator. The following statements hold:
  1. (i)
    If $T$ is linear (resp. conjugate-linear) and $TN=MT$ (resp. $TN=M^{*}T$), then
    $TE_{N}(\Delta)=E_{M}(\Delta)T\quad\mbox{for every Borel subset }\Delta\subseteq\mathbb{C}.$
  2. (ii)
    If $T$ is linear (resp. conjugate-linear) and $TN=-MT$ (resp. $TN=-M^{*}T$), then
    $TE_{N}(\Delta)=E_{M}(-\Delta)T\quad\mbox{for every Borel subset }\Delta\subseteq\mathbb{C}.$
Proof.
(i) Suppose first that $T$ is linear and $TN=MT$, and put
$L=\begin{bmatrix}N&0\\ 0&M\end{bmatrix}\begin{array}[]{l}H\\ K\end{array}\quad\mbox{and}\quad A=\begin{bmatrix}0&0\\ T&0\end{bmatrix}\begin{array}[]{l}H\\ K\end{array}.$
Then, $L$ and $A$ are bounded linear operators acting on the Hilbert space $H\oplus K$. Furthermore, one can verify that $AL=LA$. Since $L$ is normal, by the Fuglede Theorem, $AE_{L}(\Delta)=E_{L}(\Delta)A$ for every Borel subset $\Delta\subseteq\mathbb{C}$.
Fix a Borel subset $\Delta\subseteq\mathbb{C}$, and note that
$E_{L}(\Delta)=\begin{bmatrix}E_{N}(\Delta)&0\\ 0&E_{M}(\Delta)\end{bmatrix}\begin{array}[]{l}H\\ K\end{array}.$
Therefore, the matrix equality $AE_{L}(\Delta)=E_{L}(\Delta)A$ implies $TE_{N}(\Delta)=E_{M}(\Delta)T$.
Assume now that $T$ is a conjugate-linear operator such that $TN=M^{*}T$, and let $\Delta\subseteq\mathbb{C}$ be a Borel subset. Since $M$ is normal, there exists a conjugation $C$ on $K$ such that $CMC=M^{*}$. Hence, $TN=M^{*}T=(CMC)T$, and since $C^{2}=I$, we obtain
(2.7) $(CT)N=C^{2}MCT=M(CT).$
Clearly, $CT:H\to K$ is a bounded linear operator. By our previous result, we have
(2.8) $CTE_{N}(\Delta)=E_{M}(\Delta)CT.$
By [8, Page 287, Lemma 8.7], $E_{M}(\Delta)=\chi_{\Delta}(M)\in W^{*}(M)$. Hence, there exists a net $\{p_{i}\}_{i}$ in $\mathbb{C}[x,y]$ such that $p_{i}(M,M^{*})$ converges to $E_{M}(\Delta)$ in the weak operator topology. One can easily check that $Cp_{i}(M,M^{*})C=p_{i}(M,M^{*})^{*}$. Consequently, taking the limit in the weak operator topology, we get $CE_{M}(\Delta)C=E_{M}(\Delta)^{*}=E_{M}(\Delta).$ Therefore, from (2.8), we obtain
$TE_{N}(\Delta)=C^{2}TE_{N}(\Delta)=CE_{M}(\Delta)CT=E_{M}(\Delta)T,$
the desired equality.
(ii) Suppose that $TN=-MT$ if $T$ is linear and that $TN=-M^{*}T$ if $T$ is conjugate-linear. Applying the first part of lemma to $N$ and $-M$, we get $TE_{N}(\Delta)=E_{-M}(\Delta)T$ for every Borel subset $\Delta\subseteq\mathbb{C}$. Since $E_{-M}(\Delta)=\chi_{\Delta}(-M)$ and $\chi_{\Delta}(-z)=\chi_{-\Delta}(z)$ for every $z\in\mathbb{C}$, we have
$TE_{N}(\Delta)=E_{-M}(\Delta)T=\chi_{\Delta}(-M)T=\chi_{-\Delta}(M)T=E_{M}(-\Delta)T.$
This completes the proof. ∎
We are now in a position to prove Theorems 2.1 and 2.3 and 2.4.
Proof of Theorem 2.1.
The implication (i)$\implies$(ii) follows directly from the properties of $C$ as a conjugate-linear isometry, satisfying $Cx_{i}=y_{i}$, $Cy_{i}=C^{2}x_{i}=x_{i}$ and $CT=CTC^{2}=T^{*}C$ for all $i\in I$ and $T\in\mathcal{M}$.
(ii)$\implies$(iii). Let $V_{0}:{\rm ker}(V)^{\perp}\to{\rm ran}(V)$ be the invertible conjugate-linear isometry given by $V_{0}h=Vh$. Define the conjugate-linear map $V^{\#}:H\to H$ by $V^{\#}h=V_{0}^{-1}h$ if $h\in{\rm ran}(V)$ and $V^{\#}h=0$ if $h\in{\rm ran}(V)^{\perp}$. Then, $V^{\#}V$ is the orthogonal projection onto ${\rm ker}(V)^{\perp}$. Moreover, using the polarization identity, it follows that $\langle Vh,k\rangle=\langle V^{\#}k,h\rangle$ for all $h\in{\rm ker}(V)^{\perp}$ and $k\in{\rm ran}(V)$. Thus,
(2.9) $\langle Vh,k\rangle=\langle V^{\#}k,h\rangle\quad\mbox{for all }h,k\in H.$
Now, let $P\in\mathcal{M}$ be an orthogonal projection. Using the above relation, for all $i,j\in I$, we obtain
(2.10) $\begin{split}\langle Py_{j},y_{i}\rangle=\langle PVx_{j},Vx_{i}\rangle&=\langle VPx_{j},Vx_{i}\rangle=\langle V^{\#}Vx_{i},Px_{j}\rangle\\ &=\langle x_{i},Px_{j}\rangle=\langle Px_{i},x_{j}\rangle,\end{split}$
and similarly, we have
(2.11) $\begin{split}\langle Px_{j},y_{i}\rangle=\langle PVy_{j},Vx_{i}\rangle&=\langle VPy_{j},Vx_{i}\rangle=\langle V^{\#}Vx_{i},Py_{j}\rangle\\ &=\langle x_{i},Py_{j}\rangle=\langle Px_{i},y_{j}\rangle.\end{split}$
(iii)$\implies$(ii). According to [10, Page 52, Lemma II.2.8], there exists a self-adjoint operator $A\in{\mathcal{B}}(H)$ such that $\mathcal{M}=W^{*}(A)$. It suffices to show that there exists a conjugation $C$ on $H$ such that $CAC=A$ and $Cx_{i}=y_{i}$ for every $i\in I$. In fact, any polynomial in $A$ is $C$-symmetric; moreover, since the class of $C$-symmetric operators is closed in the weak operator topology, every operator in $W^{*}(A)$, and thus in $\mathcal{M}$, is $C$-symmetric.
Let $\Delta\subseteq\mathbb{C}$ be a Borel subset. Then, by [8, Page 287, Lemma 8.7], the orthogonal projection $E_{A}(\Delta)$ belongs to $\mathcal{M}$. Therefore, by hypothesis, for all $i,j\in I$, we have $\langle E_{A}(\Delta)x_{i},x_{j}\rangle=\langle E_{A}(\Delta)y_{j},y_{i}\rangle$ and $\langle E_{A}(\Delta)x_{i},y_{j}\rangle=\langle E_{A}(\Delta)x_{j},y_{i}\rangle$.
For convenience, put $Z=\{x_{i},y_{i}\}_{i\in I}$, $H_{0}=L_{A,Z}$ and $H_{1}=L_{A,Z}^{\perp}$. Then, $H_{0}$ and $H_{1}$ are reducing subspaces of $A$ and $Z\subset H_{0}$. According to Theorem 2.10, there exists a conjugation $C_{0}$ on $H_{0}$ such that $C_{0}A_{|H_{0}}C=A_{0}$ and $C_{0}x_{i}=y_{i}$ for every $i\in I$. Since $A_{|H_{1}}$ is normal, there exists a conjugation $C_{1}$ on $H_{1}$ such that $C_{1}A_{|H_{1}}C_{1}=A_{|H_{1}}$. Clearly, $C=C_{0}\oplus C_{1}$ is a conjugation on $H$; additionally, $CAC=A$ and $Cx_{i}=C_{0}x_{i}=y_{i}$ for every $i\in I$. ∎
Proof of Theorem 2.3.
The implication (i)$\implies$(ii) is obvious, and the implication (ii)$\implies$(iii) follows a similar argument as (i)$\implies$(ii) in Theorem 2.1.
To show that (iii)$\implies$(iv), first note that by Lemma 2.11, for all Borel subsets $\Delta\subseteq\mathbb{C}$ and $k\in K$, we have $VE_{N_{k}}(\Delta)=E_{N_{k}}(\Delta)V$. Hence, for any finite subset $K_{0}\subseteq K$ and Borel subsets $\{\Delta_{k}\}_{k\in K_{0}}$ of $\mathbb{C}$, we have $V\left(\prod_{k\in K_{0}}E_{N_{k}}(\Delta_{k})\right)=\left(\prod_{k\in K_{0}}E_{N_{k}}(\Delta_{k})\right)V$. Put $P=\prod_{k\in K_{0}}E_{N_{k}}(\Delta_{k})$; the remainder of the proof follows similarly to (2.10) and (2.11).
(iv)$\implies$(i). For convenience, put $\mathcal{M}=W^{*}(N_{k}:k\in K)$. By the Fuglede Theorem, we have $N_{k}N_{l}^{*}=N_{l}^{*}N_{k}$ for any $k,l\in K$. Therefore, it is not hard to see that $\mathcal{M}$ is an abelian von Neumann algebra. Hence, by Theorem 2.1, we need to show that for all orthogonal projections $P\in\mathcal{M}$ and $i,j\in I$,
(2.12) $\langle Px_{i},x_{j}\rangle=\langle Py_{j},y_{i}\rangle\quad\mbox{and}\quad\langle Px_{i},y_{j}\rangle=\langle Px_{j},y_{i}\rangle.$
First, we show that $\mathcal{M}={\rm cl-WOT}\mathcal{M}_{0}$; i.e., the closure in the weak operator topology of
$\mathcal{M}_{0}={\rm span}\left\{\prod_{k\in K_{0}}E_{N_{k}}(\Delta_{k}):K_{0}\subseteq K\mbox{ is finite}\mbox{ and }\Delta_{k}\mbox{ is a Borel subset of }\mathbb{C}\right\}.$
By [8, Page 287, Lemma 8.7], any spectral projection $E_{N_{k}}(\Delta)=\chi_{\Delta}(N_{k})\in W^{*}(N_{k})\subseteq\mathcal{M}$. Hence, $\mathcal{M}_{0}\subseteq\mathcal{M}$, and consequently ${\rm cl-WOT}\mathcal{M}_{0}\subseteq\mathcal{M}$. Let us show the other containment. According to spectral theory, for every $k\in K$, we have
$N_{k}\in\bigvee\big{\{}E_{N_{k}}(\Delta):\Delta\mbox{ is a Borel subset of }\mathbb{C}\big{\}}.$
This implies that $N_{k}\in{\rm cl-WOT}\mathcal{M}_{0}$ for every $k\in K$. It is easy to see that ${\rm cl-WOT}\mathcal{M}_{0}$ is a self-adjoint subalgebra of ${\mathcal{B}}(H)$ that contains the identity. Moreover, since, by the Fuglede Theorem, the spectral measures $\{E_{N_{k}}\}_{k\in K}$ mutually commute and $E_{N_{k}}(\Delta)E_{N_{k}}(D)=E_{N_{k}}(\Delta\cap D)$ for any Borel subsetes $\Delta$ and $D$ of $\mathbb{C}$, one can easily show that $\mathcal{M}_{0}$, and thus ${\rm cl-WOT}\mathcal{M}_{0}$, is also multiplicative. Hence, ${\rm cl-WOT}\mathcal{M}_{0}$ is a von Neumann algebra containing $\{N_{k}\}_{k\in K}$. Therefore, by definition of $\mathcal{M}$, we must have $\mathcal{M}\subseteq{\rm cl-WOT}\mathcal{M}_{0}$.
Now, by hypothesis, any $T\in\mathcal{M}_{0}$ satisfies the equalities (2.12). Then, by taking limits, in the weak operator topology, of operators in $\mathcal{M}_{0}$, we obtain that (2.12) is true for every $T\in\mathcal{M}$; in particular, for orthogonal projections in $\mathcal{M}$. ∎
Proof of Theorem 2.4.
The implication (ii)$\implies$(iii) is obvious.
The implication (iv)$\implies$(i) follows from Theorem 2.3 (iv)$\implies$(i). In fact, for any finite subset $K_{0}\subseteq K$ and Borel subsets $\{\Delta_{k}\}_{k\in K_{0}}$ of $\mathbb{C}$, the orthogonal projections $\{E_{N_{k}}(\Delta_{k})\}_{k\in K_{0}}$ are mutually commuting. Thus, their product $P$ is also an orthogonal projection; which gives
$\langle Px,x\rangle=\langle P^{2}x,x\rangle=\langle Px,Px\rangle=\|Px\|^{2}=\|Py\|^{2}=\langle Py,y\rangle.$
(iii)$\implies$(iv). Let $\Delta\subseteq\mathbb{C}$ be a Borel subset and $K_{0}\subseteq K$ a finite set. Then, $V\left(\prod_{k\in K_{0}}E_{N_{k}}(\Delta)\right)=\left(\prod_{k\in K_{0}}E_{N_{k}}(\Delta)\right)V$. Noting that $V^{*}V$ is the orthogonal projection onto ${\rm ker}(V)^{\perp}$ and letting $P=\prod_{k\in K_{0}}E_{N_{k}}(\Delta)$, we get
$\|Py\|^{2}=\langle Py,y\rangle=\langle PVx,Vx\rangle=\langle Px,V^{*}Vx\rangle=\langle Px,x\rangle=\|Px\|^{2}.$
(i)$\implies$(v). Clearly, $\|x\|=\|Cx\|=\|y\|$. It is easy to check that $C(x\otimes y)C=(Cx)\otimes(Cy)$; hence, for any $k\in K$ and $\lambda\in\mathbb{C}$, we have
$C\left(N_{k}+\lambda x\otimes y\right)C=CN_{K}C+\overline{\lambda}y\otimes x=N_{k}^{*}+\overline{\lambda}(x\otimes y)^{*}=\left(N_{k}+\lambda x\otimes y\right)^{*};$
that is, $N_{k}+\lambda x\otimes y$ is $C$-symmetric.
(v)$\implies$(ii). Since the class of $C$-symmetric operators forms a subspace of ${\mathcal{B}}(H)$, we obtain, for every $k\in K$, $CN_{k}C=N_{k}^{*}$ and $C(x\otimes y)C=(x\otimes y)^{*}=y\otimes x$. Given that $x$ and $y$ have the same norm, the latter equality implies the existence of a unimodular $\alpha\in\mathbb{C}$ such that $Cx=\alpha y$.
Applying Theorem 2.3 (iv)$\implies$(ii) to $\{x_{i}\}_{i}=\{y_{i}\}_{i}=\{x\}$, we obtain a conjugation $J$ on $H$ such that $JN_{k}J=N_{k}^{*}$ for every $k\in K$ and $Jx=x$. Put $U=\overline{\alpha}CJ$; then, since $C$ and $J$ are both surjective isometries and $\alpha$ is unimodular, the linear operator $U:H\to H$ must be unitary. Moreover, we have
$UN_{k}=\overline{\alpha}CJN_{k}=\overline{\alpha}CN_{k}^{*}J=\overline{\alpha}N_{k}CJ=N_{k}U$
and $Ux=\overline{\alpha}CJx=\overline{\alpha}Cx=\overline{\alpha}\alpha y=y$. This completes the proof of the theorem. ∎
3. Interpolation theorems for conjugations : C-skew symmetric versions
A conjugate-linear operator $V$ on $H$ is said to be anti-unitary if it is a surjective isometry. Clearly, every conjugation on $H$ is anti-unitary.
For two operators $A$ and $B$ on Hilbert spaces, we denote $A\cong B$ if there exists an isometric isomorphism $U$ between the underlying Hilbert spaces such that $UAU^{-1}=B$; or equivalently, $UA=BU$. It is worth mentioning that both complex symmetry and skew symmetry are preserved under the equivalence relation $\cong$.
For $k\in\{1,2,\dots,\infty\}$, where $\infty=\aleph_{0}$, $H^{(k)}$ denotes the direct sum of $H$ with itself $k$ times. If $T\in{\mathcal{B}}(H)$, then $T^{(k)}\in\mathcal{B}(H^{(k)})$ is the direct sum of $T$ with itself $k$ times.
3.1. Main results
Theorem 3.1.
Let $N\in{\mathcal{B}}(H)$ be normal, and let $Z=X\cup Y$ where $X=\{x_{i}\}_{i\in I}$ and $Y=\{y_{i}\}_{i\in I}$ are orthogonal sets of non-zero vectors in $H$. Then, the following statements are equivalent:
  1. (i)
    There exists a conjugation $C$ on $H$ such that $CNC=-N^{*}$ and $Cx_{i}=y_{i}$ for every $i\in I$.
  2. (ii)
    There exists an anti-unitary operator $V$ on $H$ such that $VNV^{-1}=-N^{*}$, $Vx_{i}=y_{i}$ and $Vy_{i}=x_{i}$ for every $i\in I$.
  3. (iii)
    $N_{|L_{N,Z}^{\perp}}$ is skew-symmetric and, for all Borel subsets $\Delta\subseteq\mathbb{C}$ and $i,j\in I$, we have
    $\langle E_{N}(\Delta)x_{i},x_{j}\rangle=\langle E_{N}(-\Delta)y_{j},y_{i}\rangle\quad\mbox{and}\quad\langle E_{N}(\Delta)x_{i},y_{j}\rangle=\langle E_{N}(-\Delta)x_{j},y_{i}\rangle.$
Before stating the second main result of this section, we recall some facts about normal operators on separable complex Hilbert spaces. Given a finite measure with compact support $\mu$ on $\mathbb{C}$, denote by $M_{\mu}$ the operator defined on $L^{2}(\mu)$ by $[M_{\mu}f](z)=zf(z)$ for all $f\in L^{2}(\mu)$. According to spectral multiplicity theory (see, for instance, [8, Page 298, Theorem 10.16]), each normal operator $N$ acting on $H$ is unitarily equivalent to
(3.1) $M_{\mu_{\infty}}^{(\infty)}\oplus\bigoplus_{1\leq i<\infty}M_{\mu_{i}}^{(i)}$
where $\mu_{i}$, $1\leq i\leq\infty$ (some of which may be zero), are mutually singular finite measures on Borel subsets of $\mathbb{C}$, each with compact support. Moreover, the previous decomposition is unique in the sense that if $M$ is another normal operator with corresponding measures $\nu_{i}$, $1\leq i\leq\infty$, then $N\cong M$ if and only if, for every $i\in\{1,2,\dots,\infty\}$,
(3.2) $\mu_{i}(\Delta)=0\;\;\Longleftrightarrow\;\;\nu_{i}(\Delta)=0\quad\mbox{for every Borel subset }\Delta\subseteq\mathbb{C}.$
In [17, Theorem 1.11], Li et al. proved that $N$ is skew-symmetric if and only if the sequence $\{\mu_{i}\}_{1\leq i\leq\infty}$ can be chosen so that, for every $i\in\{1,2,\dots,\infty\}$,
(3.3) $\mu_{i}(\Delta)=\mu_{i}(-\Delta)\quad\mbox{for every Borel subset }\Delta\subseteq\mathbb{C}.$
Given a normal operator $N\in{\mathcal{B}}(H)$, we write $m(N)<\infty$ if the measure $\mu_{\infty}$ in (3.1) is zero; in other words, $M_{\mu_{\infty}}^{(\infty)}$ can be omitted from the decomposition (3.1). This definition makes sense due to the uniqueness property described in (3.2).
Theorem 3.2.
Let $N\in{\mathcal{B}}(H)$ be a skew-symmetric normal operator, and let $Z$ be the set in Theorem 3.1 such that $m\left({N_{Z}}_{|{\rm ker}(N_{Z})^{\perp}}\right)<\infty$. Then, the following statements are equivalent:
  1. (i)
    There exists a conjugation $C$ on $H$ such that $CNC=-N^{*}$ and $Cx_{i}=y_{i}$ for every $i\in I$.
  2. (ii)
    For every Borel subset $\Delta\subseteq\mathbb{C}$ and all $i,j\in I$, we have
    $\langle E_{N}(\Delta)x_{i},x_{j}\rangle=\langle E_{N}(-\Delta)y_{j},y_{i}\rangle\quad\mbox{and}\quad\langle E_{N}(\Delta)x_{i},y_{j}\rangle=\langle E_{N}(-\Delta)x_{j},y_{i}\rangle.$
In particular, if $I$ is finite, then $m\left({N_{Z}}_{|{\rm ker}(N_{Z})^{\perp}}\right)<\infty$ and the previous statements are equivalent.
The reader can see from Theorem 3.1 that the implication (i)$\implies$(ii), in the previous theorem, is true without the assumption that $m\left({N_{Z}}_{|{\rm ker}(N_{Z})^{\perp}}\right)<\infty$. However, the following example shows that this condition is not superfluous for the implication (ii)$\implies$(i).
Example 3.3.
Suppose that $H=K_{1}\oplus K_{2}$, where $K_{1}$ and $K_{2}$ are infinite-dimensional separable complex Hilbert spaces, and let $\{e_{i}\}_{i\geq 1}$ and $\{f_{i}\}_{i\geq 1}$ be orthonormal bases of $K_{1}$ and $K_{2}$, respectively. For each $i\geq 1$, put $x_{i}=e_{i}\oplus 0$ and $y_{i}=0\oplus f_{i+1}$. Let $N$ be the normal operator on $H$ given by $N=I_{K_{1}}\oplus-I_{K_{2}}$. Clearly, $N\cong-N$; hence, by Lemma 3.5, $N$ is skew-symmetric.
For convenience, put $H_{1}=\vee\{x_{i}:i\geq 1\}$, $H_{2}=\vee\{y_{i}:i\geq 1\}$ and $H_{3}=\vee\{f_{1}\}$. Then, $H=H_{1}\oplus H_{2}\oplus H_{3}$ and $L_{N,Z}=H_{1}\oplus H_{2}$. Furthermore,
${N_{Z}}_{|{\rm ker}(N_{Z})^{\perp}}={N_{Z}}=\begin{bmatrix}I_{H_{1}}&0\\ 0&-I_{H_{2}}\end{bmatrix}\begin{array}[]{l}H_{1}\\ H_{2}\end{array}\cong M_{\delta_{1}}^{(\infty)}\oplus M_{\delta_{-1}}^{(\infty)}\cong\left(M_{\delta_{1}}\oplus M_{\delta_{-1}}\right)^{(\infty)}\cong M_{\mu}^{(\infty)}$
where $\mu=\delta_{1}+\delta_{-1}$ and $\delta_{a}$ denotes the Dirac measure, at $a\in\mathbb{C}$, on Borel subsets of $\mathbb{C}$. Thus, $m\left({N_{Z}}_{|{\rm ker}(N_{Z})^{\perp}}\right)\nless\infty$.
Let us check that $N$ satisfies the conditions in Theorem 3.2 (ii). One can easily see that, for every Borel subset $\Delta\subseteq\mathbb{C}$, the spectral projections $E_{N}(\Delta)$ and $E_{N}(-\Delta)$ fall into one of the following possibilities:
  1. (a)
    $E_{N}(\Delta)=E_{N}(-\Delta)=0$;
  2. (b)
    $E_{N}(\Delta)=E_{N}(-\Delta)=I_{H}$;
  3. (c)
    $E_{N}(\Delta)=I_{K_{1}}\oplus 0$ and $E_{N}(-\Delta)=0\oplus I_{K_{2}}$;
  4. (d)
    $E_{N}(\Delta)=0\oplus I_{K_{2}}$ and $E_{N}(-\Delta)=I_{K_{1}}\oplus 0$.
In all cases, we have
$\langle E_{N}(\Delta)x_{i},x_{j}\rangle=\langle E_{N}(-\Delta)y_{j},y_{i}\rangle\quad\mbox{and}\quad\langle E_{N}(\Delta)x_{i},y_{j}\rangle=\langle E_{N}(-\Delta)x_{j},y_{i}\rangle.$
Now, let us show that there exists no conjugation $C$ on $H$ satisfying the first assertion of Theorem 3.2. Suppose for contradiction that such a conjugation exists. Then,
$CH_{1}=C\vee\{x_{i}:i\geq 1\}=\vee\{Cx_{i}:i\geq 1\}=\vee\{y_{i}:i\geq 1\}=H_{2},$
and since $C^{2}=I_{H}$, we also have $CH_{2}=H_{1}$. As $C$ is a surjective isometry, it follows that $CH_{3}=C(H_{1}+H_{2})^{\perp}=(H_{1}+H_{2})^{\perp}=H_{3}$. Therefore, $C$ takes the form
$C=\begin{bmatrix}0&*&0\\ *&0&0\\ 0&0&C_{3}\end{bmatrix}\begin{array}[]{l}H_{1}\\ H_{2}\\ H_{3}\end{array}$
where $C_{3}$ is a conjugation on $H_{3}$. Finally, since $H_{3}$ reduces $N$ and $CNC=-N^{*}$, we obtain $-I_{H_{3}}=C_{3}(-I_{H_{3}})C_{3}=C_{3}N_{|H_{3}}C_{3}=-N_{|H_{3}}^{*}=I_{H_{3}}$; a contradiction.
When $X=\{x\}$, $Y=\{y\}$, and $\{x,y\}$ are linearly independent, we can add the following equivalent perturbation-type statement to Theorem 3.1.
Theorem 3.4.
Let $N\in{\mathcal{B}}(H)$ be normal, and let $\{x,y\}$ be linearly independent vectors in $H$. The following statements are equivalent:
  1. (i)
    There exists a conjugation $C$ on $H$ satisfying $CNC=-N^{*}$ and $Cx=y$.
  2. (ii)
    There exists a conjugation $C$ on $H$ such that $N+\lambda(x\otimes x-y\otimes y)$ is $C$-skew symmetric for every $\lambda\in\mathbb{C}$.
The proof of the previous theorem does not require preliminary results and can be presented here.
Proof of Theorem 3.4.
(i)$\implies$(ii). Let $C$ be a conjugation satisfying (i). Then, for every $\lambda\in\mathbb{C}$, we have
$C\big{(}N+\lambda(x\otimes x-y\otimes y)\big{)}C$ $=CNC+\overline{\lambda}\big{(}(Cx)\otimes(Cx)-(Cy)\otimes(Cy)\big{)}$
$=-N^{*}+\overline{\lambda}(y\otimes y-x\otimes x)$
$=-\left(N+\lambda(x\otimes x-y\otimes y)\right)^{*}.$
(ii)$\implies$(i). For convenience, put $T=x\otimes x-y\otimes y$. By the fact that the class of $C$-skew symmetric operators is a subspace of ${\mathcal{B}}(H)$, we clearly have $CNC=-N^{*}$ and $CTC=-T^{*}=-T$. Noting that $T$ is skew-symmetric, it must have a skew-symmetric matrix with respect to some orthonormal basis. In particular, since $T$ is a rank-two operator, its trace is well-defined and should be zero. Combining this with the fact that $T$ is self-adjoint, it follows that there exists orthonormal vectors $e_{1},e_{2}\in H$ and $\alpha\in\mathbb{R}$ such that
${\rm ran}(T)={\rm ker}(T-\alpha I)\oplus{\rm ker}(T+\alpha I),$
${\rm ker}(T-\alpha I)=\vee\{e_{1}\}\quad\mbox{and}\quad{\rm ker}(T+\alpha I)=\vee\{e_{2}\}.$
As $CTC=-T$, we get $C(T+\alpha I)C=-(T-\alpha I)$; hence
(3.4) $C(\vee\{e_{1}\})=C{\rm ker}(T-\alpha I)={\rm ker}(T+\alpha I)=\vee\{e_{2}\}.$
Note that $x$ and $y$ have the same norm. In fact, $\langle x,x\rangle-\langle y,y\rangle$ is the trace of $T$. Hence, using [29, Theorem 2.1], we obtain a conjugation $J$ on $H$ satisfying $Jx=y$; one can check that $JTJ=-T$. The argument in (3.4) shows that $J(\vee\{e_{1}\})=\vee\{e_{2}\}$. Then, since $C$ and $J$ are isometries, there exist unimodular scalars $a,b\in\mathbb{C}$ such that $Ce_{1}=ae_{2}$ and $Je_{1}=be_{2}$. Furthermore, since $C^{2}=J^{2}=I$, we get $Ce_{2}=ae_{1}$ and $Je_{2}=be_{1}$. Noting that $C$ and $J$ are both conjugate-linear, we get $Ch=a\overline{b}Jh$ for every $h\in\vee\{e_{1},e_{2}\}$. In particular, since $x,y\in{\rm ran}(T)=\vee\{e_{1},e_{2}\}$, we obtain $Cx=a\overline{b}Jx=a\overline{b}y$. Put $\widetilde{C}=\overline{a}bC$; one can easily check that $\widetilde{C}$ is a conjugation that satisfies $\widetilde{C}N\widetilde{C}=-N^{*}$ and $\widetilde{C}x=y$. This ends the proof of the theorem. ∎
The remainder of this section is devoted to proving Theorems 3.1 and 3.2.
3.2. Proof of Theorem 3.1 and 3.2
First, we begin with the following characterization of skew-symmetric normal operators.
Lemma 3.5.
Let $N\in{\mathcal{B}}(H)$ be a normal operator. The following statements are equivalent:
  1. (i)
    There exists a unitary operator $U$ on $H$ such that $UNU^{*}=-N$.
  2. (ii)
    There exists a conjugation $C$ on $H$ such that $CNC=-N^{*}$.
  3. (iii)
    There exists an anti-unitary operator $V$ on $H$ such that $VNV^{-1}=-N^{*}$.
Proof.
The implication (ii)$\implies$(iii) is obvious.
(iii)$\implies$(i). Let $V$ be an anti-unitary operator on $H$ such that $VNV^{-1}=-N^{*}$. By normality of $N$, there exists a conjugation $J$ on $H$ that satisfies $JNJ=N^{*}$. Hence, we have $VNV^{-1}=-JNJ$, and so, since $J^{2}=I$, we get
$(JV)N(V^{-1}J)=-J^{2}NJ^{2}=-N.$
Put $U=JV$; then $U\in{\mathcal{B}}(H)$ is unitary and $U^{*}=U^{-1}=(JV)^{-1}=V^{-1}J^{-1}=V^{-1}J$. Therefore, $UNU^{*}=-N$.
(i)$\implies$(ii). Let $U\in{\mathcal{B}}(H)$ be a unitary operator satisfying $UN=-NU$. Fix a Borel subset $\Delta_{0}\subset\mathbb{C}$ such that $\mathbb{C}\setminus\Delta_{0}=-\Delta_{0}\cup\{0\}$. By Lemma 2.11, we have $UE_{N}(\Delta_{0})=E_{N}(-\Delta_{0})U$ and $UE_{N}(-\Delta_{0})=E_{N}(\Delta_{0})U$. Consequently,
$U{\rm ran}(E_{N}(\Delta_{0}))={\rm ran}(E_{N}(-\Delta_{0}))\quad\mbox{and}\quad U{\rm ran}(E_{N}(-\Delta_{0}))={\rm ran}(E_{N}(\Delta_{0})).$
Given that $\mathbb{C}=\Delta_{0}\cup(-\Delta)\cup\{0\}$ is a partition, it follows from the above equalities that $U$ has the form
$U=\begin{bmatrix}0&*&0\\ W&0&0\\ 0&0&*\end{bmatrix}\begin{array}[]{l}{\rm ran}(E_{N}(\Delta_{0}))\\ {\rm ran}(E_{N}(-\Delta_{0}))\\ {\rm ran}(E_{N}(\{0\}))\end{array}$
Where $W$ is an isometric isomorphism between Hilbert spaces. Noting that $N$ can also be written as
$N=\begin{bmatrix}N_{1}&0&0\\ 0&N_{2}&0\\ 0&0&0\end{bmatrix}\begin{array}[]{l}{\rm ran}(E_{N}(\Delta_{0}))\\ {\rm ran}(E_{N}(-\Delta_{0}))\\ {\rm ran}(E_{N}(\{0\}))\end{array},$
and since $UN=-NU$, simple matrix calculations show that $WN_{1}=-N_{2}W$. This implies that $N_{1}\cong-N_{2}$, and hence $N_{|{\rm ker}(N)^{\perp}}=N_{1}\oplus N_{2}\cong N_{1}\oplus-N_{1}$. Therefore, $N$ is skew symmetric by [17, Theorem 1.10]. ∎
Note that every bounded linear $T$ acting on a Hilbert space can be written as $T=N\oplus A$ where $N$ is normal and $A$ is an operator that has no reducing subspace for which the restriction of $A$ is normal (of course, either $N$ or $A$ could be absent). In fact, $N$ is the restriction of $T$ to the reducing subspace
(3.5) $H_{\rm nor}=\bigcap_{1\leq n,m<\infty}{\rm ker}({T^{*}}^{n}T^{m}-T^{m}{T^{*}}^{n}).$
The operator $N$ is called the normal part of $T$, while $A$ is called the abnormal part of $T$. In this case, we denote $T_{\rm nor}=N$ and $T_{\rm abnor}=A$.
Using (3.5), one can easily see that if $M$ is a normal operator, then
(3.6) $(M\oplus T)_{\rm nor}=M\oplus T_{\rm nor}\quad\mbox{and}\quad(M\oplus T)_{\rm abnor}=T_{\rm abnor}.$
The following lemma is an immediate consequence of [28, Proposition 3.10] and (3.6).
Lemma 3.6.
Let $M\in{\mathcal{B}}(H)$ be normal and $T\in{\mathcal{B}}(K)$ such that $M\oplus T$ is skew-symmetric. Then, $M\oplus T_{\rm nor}$ and $T_{\rm abnor}$ are skew-symmetric.
Lemma 3.7.
Let $N\in{\mathcal{B}}(H)$ be a skew-symmetric normal operator and $\Delta\subseteq\mathbb{C}$ a Borel subset such $\Delta=-\Delta$. Then, the restriction of $N$ to the reducing subspace ${\rm ran}(E_{N}(\Delta))$ is skew-symmetric.
Proof.
Let $C$ be a conjugation on $H$ such that $CNC=-N^{*}$. By Lemma 2.11, we have $CE_{N}(\Delta)=E_{N}(-\Delta)C=E_{N}(\Delta)C$; hence, $C{\rm ran}(E_{N}(\Delta))={\rm ran}(E_{N}(\Delta))$. Let $C_{0}$ and $N_{0}$ denote the restrictions of $C$ and $N$ to ${\rm ran}(E_{N}(\Delta))$, respectively. Then, $C_{0}$ is a conjugation on $E_{N}(\Delta)$. Moreover, it follows immediately from the equality $CNC=-N^{*}$ that $C_{0}N_{0}C_{0}=-N_{0}^{*}$. ∎
In [16, Lemma 3.2] and [2, Corollary 3.4], the authors proved that if $N$ is normal, then a direct sum $N\oplus T$ is complex symmetric (resp. conjugate-normal) if and only if $T$ is complex-symmetric (resp. conjugate-normal). However, this property does not hold for skew-symmetry even if $N$ is skew-symmetric. The next result provides a complete description of normal operators $N$ that satisfy this property.
Theorem 3.8.
Let $N\in{\mathcal{B}}(H)$ be normal. The followings statements are equivalents:
  1. (i)
    For every separable complex Hilbert space $K$ and every $T\in\mathcal{B}(K)$,
[…]
  2. (ii)
    $N$ is skew-symmetric and $m(N_{|{\rm ker}(N)^{\perp}})<\infty$.
Proof.
(i)$\implies$(ii). Suppose statement (i) holds. Taking $T=0$, the zero operator on any Hilbert space $K$, we obtain that $N\oplus 0$ is skew-symmetric. By [17, Lemma 2.1], it follows that $N_{|({\rm ker}N)^{\perp}}$ is skew-symmetric. Therefore, $N=N_{|({\rm ker}N)^{\perp}}\oplus 0$ is skew-symmetric.
For convenience, put $N_{1}=N_{|{\rm ker}(N)^{\perp}}$, and let $N_{1}\cong M_{\mu_{\infty}}^{(\infty)}\oplus\bigoplus_{1\leq i<\infty}M_{\mu_{i}}^{(i)}$ be the canonical decomposition in (3.1) of $N_{1}$.
We need to show that $\mu_{\infty}=0$. Assume, for the sake of contradiction, that $\mu_{\infty}\neq 0$. Hence, $M_{\mu_{\infty}}$ is not absent and is not zero due to the injectivity of $N_{1}$. Since $\sigma(M_{\mu_{\infty}})\neq\{0\}$, using the spectral theorem, we can write $M_{\mu_{\infty}}=M\oplus L$, were $L$ is a normal operator with $\sigma(L)\neq-\sigma(L)$; in particular, $L\ncong-L$ and thus not skew-symmetric by Lemma 3.5.
Let us verify that $N\oplus L$ is skew symmetric. We have
$N_{1}\oplus L\cong M_{\mu_{\infty}}^{(\infty)}\oplus\bigoplus_{1\leq i<\infty}M_{\mu_{i}}^{(i)}\oplus L$ $\cong(M\oplus L)^{(\infty)}\oplus L\oplus\bigoplus_{1\leq i<\infty}M_{\mu_{i}}^{(i)}$
$\cong M^{(\infty)}\oplus L^{(\infty)}\oplus L\oplus\bigoplus_{1\leq i<\infty}M_{\mu_{i}}^{(i)}$
$\cong M^{(\infty)}\oplus L^{(\infty)}\oplus\bigoplus_{1\leq i<\infty}M_{\mu_{i}}^{(i)}$
$\cong N_{1}.$
Thus, $N\oplus L\cong N$; in particular, $N\oplus L$ is skew-symmetric; a contradiction. Therefore, $m(N_{1})<\infty$.
(ii)$\implies$(i). Since $N$ is skew-symmetric, the reverse implication in (i) is always true. Fix a bounded linear operator $T$, assume that $N\oplus T$ is skew-symmetric, and let us show that $T$ is skew-symmetric.
In view of Lemma 3.6 and [17, Lemma 2.1], we can assume without loss of generality that $T$ is normal and $N=N_{|{\rm ker}(N)^{\perp}}$. Moreover, by (3.1) and (3.3), we may directly assume that
$N=\bigoplus_{1\leq i<\infty}M_{\mu_{i}}^{(i)}\quad\mbox{and}\quad H=\bigoplus_{1\leq i<\infty}L^{2}(\mu_{i})^{(i)}$
where the $\mu_{i}$’s are mutually singular and satisfy $\mu_{i}(\Delta)=\mu_{i}(-\Delta)$ for every Borel subset $\Delta\subseteq\mathbb{C}$. Since the measures $\mu_{i}$’s are mutually singular and satisfy (3.3), it is elementary to see that there exist mutually disjoint Borel subsets $\{\Delta_{i}\}_{i\in\mathbb{Z}^{+}}$ of $\mathbb{C}$ such that for all $i,j\in\mathbb{Z}^{+}$ with $i\neq j$,
(3.7) $\Delta_{i}=-\Delta_{i},\quad\mu_{i}(\Delta)=\mu_{i}(\Delta\cap\Delta_{i}),\quad\mu_{i}(\Delta_{j})=0\;\;\mbox{ for every Borel }\Delta\subseteq\mathbb{C}.$
Let $\Delta_{0}=\mathbb{C}\setminus\cup_{i\in\mathbb{Z}^{+}}\Delta_{i}$; clearly, $\Delta_{0}=-\Delta_{0}$. For convenience, put $N_{i}=M_{\mu_{i}}^{(i)}$, $H_{i}=L^{2}(\mu_{i})^{(i)}$ for $i\in\mathbb{Z}^{+}$, and let $T=T_{0}\oplus\bigoplus_{i\in\mathbb{Z}^{+}}T_{i}$ be the decomposition of $T$ with respect to the orthogonal decomposition $K={\rm ran}(E_{T}(\Delta_{0}))\oplus\bigoplus_{i\in\mathbb{Z}^{+}}{\rm ran}(E_{T}(\Delta_{i}))$. For $0\leq i<\infty$, put $K_{i}={\rm ran}(E_{T}(\Delta_{i}))$.
Since by (3.7), $\mu_{i}(\Delta_{0})=\mu_{i}(\Delta_{j})=0$ all $i,j\in\mathbb{Z}^{+}$ with $i\neq j$, it follows that $E_{N_{i}}(\Delta_{0})=E_{N_{i}}(\Delta_{j})=0$. Therefore,
${\rm ran}(E_{N\oplus T}(\Delta_{0}))=K_{0}\quad\mbox{and}\quad{\rm ran}(E_{N\oplus T}(\Delta_{i}))=H_{i}\oplus K_{i}\;\;\mbox{for every }i\in\mathbb{Z}^{+}.$
It follows that
$(N\oplus T)_{|{\rm ran}(E_{N\oplus T}(\Delta_{0}))}=T_{0}\quad\mbox{and}\quad(N\oplus T)_{|{\rm ran}(E_{N\oplus T}(\Delta_{i}))}=N_{i}\oplus T_{i}\;\;\mbox{for every }i\in\mathbb{Z}^{+}.$
Consequently, by Lemma 3.7, $T_{0}$ and $N_{i}\oplus T_{i}$, for $i\in\mathbb{Z}^{+}$, are skew-symmetric operators. In particular, by Lemma 3.5, we have
(3.8) $N_{i}\oplus T_{i}\cong-N_{i}\oplus-T_{i}\quad\mbox{for every }i\in\mathbb{Z}^{+}.$
Fix $i\in\mathbb{Z}^{+}$, and let us prove that $T_{i}$ is skew-symmetric. Given $\mu_{i}(\Delta)=\mu_{i}(-\Delta)$ for every Borel subset $\Delta\subseteq\mathbb{C}$, [17, Theorem 1.11] implies that $N_{i}$ is skew-symmetric; hence, $N_{i}\cong-N_{i}$ by Lemma 3.5. Combining this with (3.8), we obtain $N_{i}\oplus T_{i}\cong N_{i}\oplus-T_{i}$. Since $M_{\mu_{i}}$ is a star-cyclic normal operator ([8, Page 269, Theorem 3.4]), applying [8, Page 295, Proposition 10.6] $i$ times to $M_{\mu_{i}}$, we conclude that $T_{i}\cong-T_{i}$, and so $T_{i}$ is skew-symmetric by Lemma 3.5.
Thus, $T=T_{0}\oplus\bigoplus_{i\in\mathbb{Z}^{+}}T_{i}$ is skew-symmetric as a direct sum of skew-symmetric operators. ∎
Lemma 3.9.
Let $N\in{\mathcal{B}}(H)$ be normal, and suppose that there exist reducing subspaces $\{H_{i}\}_{1\leq i\leq r}$, $r<\infty$, of $N$ such that $H=\vee(\sum_{i=1}^{r}H_{i})$ and $N_{|H_{i}}$ is star-cyclic for each $i\in\{1,2,\dots,r\}$. Then, for every reducing subspace $L$ of $N$, we have $m(N_{|L})<\infty$.
Proof.
First, let us show that $N$ is a finite direct sum of star-cyclic normal operators. We will proceed by induction on $r$. The result is clear when $r=1$. Assume the lemma holds for some integer $r\geq 1$. Let $H_{1},H_{2},\dots H_{r+1}$ be $r+1$ reducing subspaces of $N$ such that $H=\vee(\sum_{i=1}^{r+1}H_{i})$ and $N_{|H_{i}}$ is star cyclic for $i\in\{1,2,\dots,r+1\}$. Let $K=\vee(\sum_{i=1}^{r}H_{i})$ and $P\in{\mathcal{B}}(H)$ be the orthogonal projection onto $K$. Then, $K$ reduces $N$; moreover, by the induction hypothesis, $N_{|K}$ is a finite direct sum of star-cyclic operators.
By star-cyclicity of $N_{|H_{r+1}}$, there exists $h\in H$ such that $H_{r+1}=L_{N,h}$. To complete the proof, it is clear that it suffices to show that $H=K\oplus L_{N,(I-P)h}$. Note that $K$ and $L_{N,(I-P)h}$ are orthogonal. In fact, since $K$ reduces $N$, its orthogonal projection $P$, and hence $(I-P)$, commutes with $N$ and $N^{*}$. Consequently,
$L_{N,(I-P)h}$ =$=\bigvee\big{\{}p(N,N^{*})(I-P)h:p\in\mathbb{C}[x,y]\big{\}}$
$=(I-P)\bigvee\big{\{}p(N,N^{*})h:p\in\mathbb{C}[x,y]\big{\}}\subseteq K^{\perp}.$
It remains to prove that $H=K+L_{N,(I-P)h}$; we need only to show that $H_{r+1}\subseteq K+L_{N,(I-P)h}$. For $p\in\mathbb{C}[x,y]$, we have $p(N,N^{*})h=p(N,N^{*})Ph+p(N,N^{*})(I-P)h$. Since $Ph\in K$ and $K$ reduces $p(N,N^{*})$, we have $p(N,N^{*})Ph\in K$. Hence, $p(N,N^{*})h\in K+L_{N,(I-P)h}$, and thus $H_{r+1}\subseteq K+L_{N,(I-P)h}$.
Now, let $L$ be a reducing subspace of $N$ and assume, for contradiction, that $m(N_{|L})\nless\infty$. According to (3.1), $N_{|L}$ has a reducing subspace for which its restriction is unitarily equivalent to $M_{\mu}^{(\infty)}$ for some finite non-zero compactly supported measure $\mu$ on $\mathbb{C}$. Thus, we can write $N\cong N_{0}\oplus M_{\mu}^{(\infty)}$ for some normal operator $N_{0}$. This implies
$N\oplus M_{\mu}^{(2)}\cong N_{0}\oplus M_{\mu}^{(\infty)}\oplus M_{\mu}^{(2)}\cong N_{0}\oplus M_{\mu}^{(\infty)}\oplus M_{\mu}\cong N\oplus M_{\mu}.$
Since $N$ is a direct sum of $n$ star-cyclic normal operators $M_{1},M_{2},\dots,M_{n}$ with $n<\infty$, applying [8, Page 295, Proposition 10.6] to each $M_{i}$, we conclude $M_{\mu}\cong M_{\mu}^{(2)}$; which contradicts (3.2). ∎
We are now ready to prove Theorems 3.1 and 3.2.
Proof of Theorem 3.1.
The implication (i)$\implies$(ii) is clear.
(ii)$\implies$(iii). First, let us show that $N_{|L_{N,Z}^{\perp}}$ is skew-symmetric. By the argument in (2.7), there exists a conjugation $J$ on $H$ so that $(JV)N=-N(JV)$ and $JN^{*}J=N$. By the Fuglede-Putnam Theorem, it follows that $(JV)N^{*}=-N^{*}(JV)$; hence, $VN^{*}=J^{2}VN^{*}=-JN^{*}JV=-NV$. Consequently, for any polynomial $p\in\mathbb{C}[x,y]$, one can easily see that $Vp(N,N^{*})=p(-N,-N^{*})^{*}V$. Therefore, for every $i\in I$,
$Vp(N,N^{*})x_{i}=p(-N,-N^{*})^{*}Vx_{i}=p(-N,-N^{*})^{*}y_{i}$
and
$Vp(N,N^{*})y_{i}=p(-N,-N^{*})^{*}Vy_{i}=p(-N,-N^{*})^{*}x_{i}.$
This implies that if $L_{0}={\rm span}\big{\{}p(N,N^{*})x_{i}+q(N,N^{*})y_{i}:p,q\in\mathbb{C}[x,y],i\in I\big{\}}$, then $VL_{0}=L_{0}$. Taking the closure of $L_{0}$ in $H$, we obtain $VL_{N,Z}=L_{N,Z}$; moreover, since $V$ is anti-unitary, we get $VL_{N,Z}^{\perp}=L_{N,Z}^{\perp}$.
Let $V_{1}$ and $N_{1}$ denote the restrictions of $V$ and $N$ to $L_{N,Z}^{\perp}$, respectively. Then, $V_{1}$ is anti-unitary on $L_{N,Z}^{\perp}$, and the relation $VN=-N^{*}V$ implies $V_{1}N_{1}=-N_{1}^{*}V_{1}$. Consequently, $N_{1}$ is skew-symmetric by Lemma 3.5.
Now, by Lemma 2.11, for every Borel subset $\Delta\subseteq\mathbb{C}$, we have $VE_{N}(-\Delta)=E_{N}(\Delta)V$. Hence, using (2.9), for all $i,j\in I$,
$\langle E_{N}(\Delta)x_{i},y_{j}\rangle$ $=\langle E_{N}(\Delta)Vy_{i},Vx_{j}\rangle=\langle VE_{N}(-\Delta)y_{i},Vx_{j}\rangle$
$=\langle V^{-1}Vx_{j},E_{N}(-\Delta)y_{i}\rangle=\langle E_{N}(-\Delta)x_{j},y_{i}\rangle,$
and similarly, $\langle E_{N}(\Delta)x_{i},x_{j}\rangle=\langle E_{N}(-\Delta)y_{j},y_{i}\rangle$.
(iii)$\implies$(i). Let $M=-N$. For every Borel subset $\Delta\subseteq\mathbb{C}$ and for all $i,j\in I$,
$\langle E_{N}(\Delta)x_{i},y_{j}\rangle=\langle E_{N}(-\Delta)x_{j},y_{i}\rangle=\langle E_{-N}(\Delta)x_{j},y_{i}\rangle=\langle E_{M}(\Delta)x_{j},y_{i}\rangle,$
and similarly, $\langle E_{N}(\Delta)x_{i},x_{j}\rangle=\langle E_{M}(\Delta)y_{j},y_{i}\rangle$. For convenience, put $H_{0}=L_{N,Z}$, $H_{1}=H_{0}^{\perp}$, $N_{0}=N_{|H_{0}}$ and $N_{1}=N_{|H_{1}}$. By Theorem 2.10, there exists a conjugation $C_{0}$ on $H_{0}$ such that $C_{0}N_{0}C_{0}=-N_{0}^{*}$ and $C_{0}x_{i}=y_{i}$ for every $i\in I$. Since $N_{1}$ is skew-symmetric, there exists a conjugation $C_{1}$ on $H_{1}$ such that $C_{1}N_{1}C_{1}=-N_{1}^{*}$, and thus the conjugation $C=C_{0}\oplus C_{1}$ satisfies all required properties. ∎
Proof of Theorem 3.2.
The implication (i)$\implies$(ii) follows directly from Theorem 3.1.
(ii)$\implies$(i). Let $N_{0}$ and $N_{1}$ be the normal operators in the proof of Theorem 3.2. In the light of the same theorem, we only need to show that $N_{1}$ is skew-symmetric. This follows immediately from Theorem 3.8 and the facts that $m\left({N_{0}}_{|{\rm ker}(N_{0})^{\perp}}\right)<\infty$ and $N=N_{0}\oplus N_{1}$ is skew-symmetric.
Let us now prove the second part of the theorem. Assume that $I$ is finite. Since
$H_{0}=\bigvee\left(\sum_{i\in I}L_{N,\{x_{i}\}}+L_{N,\{y_{i}\}}\right)=\bigvee\left(\sum_{i\in I}L_{N_{0},\{x_{i}\}}+L_{N_{0},\{y_{i}\}}\right)$
and the restrictions of $N_{0}$ to $L_{N_{0},\{x_{i}\}}$ and $L_{N_{0},\{y_{i}\}}$ are star-cyclic, Lemma 3.9 implies that $m({N_{0}}_{|{\rm ker}(N_{0})^{\perp}})<\infty$. This concludes the proof. ∎
4. Applications to some equations in $L^{\infty}(\mu)$
Throughout this section, let $\mu$ denote a finite measure on Borel subsets of $\mathbb{C}$ with compact support, and let $n$ be a positive integer. All functions are considered defined on $\mathbb{C}$. For a given function $f$, the notation $\widetilde{f}$ stands for the function defined by $\widetilde{f}(z)=\overline{f(-z)}$ for $z\in\mathbb{C}$.
Let $[\phi_{i,j}]_{1\leq i,j\leq n}$ be a matrix of functions in $L^{\infty}(\mu)$. We denote $[\phi_{i,j}]_{1\leq i,j\leq n}\in\mathcal{U}(\mu,n)$ if the matrix $[\phi_{i,j}(z)]_{1\leq i,j\leq n}$ is unitary for $\mu$-almost every $z\in\mathbb{C}$. If, in addition, the functions $\phi_{i,j}$ satisfy the condition $\phi_{i,j}(z)=\phi_{j,i}(-z)$ for $\mu$-almost every $z\in\mathbb{C}$, then we write $[\phi_{i,j}]_{1\leq i,j\leq n}\in\mathcal{SU}(\mu,n)$.
For a function $\phi\in L^{\infty}(\mu)$, $M_{\phi}$ denotes the multiplication operator by $\phi$ on $L^{2}(\mu)$; that is, $[M_{\phi}f](z)=\phi(z)f(z)$ for every $f\in L^{2}(\mu)$.
Let $\{f_{i}\}_{1\leq i\leq n}$ and $\{g_{i}\}_{1\leq i\leq n}$ be functions in $L^{2}(\mu)$ so that $\sum_{k=1}^{n}|f_{k}|^{2}=\sum_{k=1}^{n}|g_{k}|^{2}$ $\mu$-almost everywhere ($\mu$-a.e.). It is clear that we can find functions $\phi_{i,j}$ so that, for $\mu$-almost every $z\in\mathbb{C}$, the matrix $[\phi_{i,j}(z)]_{1\leq i,j\leq n}$ is unitary and
$\left[\phi_{i,j}(z)\right]_{1\leq i,j\leq n}\begin{bmatrix}f_{1}(z),f_{2}(z),\cdots,f_{n}(z)\end{bmatrix}^{\rm tr}=\begin{bmatrix}g_{1}(z),g_{2}(z),\cdots,g_{n}(z)\end{bmatrix}^{\rm tr}.$
However, it is not clear whether the functions $\phi_{i,j}$ can be chosen to be Borel measurable. The next theorem provides a positive answer.
Theorem 4.1.
Let $\{f_{i}\}_{1\leq i\leq n}$ and $\{g_{i}\}_{1\leq i\leq n}$ be two families of functions in $L^{2}(\mu)$. Then, the following equation
[…]
has a solution $[\phi_{i,j}]_{1\leq i,j\leq n}\in\mathcal{U}(\mu,n)$ if and only if $\sum_{k=1}^{n}|f_{k}|^{2}=\sum_{k=1}^{n}|g_{k}|^{2}$ $\mu$-a.e.
Lemma 4.2.
Let $U\in\mathcal{B}\left(L^{2}(\mu)^{(n)}\right)$. Then, $U$ is unitary and commutes with $M_{\mu}^{(n)}$ if and only if $U=\left[M_{\phi_{i,j}}\right]_{1\leq i,j\leq n}$ for some $[\phi_{i,j}]_{1\leq i,j\leq n}\in\mathcal{U}(\mu,n)$.
Proof.
Suppose $U=\left[M_{\phi_{i,j}}\right]_{1\leq i,j\leq n}$ for some $[\phi_{i,j}]_{1\leq i,j\leq n}\subset L^{\infty}(\mu)$. Then,
$UU^{*}$ $=\left[\sum_{k=1}^{n}M_{\phi_{i,k}}M_{\phi_{j,k}}^{*}\right]_{1\leq i,j\leq n}=\left[\sum_{k=1}^{n}M_{\phi_{i,k}}M_{\overline{\phi_{j,k}}}\right]_{1\leq i,j\leq n}$
$=\left[\sum_{k=1}^{n}M_{\phi_{i,k}\overline{\phi_{j,k}}}\right]_{1\leq i,j\leq n}=\left[M_{\sum_{k=1}^{n}\phi_{i,k}\overline{\phi_{j,k}}}\right]_{1\leq i,j\leq n}.$
Similarly, $U^{*}U=\left[M_{\sum_{k=1}^{n}\overline{\phi_{k,i}}\phi_{k,j}}\right]_{1\leq i,j\leq n}$. Hence, the operator $U$ is unitary if and only if, for all $1\leq i,j\leq n$,
[…]
where $\delta_{ij}$ denotes the Kronecker delta. That is, $[\phi_{i,j}]_{1\leq i,j\leq n}\in\mathcal{U}(\mu,n)$. The desired equivalence follows now directly from [8, Page 276, Proposition 6.1 (b)] and [8, Page 279, Corollary 6.9]. ∎
Proof of Theorem 4.1.
In the light of the previous lemma, the equation (4.1) has a solution in $\mathcal{U}(\mu,n)$ if and only if there exists a unitary operator $U$ commuting with $M_{\mu}^{(n)}$ such that $UF=G$, where
$F=\begin{bmatrix}f_{1},f_{2},\cdots,f_{n}\end{bmatrix}^{\rm tr}\quad\mbox{and}\quad G=\begin{bmatrix}g_{1},g_{2},\cdots,g_{n}\end{bmatrix}^{\rm tr}.$
By Theorem 2.4, this is true if and only if $\left\|E_{M_{\mu}^{(n)}}(\Delta)F\right\|_{L^{2}(\mu)^{(n)}}=\left\|E_{M_{\mu}^{(n)}}(\Delta)G\right\|_{L^{2}(\mu)^{(n)}}$ for every Borel subset $\Delta\subseteq\mathbb{C}$. Note that
$E_{M_{\mu}^{(n)}}(\Delta)=\bigoplus_{k=1}^{n}E_{M_{\mu}}(\Delta)=\bigoplus_{k=1}^{n}M_{\chi_{\Delta}}$
(see, for instance, [9, Page 49, Proposition 10.4 (a)]). Consequently, (4.1) has a solution in $\mathcal{U}(\mu,n)$ if and only if, for every Borel subset $\Delta\subseteq\mathbb{C}$, we have
$\int\sum_{k=1}^{n}|f_{k}|^{2}\chi_{\Delta}\,{\rm d}\mu$ $=\sum_{k=1}^{n}\int|f_{k}|^{2}\chi_{\Delta}\,{\rm d}\mu=\sum_{k=1}^{n}\left\|E_{M_{\mu}}(\Delta)f_{k}\right\|_{L^{2}(\mu)}^{2}=\left\|E_{M_{\mu}^{(n)}}(\Delta)F\right\|_{L^{2}(\mu)^{(n)}}^{2}$
$=\left\|E_{M_{\mu}^{(n)}}(\Delta)G\right\|_{L^{2}(\mu)^{(n)}}^{2}=\int\sum_{k=1}^{n}|g_{k}|^{2}\chi_{\Delta}\,{\rm d}\mu.$
That is, $\sum_{k=1}^{n}|f_{k}|^{2}=\sum_{k=1}^{n}|g_{k}|^{2}$ $\mu$-a.e. ∎
The measure $\mu$ is called a symmetric measure if $\mu(\Delta)=\mu(-\Delta)$ for every Borel subset $\Delta\subseteq\mathbb{C}$.
Theorem 4.3.
Assume that $\mu$ is symmetric, and let $\{f_{i}\}_{1\leq i\leq n}$ and $\{g_{i}\}_{1\leq i\leq n}$ be two families of functions in $L^{2}(\mu)$. Then, the following equation
[…]
has a solution $[\phi_{i,j}]_{1\leq i,j\leq n}\in\mathcal{SU}(\mu,n)$ if and only if $\sum_{k=1}^{n}|f_{k}(z)|^{2}=\sum_{k=1}^{n}|g_{k}(z)|^{2}$ and $\sum_{k=1}^{n}f_{k}(z)g_{k}(-z)=\sum_{k=1}^{n}f_{k}(-z)g_{k}(z)$ for $\mu$-almost every $z\in\mathbb{C}$.
Lemma 4.4.
Let $U\in{\mathcal{B}}(H)$ be a unitary operator and $J$ a conjugation on $H$. Then, $U=JC$ for some conjugation $C$ on $H$ if and only if $JUJ=U^{*}$.
Proof.
In the light of the facts that $J^{2}=I$ and $\langle JUh,JUk\rangle=\langle Uk,Uh\rangle=\langle k,h\rangle$ for all $h,k\in H$, it is easy to see that
[…]
$\iff(JU)=(JU)^{-1}=U^{*}J^{-1}$
$\iff JUJ=U^{*}.$
Proof of Theorem 4.3.
First, note that the map $J$ defined by $Jf(z)=\overline{f(-z)}$ is a conjugation on $L^{2}(\mu)$. In fact, $J$ is a well-defined conjugate-linear map with $J^{2}=I_{L^{2}(\mu)}$. Additionally, since $\mu$ is symmetric, we have
$\langle Jf,Jg\rangle_{L^{2}(\mu)}=\int\overline{f(-z)}g(-z)\,{\rm d}\mu=\int\overline{f(z)}g(z)\,{\rm d}\mu=\langle g,f\rangle_{L^{2}(\mu)}$
for all $f,g\in L^{2}(\mu)$.
Also, for every $\phi\in L^{\infty}(\mu)$, we have
$JM_{\phi}Jf(z)=JM_{\phi}\overline{f(-z)}=J\phi(z)\overline{f(-z)}=\overline{\phi(-z)}f(z)$
for every $f\in L^{2}(\mu)$. Hence,
(4.3) $JM_{\phi}J=M_{\widetilde{\phi}}.$
In particular, $J^{(n)}M_{\mu}^{(n)}J^{(n)}=-M_{\mu}^{(n)*}$, and $M_{\mu}^{(n)}$ is skew-symmetric since $J^{(n)}$ is a conjugation on $L^{2}(\mu)^{(n)}$.
Claim: An operator $U\in\mathcal{B}\left(L^{2}(\mu)^{(n)}\right)$ is unitary, commutes with $M_{\mu}^{(n)}$ and satisfies $J^{(n)}UJ^{(n)}=U^{*}$ if and only if $U=\left[M_{\phi_{i,j}}\right]_{1\leq i,j\leq n}$ for some $[\phi_{i,j}]_{1\leq i,j\leq n}\in\mathcal{SU}(\mu,n)$. By Lemma 4.2, it suffices to show that if $U=\left[M_{\phi_{i,j}}\right]_{1\leq i,j\leq n}$, then
$J^{(n)}UJ^{(n)}=U^{*}\quad\Longleftrightarrow\quad\phi_{i,j}(z)=\phi_{j,i}(-z)\;\mu\mbox{-a.e. for all }1\leq i,j\leq n.$
In fact, simple matrix calculations show that
$J^{(n)}UJ^{(n)}=U^{*}$ $\quad\Longleftrightarrow\quad\left(JM_{\phi_{i,j}}J\right)_{1\leq i,j\leq n}=\left(M_{\phi_{j,i}}^{*}\right)_{1\leq i,j\leq n}$
$\quad\Longleftrightarrow\quad\left(JM_{\phi_{i,j}}J\right)_{1\leq i,j\leq n}=\left(M_{\overline{\phi_{j,i}}}\right)_{1\leq i,j\leq n}$
[…]
$\quad\Longleftrightarrow\quad\phi_{i,j}(z)=\phi_{j,i}(-z)\;\;\mu\mbox{-a.e. for all }1\leq i,j\leq n.$
This proves the claim.
Now, one can see that the equation (4.2) has a solution in $\mathcal{SU}(\mu,n)$ if and only if, there exists a unitary operator $U$ on $L^{2}(\mu)^{(n)}$ commuting with $M_{\mu}^{(n)}$ such that $J^{(n)}UJ^{(n)}=U^{*}$ and
$U\begin{bmatrix}f_{1},f_{2},\cdots,f_{n}\end{bmatrix}^{\rm tr}=\begin{bmatrix}g_{1},g_{2},\cdots,g_{n}\end{bmatrix}^{\rm tr}.$
By Lemma 4.4, this true if and only if there exists a conjugation $C$ on $L^{2}(\mu)^{(n)}$ such that $J^{(n)}C$ commutes with $M_{\mu}^{(n)}$ and
(4.4) $J^{(n)}C\begin{bmatrix}f_{1},f_{2},\cdots,f_{n}\end{bmatrix}^{\rm tr}=\begin{bmatrix}g_{1},g_{2},\cdots,g_{n}\end{bmatrix}^{\rm tr}.$
Note that
$J^{(n)}CM_{\mu}^{(n)}=M_{\mu}^{(n)}J^{(n)}C$ $\quad\Longleftrightarrow\quad CM_{\mu}^{(n)}C=J^{(n)}M_{\mu}^{(n)}J^{(n)}$
[…]
Hence, taking into account (4.4), the equation (4.2) has a solution in $\mathcal{SU}(\mu,n)$ if and only if there exists a conjugation $C$ on $L^{2}(\mu)^{(n)}$ such that $CM_{\mu}^{(n)}C=-M_{\mu}^{(n)*}$ and
$C\begin{bmatrix}f_{1},f_{2},\cdots,f_{n}\end{bmatrix}^{\rm tr}=J^{(n)}\begin{bmatrix}g_{1},g_{2},\cdots,g_{n}\end{bmatrix}^{\rm tr}=\begin{bmatrix}\widetilde{g_{1}},\widetilde{g_{2}},\cdots,\widetilde{g_{n}}\end{bmatrix}^{\rm tr}.$
For convenience, put $F=\begin{bmatrix}f_{1},f_{2},\cdots,f_{n}\end{bmatrix}^{\rm tr}$ and $\widetilde{G}=\begin{bmatrix}\widetilde{g_{1}},\widetilde{g_{2}},\cdots,\widetilde{g_{n}}\end{bmatrix}^{\rm tr}$. Then, by Theorem 3.2, the equation (4.2) has a solution in $\mathcal{SU}(\mu,n)$ if and only if
$\left\langle E_{M_{\mu}^{(n)}}(\Delta)F,\widetilde{G}\right\rangle_{L^{2}(\mu)^{(n)}}=\left\langle E_{M_{\mu}^{(n)}}(-\Delta)F,\widetilde{G}\right\rangle_{L^{2}(\mu)^{(n)}}$
and
$\left\langle E_{M_{\mu}^{(n)}}(\Delta)F,F\right\rangle_{L^{2}(\mu)^{(n)}}=\left\langle E_{M_{\mu}^{(n)}}(-\Delta)\widetilde{G},\widetilde{G}\right\rangle_{L^{2}(\mu)^{(n)}}$
for every Borel subset $\Delta\subseteq\mathbb{C}$. That is,
$\left\langle\left(\bigoplus_{i=1}^{n}M_{\chi_{\Delta}}\right)F,\widetilde{G}\right\rangle_{L^{2}(\mu)^{(n)}}=\left\langle\left(\bigoplus_{i=1}^{n}M_{\chi_{-\Delta}}\right)F,\widetilde{G}\right\rangle_{L^{2}(\mu)^{(n)}}$
and
$\left\langle\left(\bigoplus_{i=1}^{n}M_{\chi_{\Delta}}\right)F,F\right\rangle_{L^{2}(\mu)^{(n)}}=\left\langle\left(\bigoplus_{i=1}^{n}M_{\chi_{-\Delta}}\right)\widetilde{G},\widetilde{G}\right\rangle_{L^{2}(\mu)^{(n)}}$
for every Borel subset $\Delta\subseteq\mathbb{C}$; which means that
$\int\sum_{i=1}^{n}f_{i}(z)g_{i}(-z)\chi_{\Delta}\,{\rm d}\mu=\int\sum_{i=1}^{n}f_{i}(z)g_{i}(-z)\chi_{-\Delta}\,{\rm d}\mu$
and
$\int\sum_{i=1}^{n}|f_{i}(z)|^{2}\chi_{\Delta}\,{\rm d}\mu=\int\sum_{i=1}^{n}|g_{i}(-z)|^{2}\chi_{-\Delta}\,{\rm d}\mu$
for every Borel subset $\Delta\subseteq\mathbb{C}$. Therefore, by symmetry of $\mu$, for every Borel subset $\Delta\subseteq\mathbb{C}$,
$\int\sum_{i=1}^{n}f_{i}(z)g_{i}(-z)\chi_{\Delta}\,{\rm d}\mu=\int\sum_{i=1}^{n}f_{i}(-z)g_{i}(z)\chi_{\Delta}\,{\rm d}\mu$
and
$\int\sum_{i=1}^{n}|f_{i}(z)|^{2}\chi_{\Delta}\,{\rm d}\mu=\int\sum_{i=1}^{n}|g_{i}(z)|^{2}\chi_{\Delta}\,{\rm d}\mu;$
which lead to the desired conditions. ∎

### v3

@@
 Abstract.
-Let $H$ be a separable complex Hilbert space.
-A conjugate-linear map $C:H\to H$ is a conjugation if it is an involutive isometry.
-In this paper, we consider the following interpolation problems: Let $\{x_{i}\}_{i\in I}$ and $\{y_{i}\}_{i\in I}$ be two orthogonal sets of vectors in $H$, and let $N$ and $\{N_{k}\}_{k\in K}$ be normal operators such that the $N_{k}$’s mutually commute.
-Then, under which conditions does there exist a conjugation $C$ on $H$ such that
+Let $\mathcal{H}$ be a separable complex Hilbert space.
+A conjugate-linear map $C:\mathcal{H}\to\mathcal{H}$ is called a conjugation if it is an involutive isometry.
+In this paper, we focus on the following interpolation problems: Let $\{x_{i}\}_{i\in I}$ and $\{y_{i}\}_{i\in I}$ be orthonormal sets of vectors in $\mathcal{H}$, and let $\{N_{k}\}_{k\in K}$ be a set of mutually commuting normal operators.
+We seek to determine under which conditions there exists a conjugation $C$ on $\mathcal{H}$ such that
   1. (a)
@@
   2. (b)
-    $Cx_{i}=y_{i}$, for every $i\in I$, and $CNC=-N^{*}$.
-We provide complete answers to problems (a) and (b). As a consequence of our results, we give necessary and sufficient conditions for the existence of solutions of some equations in $L^{\infty}(\mu)$.
+    $Cx_{i}=y_{i}$ and $CN_{k}C=-N_{k}^{*}$ for all $i\in I$ and $k\in K$.
+We provide complete answers to problems (a) and (b) using the spectral projections of normal operators.
+Our results are then applied to the study of complex symmetric and skew symmetric operators, as well as to the characterization of hyperinvariant subspaces of normal operators through conjugations.
 Key words and phrases:
@@
 2020 Mathematics Subject Classification:
-47B15, 47B38, 47A15.
+47B15, 47A15.
 1. Introduction
-Throughout this paper, $H$ and $K$ are separable complex Hilbert spaces.
+Throughout this paper, $\mathcal{H}$ and $\mathcal{K}$ are separable complex Hilbert spaces.
 We use the notation .,$\langle$.,$\rangle$ .
 , .
-⟩ for the inner product of both $H$ and $K$.
-The algebra of all bounded linear operators on $H$ is denoted by ${\mathcal{B}}(H)$, and the identity operator on $H$ is denoted by $I_{H}$; if there is no ambiguity, we simply write $I$.
+⟩ for the inner product of both $\mathcal{H}$ and $\mathcal{K}$.
+The algebra of all bounded linear operators on $\mathcal{H}$ is denoted by ${\mathcal{B}}(\mathcal{H})$, and the identity operator on $\mathcal{H}$ is denoted by $I_{\mathcal{H}}$.
 The range and the null space of an operator $T$ are denoted by ${\rm ran}(T)$ and ${\rm ker}(T)$, respectively.
 All direct sums in this paper are orthogonal.
-A conjugation $C$ on $H$ is a conjugate-linear map $C:H\to H$ that satisfies
+The sets of non-negative integers and positive integers are denoted, as customary, by $\mathbb{N}$ and $\mathbb{Z}^{+}$, respectively.
+Throughout the paper, $\mathcal{N}=\{N_{k}\}_{k\in K}$ and =$\mathcal{M}=\{M_{k}\}_{k\in K}$ are two sets of normal operators on $\mathcal{H}$ such that each two operators in their union are commuting, and $X=\{x_{i}\}_{i\in I}$ and $Y=\{y_{i}\}_{i\in I}$ will always denote two orthonormal sets of non-zero vectors in $\mathcal{H}$.
+Definition 1.1.
+A conjugation $C$ on $\mathcal{H}$ is a conjugate-linear map $C:\mathcal{H}\to\mathcal{H}$ that satisfies
   1. (i)
-    $C^{2}=I$;
+    $C^{2}=I_{\mathcal{H}}$;
   2. (ii)
-    $\langle Ch,Ck\rangle=\langle k,h\rangle$ for all $h,k\in H$.
-The simplest example of a conjugation on $H$ is the map given by $C\left(\sum_{i}\lambda_{i}e_{i}\right)=\overline{\lambda_{i}}e_{i}$, where $\{e_{i}\}_{i}$ is an orthonormal basis of $H$.
-In fact, according to [13, Lemma 1], any conjugation can be represented in this manner with respect to some orthonormal basis.
+    $\langle Ch,Ck\rangle=\langle k,h\rangle$ for all $h,k\in\mathcal{H}$.
+The simplest example of a conjugation on $\mathcal{H}$ is the map given by $C\left(\sum_{i}\lambda_{i}e_{i}\right)=\overline{\lambda_{i}}e_{i}$ where $\{e_{i}\}_{i}$ is an orthonormal basis of $\mathcal{H}$.
+In fact, according to [14, Lemma 1], any conjugation can be represented in this manner with respect to some orthonormal basis.
 Conjugations have been the subject of intensive study in recent years [6, 7, 12, 13, 21, 22], with origins in physics.
@@
 Recent investigations into conjugations have largely stemmed from their connections to complex-symmetric and skew-symmetric operators.
-An operator $T\in{\mathcal{B}}(H)$ is called complex-symmetric (resp.
-skew-symmetric) if there exists a conjugation $C$ on $H$ such that $CTC=T^{*}$ (resp.
+An operator $T\in{\mathcal{B}}(\mathcal{H})$ is called complex-symmetric (resp.
+skew-symmetric) if there exists a conjugation $C$ on $\mathcal{H}$ such that $CTC=T^{*}$ (resp.
 $CTC=-T^{*}$); in this case, we say more precisely that $T$ is $C$-symmetric (resp.
@@
 $M^{\rm tr}=-M$) where ${\rm tr}$ stands for the transpose ([15, 18]).
+Denote by ${\rm CSO}$ and ${\rm SSO}$ the classes of complex symmetric and skew symmetric operators on $\mathcal{H}$, respectively.
 These classes of linear operators have received significant attention in the last two decades.
 Applications of complex symmetric operators and skew-symmetric matrices are found in various areas (see, for instance, [14, 15, 16, 25] and [8, 23, 24]).
+Given two vectors $x,y\in\mathcal{H}$, we denote by $x\otimes y$ the operator defined by $(x\otimes y)h=\langle h,y\rangle x$ for every $h\in\mathcal{H}$.
+It is easy to see that if $C$ is a conjugation, then
+(1.1) $C(x\otimes y)C=(Cx)\otimes(Cy)\quad\mbox{for all }x,y\in\mathcal{H}.$
 In [19, 20, 27, 30], the authors investigated various interpolation problems for conjugations.
-In [18], Liu et al.
-established that for two orthogonal projections $P$ and $Q$ in ${\mathcal{B}}(H)$, a conjugation $C$ on $H$ exists with $CPC=Q$ if and only if ${\rm ran}(P)$ and ${\rm ran}(Q)$ are in symmetric position; i.e.,
-$\dim\left({\rm ran}(P)^{\perp}\cap{\rm ran}(Q)\right)=\dim\left({\rm ran}(P)\cap{\rm ran}(Q)^{\perp}\right).$
-In [26], Wang et al.
-proved that for two commuting self-adjoint operators $P,Q\in{\mathcal{B}}(H)$, there exists a conjugation $C$ on $H$ satisfying $CPC=Q$ if and only if there exists an isometric isomorphism $U$ between Hilbert spaces such that
-$UPU^{-1}=B_{1}\oplus B_{2}\oplus A\quad\mbox{and}\quad UQU^{-1}=B_{2}\oplus B_{1}\oplus A$
-where $B_{1}$, $B_{2}$ and $A$ are self-adjoint operators such that $B_{1}$ and $B_{2}$ have the same underlying Hilbert space and commute.
-It is well-known that for every normal operator $N$, there exists a conjugation $C$ such that $CNC=N^{*}$ ([13]).
-However, this is not always the case if it is also required that $Cx_{i}=y_{i}$ for some families of vectors $\{x_{i}\}_{i\in I}$ and $\{y_{i}\}_{i\in I}$.
-It is the purpose of this paper to investigate such a problem.
-The paper is structured as follows.
-In Section 2, we present our primary results regarding the interpolation problem: when does there exist a conjugation $C$ on $H$ satisfying $Cx_{i}=y_{i}$ and $CN_{k}C=N_{k}^{*}$ for all $i\in I$ and $k\in K$.
-Here, $\{x_{i}\}_{i\in I}$ and $\{y_{i}\}_{i\in I}$ are orthogonal subsets of $H$, and $\{N_{k}\}_{k\in K}$ is a collection of mutually commuting normal operators on $H$.
-We establish geometric conditions to answer this question, along with several equivalent conditions.
-As a consequence of these results, we extend a recent result by Mashreghi et al.
-[20] concerning hyperinvariant subspaces of unitary operators (see Proposition 2.5).
-In Section 3, we investigate a “skew-symmetric” version of the earlier interpolation problem.
-We aim to identify the conditions under which there exists a conjugation $C$ on $H$ that can fulfill the conditions $Cx_{i}=y_{i}$ and $CNC=-N^{*}$ for all $i\in I$, with $N$ being a normal operator.
-The simplest answer is obtained when $I$ is finite.
-In the last section, we apply our main results to give necessary and sufficient conditions for the existence of solutions of certain equations in $L^{\infty}(\mu)$ with coefficients in $L^{2}(\mu)$.
+For instance, in [19], Liu et al.
+established that for two orthogonal projections $P$ and $Q$ in ${\mathcal{B}}(\mathcal{H})$, a conjugation $C$ on $\mathcal{H}$ exists with $CPC=Q$ if and only if ${\rm ran}(P)$ and ${\rm ran}(Q)$ are in symmetric position; i.e.,
+$\dim\big{(}{\rm ran}(P)^{\perp}\cap{\rm ran}(Q)\big{)}=\dim\big{(}{\rm ran}(P)\cap{\rm ran}(Q)^{\perp}\big{)}.$
+The main motivation for our results in Section 2 stems from the study of complex symmetric operators.
+It should be mentioned that the class ${\rm CSO}$ is not stable under the addition.
+Indeed, it is well-known that every normal operator $N\in{\mathcal{B}}(\mathcal{H})$ is complex symmetric.
+However, even its sum with the simplest complex symmetric operator may fail to be in ${\rm CSO}$ (see [1, Theorem 2.4]).
+The main results in Section 2 can be used to study complex symmetry of $N+T$ for a dense class of operators $T$ in ${\mathcal{B}}(\mathcal{H})$.
+To elaborate, Zhu and Li proved in [30] that operators $T\in{\mathcal{B}}(\mathcal{H})$ of the form
+(1.2) $T=\sum_{j\in J}r_{j}e_{j}\otimes f_{j},$
+where $\{r_{j}\}_{j\in J}$ are distinct positive scalars and $\{e_{j}\}_{j\in J}$ and $\{f_{j}\}_{j\in J}$ are orthonormal subsets of $\mathcal{H}$, form a dense class in ${\mathcal{B}}(\mathcal{H})$.
+Moreover, they showed that such an operator is complex symmetric if and only if there exist unimodular scalars $\lambda_{j}\in\mathbb{C}$, $j\in J$, such that
+(1.3) $\lambda_{i}\langle e_{i},f_{j}\rangle=\lambda_{j}\langle e_{j},f_{i}\rangle\quad\mbox{for all }i,j\in J.$
+Corollary 2.7 provides geometric conditions for the existence of a conjugation $C$ such that $N_{k}+\lambda T$ is $C$-symmetric for all $k\in K$ and $\lambda\in\mathbb{C}$.
+When $N_{k}=0$ for every $k\in K$, we recover the result in (1.3) with ease.
+Furthermore, our results extend a recent finding by Mashreghi et al.
+[21] on hyperinvariant subspaces of unitary operators (see Proposition 2.10).
+In Section 3, we establish necessary and sufficient conditions for the existence of a conjugation $C$ such that $CN_{k}C=-N_{k}^{*}$ and $Cx_{i}=y_{i}$ for all $k\in K$ and $i\in I$.
+These results are used to characterize skew symmetry of operators of the form
+(1.4) $N+\sum_{j\in J}r_{j}\left(e_{j}\otimes e_{j}-f_{j}\otimes f_{j}\right),$
+where $N\in{\mathcal{B}}(\mathcal{H})$ is normal, $\{r_{j}\}_{j\in J}$ are distinct positive scalars and $\{e_{j},f_{j}\}_{j\in J}$ is an orthonormal subset.
+Notably, skew symmetric operators of the form (1.4) are dense in ${\rm SSO}$ (see Section 3).
+To prove the main results of Section 3, we will uncover certain facts about skew symmetric normal operators.
+We will show that $\{N_{k}\}_{k\in K}$ are simultaneously skew-symmetric (i.e., there exists a conjugation $C$ such that $CN_{k}C=-N_{k}^{*}$ for every $k\in K$) if and only if there exists a unitary operator $U\in{\mathcal{B}}(\mathcal{H})$ such that $UN_{k}U^{*}=-N_{k}$ for every $k\in K$.
+Furthermore, for a normal operator $N\in{\mathcal{B}}(\mathcal{H})$, we provide a condition on the multiplicity of $N$ under which skew symmetry of $N\oplus T$ implies that one of $T$, where $T$ is an arbitrary operator.
 2. Interpolation theorems for conjugations : $C$-symmetric versions
-Recall that a linear (or conjugate-linear) operator $V$ on $H$ is called partial isometry if the restriction of $V$ to ${\rm ker}(V)^{\perp}$ is an isometry.
+Given a subset $\mathcal{A}=\{A_{k}\}_{k\in K}\subset{\mathcal{B}}(\mathcal{H})$ of mutually commuting normal operators and a subset of vectors $Z\subseteq\mathcal{H}$, we denote
+$L_{\mathcal{A},Z}=\bigvee\big{\{}\prod_{k\in K_{0}}A_{k}^{n_{k}}A_{k}^{m_{k}*}h:K_{0}\subseteq K\mbox{ is finite, }n_{k},m_{k}\in\mathbb{N}\mbox{ and }h\in Z\big{\}}$
+where $\vee$ denotes the closed linear span.
+By the Fuglede Theorem, we also have $A_{k}A_{l}^{*}=A_{l}^{*}A_{k}$ for any $k,l\in K$.
+Hence, one can see that $L_{\mathcal{A},Z}$ is a reducing subspace of each $A_{k}$.
+For a normal operator $N$, we denote by $E_{N}$ the associated spectral measure defined on Borel subsets of $\mathbb{C}$. The following theorem is central to proving our main results.
 2.1. Main results
 Theorem 2.1.
-Let $\mathcal{M}$ be an abelian von Neumann algebra of operators in ${\mathcal{B}}(H)$, and let $\{x_{i}\}_{i\in I}$ and $\{y_{i}\}_{i\in I}$ be two orthogonal sets of non-zero vectors in $H$.
-The following statements are equivalents:
+Assume that $\mathcal{H}=L_{\mathcal{N}\cup\mathcal{M},X\cup Y}$. The following statements are equivalent:
   1. (i)
-    There exists a conjugation $C$ on $H$ such that
-    $CTC=T^{*}\;\;\mbox{and}\;\;Cx_{i}=y_{i}\;\;\;\mbox{for all $T\in\mathcal{M}$ and $i\in I$}.$
+    There exists a conjugation $C$ on $\mathcal{H}$ such that $CN_{k}C=M_{k}^{*}$ and $Cx_{i}=y_{i}$ for all $k\in K$ and $i\in I$.
   2. (ii)
-    There exists a conjugate-linear partial isometry $V$ on $H$ such that
-    $VT=T^{*}V,\;\;x_{i}\in{\rm ker}(V)^{\perp},\;\;Vx_{i}=y_{i}\;\;\mbox{and}\;\;Vy_{i}=x_{i}\;\;\;\mbox{for all $T\in\mathcal{M}$ and $i\in I$}.$
-  3. (iii)
-    For every orthogonal projection $P\in\mathcal{M}$ and for all $i,j\in I$, we have
-    $\langle Px_{i},x_{j}\rangle=\langle Py_{j},y_{i}\rangle\;\;\mbox{and}\;\;\langle Px_{i},y_{j}\rangle=\langle Px_{j},y_{i}\rangle.$
+    For all finite subsets $K_{0},K_{1}\subseteq K$, Borel subsets $\{\Delta_{k}\}_{k\in K_{0}}$ and $\{D_{k}\}_{k\in K_{1}}$ of $\mathbb{C}$, and $i,j\in I$, we have
+    $\left\langle\left(\prod_{k\in K_{0}}E_{N_{k}}(\Delta_{k})\right)x_{i},\left(\prod_{l\in K_{1}}E_{M_{l}}(D_{l})\right)x_{j}\right\rangle=\left\langle\left(\prod_{l\in K_{1}}E_{N_{l}}(D_{l})\right)y_{j},\left(\prod_{k\in K_{0}}E_{M_{k}}(\Delta_{k})\right)y_{i}\right\rangle$
+    and
+    $\left\langle\left(\prod_{k\in K_{0}}E_{N_{k}}(\Delta_{k})\right)x_{i},\left(\prod_{l\in K_{1}}E_{M_{l}}(D_{l})\right)y_{j}\right\rangle=\left\langle\left(\prod_{l\in K_{1}}E_{N_{l}}(D_{l})\right)x_{j},\left(\prod_{k\in K_{0}}E_{M_{k}}(\Delta_{k})\right)y_{i}\right\rangle.$
 Remark 2.2.
-Noting that $C(\lambda I)C=(\lambda I)^{*}$ for all conjugations $C$ and $\lambda\in\mathbb{C}$, one can see that when =$\mathcal{M}=\{\lambda I:\lambda\in\mathbb{C}\}$, we recapture [29, Theorem 2.1].
-This theorem states that for two orthonormal sets $\{x_{i}\}_{i\in I}$ and $\{y_{i}\}_{i\in I}$ in $H$, there exists a conjugation $C$ on $H$ satisfying $Cx_{i}=y_{i}$ for every $i\in I$ if and only if $\langle x_{i},y_{j}\rangle=\langle x_{j},y_{i}\rangle$ for all $i,j\in I$.
-It should be noted that their result is important in proving our main results.
-For a normal operator $N$, we denote by $E_{N}$ the associated spectral measure defined on Borel subsets of $\mathbb{C}$.
-If $\mathcal{M}=W^{*}(N_{k}:k\in K)$, meaning that $\mathcal{M}$ is the smallest von Neumann algebra containing a family of mutually commuting normal operators $\{N_{k}\}_{k\in K}$, then the conditions outlined in Theorem 2.1 can be refined as follows.
+Recall that an anti-automorphism $\rho:{\mathcal{B}}(\mathcal{H})\to{\mathcal{B}}(\mathcal{H})$ is a vector space automorphism satisfying $\rho(T^{*})=\rho(T)^{*}$ and $\rho(TS)=\rho(S)\rho(T)$ for all $T,S\in{\mathcal{B}}(\mathcal{H})$.
+It is well-known that each anti-automorphism $\rho:{\mathcal{B}}(\mathcal{H})\to{\mathcal{B}}(\mathcal{H})$, of order two (i.e., $\rho^{2}=I_{{\mathcal{B}}(\mathcal{H})}$), has the form $\rho(T)=CT^{*}C$ for every $T\in{\mathcal{B}}(\mathcal{H})$, where $C$ is a conjugation on $\mathcal{H}$.
+In the light of this fact, it is not hard to see that Theorem 2.1 can be used to obtain conditions under which there exists an anti-automorphism $\rho:{\mathcal{B}}(\mathcal{H})\to{\mathcal{B}}(\mathcal{H})$ of order two such that
+$\rho(T)=T\quad\mbox{and}\quad\rho(N_{k})=M_{k}\;\mbox{for every }k\in K,$
+where $T$ is an operator of the form (1.2).
+For a subspace $M$ of $\mathcal{H}$, denote by $P_{M}\in{\mathcal{B}}(\mathcal{H})$ the orthogonal projection of $\mathcal{H}$ onto $M$. The following corollary is a direct consequence of Theorem 2.1.
+Corollary 2.3.
+Assume that =$\mathcal{H}=\vee\{x_{i},y_{i}:i\in I\}$, and let $M$ and $N$ be subspaces of $\mathcal{H}$ such that $P_{M}$ and $P_{N}$ commute. The following are equivalent:
+  1. (i)
+    There exists a conjugation $C$ on $\mathcal{H}$ such that $CM=N$ and $Cx_{i}=y_{i}$ for every $i\in I$.
+  2. (ii)
+    For all $i,j\in I$ and $r,s\in\{0,1\}$, we have
+    $\left\langle P_{N}^{r}P_{M}^{s}x_{i},x_{j}\right\rangle=\left\langle P_{N}^{s}P_{M}^{r}y_{j},y_{i}\right\rangle\quad\mbox{and}\quad\left\langle P_{N}^{r}P_{M}^{s}x_{i},y_{j}\right\rangle=\left\langle P_{N}^{s}P_{M}^{r}x_{j},y_{i}\right\rangle.$
+Recall that a linear (or conjugate-linear) operator $V$ on $\mathcal{H}$ is called partial isometry if the restriction of $V$ to ${\rm ker}(V)^{\perp}$ is an isometry.
+If $N_{k}=M_{k}$ for every $k\in K$, Theorem 2.1 becomes:
 Theorem 2.4.
-Let $\{N_{k}\}_{k\in K}\subset{\mathcal{B}}(H)$ be a family of mutually commuting normal operators, and let $\{x_{i}\}_{i\in I}$ and $\{y_{i}\}_{i\in I}$ be two orthogonal sets of non-zero vectors in $H$.
 The following statements are equivalent:
   1. (i)
-    There exists a conjugation $C$ on $H$ such that $CTC=T^{*}$ and $Cx_{i}=y_{i}$ for all $T\in W^{*}(N_{k}:k\in K)$ and $i\in I$.
+    There exists a conjugation $C$ on $\mathcal{H}$ such that $CN_{k}C=N_{k}^{*}$ and $Cx_{i}=y_{i}$ for all $k\in K$ and $i\in I$.
   2. (ii)
-    There exists a conjugation $C$ on $H$ such that $CN_{k}C=N_{k}^{*}$ and $Cx_{i}=y_{i}$ for all $k\in K$ and $i\in I$.
+    There exists a conjugate-linear partial isometry $V$ on $\mathcal{H}$ such that $VN_{k}=N_{k}^{*}V$, $x_{i}\in{\rm ker}(V)^{\perp}$, $Vx_{i}=y_{i}$ and $Vy_{i}=x_{i}$ for all $k\in K$ and $i\in I$.
   3. (iii)
-    There exists a conjugate-linear partial isometry $V$ on $H$ such that $VN_{k}=N_{k}^{*}V$, $x_{i}\in{\rm ker}(V)^{\perp}$, $Vx_{i}=y_{i}$ and $Vy_{i}=x_{i}$ for all $k\in K$ and $i\in I$.
-  4. (iv)
     For all finite subsets $K_{0}\subseteq K$, Borel subsets $\{\Delta_{k}\}_{k\in K_{0}}$ of $\mathbb{C}$, and $i,j\in I$ we have
@@
     $\left\langle\left(\prod_{k\in K_{0}}E_{N_{k}}(\Delta_{k})\right)x_{i},y_{j}\right\rangle=\left\langle\left(\prod_{k\in K_{0}}E_{N_{k}}(\Delta_{k})\right)x_{j},y_{i}\right\rangle.$
-Given non-zero vectors $x,y\in H$, we denote by $x\otimes y$ the rank-one operator defined by $(x\otimes y)h=\langle h,y\rangle x$ for every $h\in H$.
-If $I$ is of cardinal one, then Theorem 2.3 reduces to:
+Remark 2.5.
+Taking $N_{k}=0$, for every $k\in K$, in Theorem 2.4, we recapture [30, Theorem 2.1] which states that there exists a conjugation $C$ on $\mathcal{H}$ satisfying $Cx_{i}=y_{i}$ for every $i\in I$ if and only if $\langle x_{i},y_{j}\rangle=\langle x_{j},y_{i}\rangle$ for all $i,j\in I$.
+Remark 2.6.
+One can use Theorem 2.4 to show that if the subspaces $M$ and $N$ in Corollary 2.3 are equal, then the condition “=$\mathcal{H}=\vee\{x_{i},y_{i}:i\in I\}$” is no longer needed.
+Corollary 2.7.
+Let $\{r_{i}\}_{i\in I}$ be a set of distinct positive scalars, and let $T=\sum_{i\in I}r_{i}x_{i}\otimes y_{i}$. The following statements are equivalent:
+  1. (i)
+    There exists a conjugation $C$ on $\mathcal{H}$ such that $N_{k}+\lambda T$ is $C$-symmetric for all $k\in K$ and $\lambda\in\mathbb{C}$.
+  2. (ii)
+    There exist unimodular scalars $\lambda_{i}\in\mathbb{C}$, for $i\in I$, such that
+    $\lambda_{i}\left\langle\left(\prod_{k\in K_{0}}E_{N_{k}}(\Delta_{k})\right)x_{i},x_{j}\right\rangle=\lambda_{j}\left\langle\left(\prod_{k\in K_{0}}E_{N_{k}}(\Delta_{k})\right)y_{j},y_{i}\right\rangle$
+    and
+    $\lambda_{i}\left\langle\left(\prod_{k\in K_{0}}E_{N_{k}}(\Delta_{k})\right)x_{i},y_{j}\right\rangle=\lambda_{j}\left\langle\left(\prod_{k\in K_{0}}E_{N_{k}}(\Delta_{k})\right)x_{j},y_{i}\right\rangle$
+    for all finite subsets $K_{0}\subseteq K$, Borel subsets $\{\Delta_{k}\}_{k\in K_{0}}$ of $\mathbb{C}$, and $i,j\in I$.
+Proof.
+(ii)$\implies$(i).
+For every $i\in I$, put $\widetilde{x_{i}}=\lambda_{i}x_{i}$.
+It is easy to see that the orthonormal sets $\{\widetilde{x_{i}}\}_{i\in I}$ and $\{y_{i}\}_{i\in I}$ satisfy the assertion (iii) in Theorem 2.4.
+Thus, there exists a conjugation $C$ on $\mathcal{H}$ such that $CN_{k}C=N_{k}^{*}$ and $C\widetilde{x_{i}}=y_{i}$ for all $k\in K$ and $i\in I$.
+Moreover, since $C^{-1}=C$, we also have $Cy_{i}=\widetilde{x_{i}}$.
+Now, by (1.1), one can check that $C(x_{i}\otimes y_{i})C=y_{i}\otimes x_{i}$ for every $i\in I$.
+Therefore, $CTC=T^{*}$, and consequently, $C(N_{k}+\lambda T)C=(N_{k}+\lambda T)^{*}$ for all $\lambda\in\mathbb{C}$ and $k\in K$.
+(i)$\implies$(ii).
+Since the class of $C$-symmetric operators is a vector space and $\lambda$ is arbitrary, we have $CN_{k}C=N_{k}^{*}$ for every $k\in K$ and $CTC=T^{*}$.
+The later equality yields $CT^{*}TC=(CT^{*}C)(CTC)=TT^{*}$, and consequently $C(T^{*}T-r_{i}I_{\mathcal{H}})C=TT^{*}-r_{i}I_{\mathcal{H}}$ for every $i\in I$.
+In particular, $C{\rm ker}(TT^{*}-r_{i}I_{\mathcal{H}})\subseteq{\rm ker}(T^{*}T-r_{i}I_{\mathcal{H}})$ for every $i\in I$.
+Therefore, $C\left(\vee\{x_{i}\}\right)=\vee\{y_{i}\}$ for every $i\in I$.
+Since $C$ is isometric, there exist unimodular scalars $\lambda_{i}$, for $i\in I$, such that $C(\lambda_{i}x_{i})=y_{i}$ for every $i\in I$.
+Now, using Theorem 2.4, one can easily see that the scalars $\lambda_{i}$, $i\in I$, satisfy all the desired equalities.
+∎
+If $I$ is of cardinal one, then we get a linear version of Theorem 2.4 as follows:
 Theorem 2.8.
-Let $\{N_{k}\}_{k\in K}\subset{\mathcal{B}}(H)$ be a family of mutually commuting normal operators, and let $x,y\in H$ be non-zero vectors. The following statements are equivalent:
+Let $x,y\in\mathcal{H}$ be unit vectors. The following statements are equivalent:
   1. (i)
-    There exists a conjugation $C$ on $H$ satisfying $CN_{k}C=N_{k}^{*}$ and $Cx=y$ for every $k\in K$.
+    There exists a conjugation $C$ on $\mathcal{H}$ satisfying $CN_{k}C=N_{k}^{*}$ and $Cx=y$ for every $k\in K$.
   2. (ii)
-    There exists a unitary operator $U\in{\mathcal{B}}(H)$ satisfying $UN_{k}=N_{k}U$ and $Ux=y$ for every $k\in K$.
+    There exists a unitary operator $U\in{\mathcal{B}}(\mathcal{H})$ satisfying $UN_{k}=N_{k}U$ and $Ux=y$ for every $k\in K$.
   3. (iii)
-    There exists a partial isometry $V\in{\mathcal{B}}(H)$ satisfying $x\in{\rm ker}(V)^{\perp}$, $VN_{k}=N_{k}V$ and $Vx=y$ for every $k\in K$.
+    There exists a partial isometry $V\in{\mathcal{B}}(\mathcal{H})$ satisfying $x\in{\rm ker}(V)^{\perp}$, $VN_{k}=N_{k}V$ and $Vx=y$ for every $k\in K$.
   4. (iv)
@@
   5. (v)
-    $\|x\|=\|y\|$ and there exists a conjugation $C$ on $H$ such that $N_{k}+\lambda x\otimes y$ is $C$-symmetric for all $k\in K$ and $\lambda\in\mathbb{C}$.
-The proofs of Theorems 2.1, 2.3, and 2.4 will be given later in this section, following the necessary preliminary preparations.
-A slight reformulation of [1, Theorem 2.1] shows that if $N\in{\mathcal{B}}(H)$ is a normal operator, $U\in{\mathcal{B}}(H)$ is a unitary operator commuting with $N$, and $x\in H$, then for every $\lambda\in\mathbb{C}$, there exists a conjugation $C$ on $H$ such that $N+\lambda x\otimes Ux$ is $C$-symmetric.
+    There exists a conjugation $C$ on $\mathcal{H}$ such that $N_{k}+\lambda x\otimes y$ is $C$-symmetric for all $k\in K$ and $\lambda\in\mathbb{C}$.
+The proofs of Theorems 2.1, 2.4, and 2.8 will be given later in this section, following the necessary preliminary preparations.
+Remark 2.9.
+A slight reformulation of [1, Theorem 2.1] shows that if $N\in{\mathcal{B}}(\mathcal{H})$ is a normal operator, $U\in{\mathcal{B}}(\mathcal{H})$ is a unitary operator commuting with $N$, and $x\in\mathcal{H}$, then for every $\lambda\in\mathbb{C}$, there exists a conjugation $C$ on $\mathcal{H}$ such that $N+\lambda x\otimes Ux$ is $C$-symmetric.
 It can be seen that the implication (ii)$\implies$(v) in Theorem 2.8 generalizes this result, as the conjugation $C$ is independent of $\lambda$.
-A subspace $M$ of $H$ is said to be hyperinvariant for $T\in{\mathcal{B}}(H)$ if $M$ is invariant under every operator that commutes with $T$.
+A subspace $M$ of $\mathcal{H}$ is said to be hyperinvariant for $T\in{\mathcal{B}}(\mathcal{H})$ if $M$ is invariant under every operator that commutes with $T$.
 Recently, in [21], Mashreghi et all.
-proved that $M$ is hyperinvariant for a unitary operator $U\in{\mathcal{B}}(H)$ if and only if $CM\subseteq M$ for every conjugation $C$ on $H$ satisfying $CUC=U^{*}$.
+proved that $M$ is hyperinvariant for a unitary operator $U\in{\mathcal{B}}(\mathcal{H})$ if and only if $CM\subseteq M$ for every conjugation $C$ on $\mathcal{H}$ satisfying $CUC=U^{*}$.
 As a consequence of Theorem 2.8, this result is extended to all normal operators.
 Proposition 2.10.
-Let $N\in{\mathcal{B}}(H)$ be normal and $M$ a subspace of $H$. Then, $M$ is hyperinvariant for $N$ if and only if $CM\subseteq M$ for every conjugation $C$ on $H$ satisfying $CNC=N^{*}$.
-The proof of the necessity is similar to the one in [20]. We include it here for the sake of completeness.
-Proof of Proposition 2.5.
-Assume that $M$ is hyperinvariant for $N$ and denote by $P\in{\mathcal{B}}(H)$ the orthogonal projection onto $M$.
-Let $C$ be a conjugation on $N$ satisfying $CNC=N^{*}$; equivalently, $CN=N^{*}C$.
-Clearly, we need to show that $CP=PC$.
-According to [25, Proposition 6.9], there exists a Borel subset $\Delta\subseteq\mathbb{C}$ such that $P=E_{N}(\Delta)$.
-Consequently, Lemma 2.11 implies that $CP=CE_{N}(\Delta)=E_{N}(\Delta)C=PC$.
-For the sufficiency, we need to show that for each $T\in\{N\}^{\prime}$, the commutant of $N$, $TM\subseteq M$.
-Additionally, since $\{N\}^{\prime}$ is a von Neumann algebra, it is spanned by its unitary operators ([9, Page 61, Proposition 13.3 (b)]).
+Let $N\in{\mathcal{B}}(\mathcal{H})$ be normal and $M$ a subspace of $\mathcal{H}$. The following statements are equivalents:
+  1. (i)
+    $M$ is hyperinvariant for $N$.
+  2. (ii)
+    $TM\subseteq M$ for every conjugate-linear operator $T$ on $\mathcal{H}$ satisfying $TN=N^{*}T$.
+  3. (iii)
+    $CM\subseteq M$ for every conjugation $C$ on $\mathcal{H}$ satisfying $CNC=N^{*}$.
+Proof.
+(i)$\implies$(ii).
+Assume that $M$ is hyperinvariant for $N$, and let $T$ be a conjugate-linear operator on $\mathcal{H}$ such that $TN=N^{*}T$.
+Clearly, it suffices to show that $TP_{M}=P_{M}T$.
+According to [26, Proposition 6.9], there exists a Borel subset $\Delta\subseteq\mathbb{C}$ such that $P_{M}=E_{N}(\Delta)$.
+Consequently, Lemma 2.15 implies that $TP_{M}=TE_{N}(\Delta)=E_{N}(\Delta)T=P_{M}T$.
+(ii)$\implies$(iii) If $C$ is a conjugation on $\mathcal{H}$ such that $CNC=N^{*}$, and since $C^{-1}=C$, it follows that $CN=N^{*}C$, which leads to the desired containment.
+(iii)$\implies$(i).
+We need to show that for each $T\in\{N\}^{\prime}$, the commutant of $N$, $TM\subseteq M$.
+Additionally, since $\{N\}^{\prime}$ is a von Neumann algebra, it is spanned by its unitary operators ([10, p.
+61, Proposition 13.3 (b)]).
 Consequently, we can assume without loss of generality that $T$ is unitary.
 Let $x\in M$ be non-zero, and put $y=Tx$.
-Then, according to Theorem 2.4, there exists a conjugation $C$ on $H$ satisfying $CNC=N^{*}$ and $Cx=y$.
+Then, according to Theorem 2.8, there exists a conjugation $C$ on $\mathcal{H}$ satisfying $CNC=N^{*}$ and $Cx=y$.
 Hence, by hypothesis, $Tx=y=Cx\in M$.
@@
 ∎
-The following example shows that the result in Proposition 2.5 cannot be extended to an arbitrary operator in ${\mathcal{B}}(H)$.
+The following example shows that the result in Proposition 2.10 cannot be extended to an arbitrary operator in ${\mathcal{B}}(\mathcal{H})$.
 Example 2.11.
-Let $x$ and $y$ be orthonormal vectors in $H$, and put $T=A+{\rm i}B$ where $A=x\otimes x$ and $B=(x+y)\otimes(x+y)$.
+Let $x$ and $y$ be orthonormal vectors in $\mathcal{H}$, and put $T=A+{\rm i}B$ where $A=x\otimes x$ and $B=(x+y)\otimes(x+y)$.
 Clearly, $A$ and $B$ are self-adjoint operators; moreover, $A=\frac{1}{2}(T+T^{*})$.
-It follows that if $C$ is a conjugation on $H$ such that $CTC=T^{*}$, then $CAC=A$.
-Hence, if $M={\rm span}\{x\}$, then $CM=M$.
-Therefore, the subspace $M$ satisfies the necessary condition in Proposition 2.5; however, $T\in\{T\}^{\prime}$ and $TM\nsubset M$.
-Given a conjugation $C$ on $H$, an operator $T\in{\mathcal{B}}(H)$ is said to be $C$-normal if $CTT^{*}C=T^{*}T$.
+It follows that if $C$ is a conjugation on $\mathcal{H}$ such that $CTC=T^{*}$, then $CAC=A$.
+Hence, if $M=\vee\{x\}$, then $CM=M$.
+Therefore, the subspace $M$ satisfies condition (iii) in Proposition 2.10; however, $T\in\{T\}^{\prime}$ and $TM\nsubset M$.
+Given a conjugation $C$ on $\mathcal{H}$, an operator $T\in{\mathcal{B}}(\mathcal{H})$ is said to be $C$-normal if $CTT^{*}C=T^{*}T$.
 When $T$ is $C$-normal for some conjugation $C$, we say that $T$ is conjugate-normal.
@@
 Corollary 2.12.
-Let $T\in{\mathcal{B}}(H)$. The following statements are equivalent:
+Let $T\in{\mathcal{B}}(\mathcal{H})$. The following statements are equivalent:
   1. (i)
@@
   2. (ii)
-    For every non-zero $x\in H$, there exists a conjugation $C$ on $H$ such that $T$ is $C$-symmetric and $Cx=x$;
+    For every unit vector $x\in\mathcal{H}$, there exists a conjugation $C$ on $\mathcal{H}$ such that $T$ is $C$-symmetric and $Cx=x$.
   3. (iii)
-    For every non-zero $x\in H$, there exists a conjugation $C$ on $H$ such that $T$ is $C$-normal and $Cx=x$.
+    For every unit vector $x\in\mathcal{H}$, there exists a conjugation $C$ on $\mathcal{H}$ such that $T$ is $C$-normal and $Cx=x$.
 Proof.
@@
 (ii)$\implies$(iii) is obvious.
-(iii)$\implies$(ii). Let $x\in H$ be non-zero. Then, by hypothesis, there exists a conjugation $C$ on $H$ such that $Cx=x$ and $CTT^{*}C=T^{*}T$. It follows that
+(iii)$\implies$(i). Let $x\in\mathcal{H}$ be a unit vector. Then, by hypothesis, there exists a conjugation $C$ on $\mathcal{H}$ such that $Cx=x$ and $CTT^{*}C=T^{*}T$. It follows that
 $\langle T^{*}Tx,x\rangle=\langle CTT^{*}Cx,x\rangle=\langle Cx,CCTT^{*}Cx\rangle=\langle x,TT^{*}x\rangle=\langle TT^{*}x,x\rangle.$
 Since $x$ was arbitrary, we infer that $T^{*}T=TT^{*}$, the desired equality. ∎
-2.2. Proofs of Theorems 2.1, 2.3 and 2.4
+2.2. Proof of Theorems 2.1, 2.4, and 2.8
 The proofs require some preparation and will be provided at the end of this section.
-In what follows, we denote the set of all complex polynomials in two variables by $\mathbb{C}[x,y]$.
+In what follows, we denote the set of all complex polynomials in $n$ variables by $\mathcal{P}_{n}(\mathbb{C})$.
 We begin with the following technical lemma.
 Lemma 2.13.
-Let $N\in{\mathcal{B}}(H)$ and $M\in\mathcal{B}(K)$ be normal operators, and let $x,y\in H$ and $h,k\in K$ be vectors such that
-$\langle E_{N}(\Delta)x,y\rangle=\langle E_{M}(\Delta)h,k\rangle\quad\mbox{for every Borel subset $\Delta\subseteq\mathbb{C}$}.$
-Then, for every $p\in\mathbb{C}[x,y]$, we have $\langle p(N,N^{*})x,y\rangle=\langle p(M,M^{*})h,k\rangle.$
+Let $\{N_{i}\}_{1\leq i\leq n}\subset{\mathcal{B}}(\mathcal{H})$ and $\{M_{i}\}_{1\leq i\leq n}\subset\mathcal{B}(\mathcal{K})$ be two finite sets of mutually commuting normal operators, and let $x,y\in\mathcal{H}$ and $h,k\in\mathcal{K}$ be vectors such that
+$\left\langle\prod_{i=1}^{n}\left(E_{N_{i}}(\Delta_{i})\right)x,y\right\rangle=\left\langle\prod_{i=1}^{n}\left(E_{M_{i}}(\Delta_{i})\right)h,k\right\rangle\quad\mbox{for every Borel subset $\Delta_{i}\subseteq\mathbb{C}$}.$
+Then, for every $p\in\mathcal{P}_{2n}(\mathbb{C})$, we have
+$\left\langle p(N_{1},N_{1}^{*},\dots,N_{n},N_{n}^{*})x,y\right\rangle=\left\langle p(M_{1},M_{1}^{*},\dots,M_{n},M_{n}^{*})h,k\right\rangle.$
 Proof.
-Let $p\in\mathbb{C}[x,y]$, and put $f(z)=p(z,\overline{z})\chi_{K}(z)$ where $K=\sigma(N)\cup\sigma(M)$ and $\chi_{K}$ is the characteristic function of $K$.
-Then, $f$ is a bounded Borel function on $\mathbb{C}$; furthermore, $f(N)=p(N,N^{*})$ and $f(M)=p(M,M^{*})$.
-This implies that
-$\langle p(N,N^{*})x,y\rangle=\langle f(N)x,y\rangle=\int f{\,\rm d}\langle E_{N}(.)x,y\rangle$
+Let $p\in\mathcal{P}_{2n}(\mathbb{C})$.
+By the commutativity assumption, we may suppose that $p(\alpha_{1},\lambda_{1},\dots,\alpha_{n},\lambda_{n})=\prod_{j=1}^{n}\alpha_{i}^{r_{i}}\lambda_{i}^{s_{i}}$ where $r_{i},s_{i}\in\mathbb{N}$.
+For each $i\in\{1,\dots,n\}$, put $f_{i}(z)=z^{r_{i}}\overline{z}^{s_{i}}\chi_{K_{i}}(z)$ where $K_{i}=\sigma(N_{i})\cup\sigma(M_{i})$ and $\chi_{K_{i}}$ is the characteristic function of $K_{i}$.
+Then, for every $i\in\{1,\dots,n\}$, $f_{i}$ is a bounded Borel function on $\mathbb{C}$.
+Additionally,
+(2.1) $f_{i}(N_{i})=N_{i}^{r_{i}}N_{i}^{*s_{i}}\quad\mbox{and}\quad f_{i}(M_{i})=M_{i}^{r_{i}}M_{i}^{*s_{i}}\quad\mbox{for every }i\in\{1,\dots,n\}.$
+Fix Borel subsets $\{\Delta_{i}\}_{2\leq i\leq n}\subseteq\mathbb{C}$ and denote by $\mu_{1}$ and $\nu_{1}$ the complex-valued measures on Borel subsets of $\mathbb{C}$ given by
+$\mu_{1}(\Delta)=\left\langle E_{N_{1}}(\Delta)\prod_{i=2}^{n}\left(E_{N_{i}}(\Delta_{i})\right)x,y\right\rangle\quad\mbox{and}\quad\nu_{1}(\Delta)=\left\langle E_{M_{1}}(\Delta)\prod_{i=2}^{n}\left(E_{M_{i}}(\Delta_{i})\right)h,k\right\rangle.$
+Then,
+$\left\langle N_{1}^{r_{1}}N_{1}^{*s_{1}}\prod_{i=2}^{n}\left(E_{N_{i}}(\Delta_{i})\right)x,y\right\rangle=\left\langle f_{1}(N_{1})\prod_{i=2}^{n}\left(E_{N_{i}}(\Delta_{i})\right)x,y\right\rangle=\int f_{1}{\,\rm d}\mu_{1}$
 and
-$\langle p(M,M^{*})h,k\rangle=\langle f(M)h,k\rangle=\int f{\,\rm d}\langle E_{M}(.)h,k\rangle.$
-As the complex-valued measures $\langle E_{N}(.)x,y\rangle$ and $\langle E_{M}(.)h,k\rangle$ are presumed equal by hypothesis, we deduce $\int f{\,\rm d}\langle E_{N}(.)x,y\rangle=\int f{\,\rm d}\langle E_{M}(.)h,k\rangle$.
-Consequently, $\langle p(N,N^{*})x,y\rangle=\langle p(M,M^{*})x,y\rangle.$
-Given a normal operator $N\in{\mathcal{B}}(H)$ and a family of vectors $X\subset H$, we denote
-$L_{N,X}=\bigvee\big{\{}p(N,N^{*})h:p\in\mathbb{C}[x,y]\mbox{ and }h\in X\big{\}}$
-where $\vee$ denotes the closed linear span. It is easy to see that $L_{N,X}$ is a reducing subspace of $N$. The restriction of $N$ to $L_{N,X}$ will be denoted by $N_{X}$.
+$\left\langle M_{1}^{r_{1}}M_{1}^{*s_{1}}\prod_{i=2}^{n}\left(E_{M_{i}}(\Delta_{i})\right)h,k\right\rangle=\left\langle f_{1}(M_{1})\prod_{i=2}^{n}\left(E_{M_{i}}(\Delta_{i})\right)h,k\right\rangle=\int f_{1}{\,\rm d}\nu_{1}.$
+As the measures $\mu_{1}$ and $\nu_{1}$ are presumed equal by hypothesis, we deduce
+$\left\langle N_{1}^{r_{1}}N_{1}^{*s_{1}}\prod_{i=2}^{n}\left(E_{N_{i}}(\Delta_{i})\right)x,y\right\rangle=\left\langle M_{1}^{r_{1}}M_{1}^{*s_{1}}\prod_{i=2}^{n}\left(E_{M_{i}}(\Delta_{i})\right)h,k\right\rangle.$
+By the Fuglede Theorem, we obtain $\mu_{2}(\Delta_{2})=\nu_{2}(\Delta_{2})$ where
+$\mu_{2}(\Delta_{2}):=\left\langle E_{N_{2}}(\Delta_{2})N_{1}^{r_{1}}N_{1}^{*s_{1}}\prod_{i=3}^{n}\left(E_{N_{i}}(\Delta_{i})\right)x,y\right\rangle$
+and
+$\nu_{2}(\Delta_{2}):=\left\langle E_{M_{2}}(\Delta_{2})M_{1}^{r_{1}}M_{1}^{*s_{1}}\prod_{i=3}^{n}\left(E_{M_{i}}(\Delta_{i})\right)h,k\right\rangle.$
+Since $\Delta_{2}$ was arbitrary, the complex-valued measures $\mu_{2}$ and $\nu_{2}$ are equal. Integrating $f_{2}$ with respect to these measures and taking into account (2.1), we obtain
+$\left\langle N_{2}^{r_{2}}N_{2}^{*s_{2}}N_{1}^{r_{1}}N_{1}^{*s_{1}}\prod_{i=3}^{n}\left(E_{N_{i}}(\Delta_{i})\right)x,y\right\rangle=\left\langle M_{2}^{r_{2}}M_{2}^{*s_{2}}M_{1}^{r_{1}}M_{1}^{*s_{1}}\prod_{i=3}^{n}\left(E_{M_{i}}(\Delta_{i})\right)h,k\right\rangle.$
+The desired equality will follow by continuing the previous process. ∎
 Proposition 2.14.
-Let $N\in{\mathcal{B}}(H)$ and $M\in\mathcal{B}(K)$ be normal operators, and let $X=\{x_{i}\}_{i\in I}\subseteq H$ and $Y=\{y_{i}\}_{i\in I}\subseteq K$ be two families of vectors such that
-$\langle E_{N}(\Delta)x_{i},x_{j}\rangle=\langle E_{M}(\Delta)y_{j},y_{i}\rangle\quad\mbox{for all Borel subsets $\Delta\subseteq\mathbb{C}$ and $i,j\in I$}.$
-Then, there exists a conjugate-linear surjective isometry $W:L_{N,X}\to L_{M,Y}$ such that
-$Wx_{i}=y_{i}\quad\mbox{and}\quad Wp(N_{X},N_{X}^{*})=p(M_{Y},M_{Y}^{*})^{*}W\quad\mbox{for all $i\in I$ and $p\in\mathbb{C}[x,y]$}.$
+Let $\mathcal{A}=\{A_{k}\}_{k\in K}\subset{\mathcal{B}}(\mathcal{H})$ and =$\mathcal{B}=\{B_{k}\}_{k\in K}\subset{\mathcal{B}}(\mathcal{K})$ be two sets of mutually commuting normal operators, and let $A=\{a_{i}\}_{i\in I}\subseteq\mathcal{H}$ and $B=\{b_{i}\}_{i\in I}\subseteq\mathcal{K}$ be two subsets of vectors such that
+$\left\langle\left(\prod_{k\in K_{0}}E_{A_{k}}(\Delta_{k})\right)a_{i},a_{j}\right\rangle=\left\langle\left(\prod_{k\in K_{0}}E_{B_{k}}(\Delta_{k})\right)b_{j},b_{i}\right\rangle$
+for all finite subsets $K_{0}\subseteq K$, Borel subsets $\{\Delta_{k}\}_{k\in K_{0}}$ of $\mathbb{C}$ and $i,j\in I$.
+Then, there exists a conjugate-linear surjective isometry $W:L_{\mathcal{A},A}\to L_{\mathcal{B},B}$ such that
+$Wa_{i}=b_{i}\quad\mbox{and}\quad Wp(T_{k_{1}},T_{k_{1}}^{*},\dots,T_{k_{n}},T_{k_{n}}^{*})=p(S_{k_{1}},S_{k_{1}}^{*},\dots,S_{k_{n}},S_{k_{n}}^{*})^{*}W$
+for all $i\in I$, $n\in\mathbb{Z}^{+}$, $p\in\mathcal{P}_{2n}(\mathbb{C})$ and $\{k_{l}\}_{1\leq l\leq n}\subseteq K$, where $T_{k_{i}}$ and $S_{k_{i}}$ are the restrictions of $A_{k_{i}}$ and $B_{k_{i}}$ to $L_{\mathcal{A},A}$ and $L_{\mathcal{B},B}$, respectively.
 Proof.
-According to Lemma 2.8, for every polynomial $p\in\mathbb{C}[x,y]$ and all $i,j\in I$, we have
-(2.1) $\left\langle p(N,N^{*})x_{i},x_{j}\right\rangle=\left\langle p(M,M^{*})y_{j},y_{i}\right\rangle.$
+According to Lemma 2.13, for every subset $\{k_{l}\}_{1\leq l\leq n}\subseteq K$, every polynomial $p\in\mathcal{P}_{2n}(\mathbb{C})$, and all $i,j\in I$, we have
+(2.2) $\left\langle p(A_{k_{1}},A_{k_{1}}^{*},\dots,A_{k_{n}},A_{k_{n}}^{*})a_{i},a_{j}\right\rangle=\left\langle p(B_{k_{1}},B_{k_{1}}^{*},\dots,B_{k_{n}},B_{k_{n}}^{*})b_{j},b_{i}\right\rangle.$
 For convenience, put
-$H_{0}={\rm span}\big{\{}p(N,N^{*})x_{i}:p\in\mathbb{C}[x,y],i\in I\big{\}}$
+$\mathcal{H}_{0}={\rm span}\big{\{}p(A_{k_{1}},A_{k_{1}}^{*},\dots,A_{k_{n}},A_{k_{n}}^{*})a_{i}:n\in\mathbb{Z}^{+},p\in\mathcal{P}_{2n}(\mathbb{C}),\{k_{l}\}_{1\leq l\leq n}\subseteq K,i\in I\big{\}}$
 and
-$K_{0}={\rm span}\big{\{}p(M,M^{*})y_{i}:p\in\mathbb{C}[x,y],i\in I\big{\}},$
-and let $W:H_{0}\to K_{0}$ be the conjugate-linear transformation given by
-$W(p(N,N^{*})x_{i})=p(M,M^{*})^{*}y_{i}\quad\mbox{for all $p\in\mathbb{C}[x,y]$and $i\in I$.}$
+$\mathcal{K}_{0}={\rm span}\big{\{}p(B_{k_{1}},B_{k_{1}}^{*},\dots,B_{k_{n}},B_{k_{n}}^{*})b_{i}:n\in\mathbb{Z}^{+},p\in\mathcal{P}_{2n}(\mathbb{C}),\{k_{l}\}_{1\leq l\leq n}\subseteq K,i\in I\big{\}},$
+and let $W:\mathcal{H}_{0}\to\mathcal{K}_{0}$ be the conjugate-linear transformation given by
+$W\left(p(A_{k_{1}},A_{k_{1}}^{*},\dots,A_{k_{n}},A_{k_{n}}^{*})a_{i}\right)=p(B_{k_{1}},B_{k_{1}}^{*},\dots,B_{k_{n}},B_{k_{n}}^{*})^{*}b_{i}$
+for all $n\geq 1$, $p\in\mathcal{P}_{2n}(\mathbb{C})$, $\{k_{l}\}_{1\leq l\leq n}\subseteq K$ and $i\in I$.
 We will first show that $W$ is well-defined.
-Arbitrarily choose $h\in H_{0}$; then, there exists a finite subset $J$ of $I$ and polynomials $p_{i}\in\mathbb{C}[x,y]$, for $i\in J$, such that $h=\sum_{i\in J}p_{i}(N,N^{*})x_{i}$.
+Arbitrarily choose $h\in\mathcal{H}_{0}$; then, there exists a finite subset $J$ of $I$, $\{k_{i,l}\}_{1\leq l\leq n_{i}}\subseteq K$, for $i\in J$, and polynomials $p_{i}\in\mathcal{P}_{2n_{i}}(\mathbb{C})$, for $i\in J$, such that
+$h=\sum_{i\in J}p_{i}(A_{k_{i,1}},A_{k_{i,1}}^{*},\dots,A_{k_{i,n_{i}}},A_{k_{i,n_{i}}}^{*})a_{i}.$
 Then,
-$\|h\|^{2}$ $=\left\langle\sum_{i\in J}p_{i}(N,N^{*})x_{i},\sum_{j\in J}p_{j}(N,N^{*})x_{j}\right\rangle$
-$=\sum_{i\in J}\sum_{j\in J}\left\langle p_{i}(N,N^{*})x_{i},p_{j}(N,N^{*})x_{j}\right\rangle$
-$=\sum_{i\in J}\sum_{j\in J}\left\langle p_{j}(N,N^{*})^{*}p_{i}(N,N^{*})x_{i},x_{j}\right\rangle$
-[…]
-$=\sum_{i\in J}\sum_{j\in J}\left\langle p_{j}(M,M^{*})^{*}y_{j},p_{i}(M,M^{*})^{*}y_{i}\right\rangle$
-$=\left\langle\sum_{j\in J}p_{j}(M,M^{*})^{*}y_{j},\sum_{i\in J}p_{i}(M,M^{*})^{*}y_{i}\right\rangle=\|Wh\|^{2}.$
-Hence, if $h,k\in H_{0}$ are such that $\|h-k\|=0$, we get $\|Wh-Wk\|=\|W(h-k)\|=\|h-k\|=0$.
+$\|h\|^{2}$ $=\left\langle\sum_{i\in J}p_{i}(A_{k_{i,1}},A_{k_{i,1}}^{*},\dots,A_{k_{i,n_{i}}},A_{k_{i,n_{i}}}^{*})a_{i},\sum_{j\in J}p_{j}(A_{k_{j,1}},A_{k_{j,1}}^{*},\dots,A_{k_{j,n_{j}}},A_{k_{j,n_{j}}}^{*})a_{j}\right\rangle$
+$=\sum_{i\in J}\sum_{j\in J}\left\langle p_{i}(A_{k_{i,1}},A_{k_{i,1}}^{*},\dots,A_{k_{i,n_{i}}},A_{k_{i,n_{i}}}^{*})a_{i},p_{j}(A_{k_{j,1}},A_{k_{j,1}}^{*},\dots,A_{k_{j,n_{j}}},A_{k_{j,n_{j}}}^{*})a_{j}\right\rangle$
+$=\sum_{i\in J}\sum_{j\in J}\left\langle p_{i}(A_{k_{i,1}},A_{k_{i,1}}^{*},\dots,A_{k_{i,n_{i}}},A_{k_{i,n_{i}}}^{*})p_{j}(A_{k_{j,1}},A_{k_{j,1}}^{*},\dots,A_{k_{j,n_{j}}},A_{k_{j,n_{j}}}^{*})^{*}a_{i},a_{j}\right\rangle.$
+It follows by (2.2) that
+$\|h\|^{2}$ $=\sum_{i\in J}\sum_{j\in J}\left\langle p_{i}(B_{k_{i,1}},B_{k_{i,1}}^{*},\dots,B_{k_{i,n_{i}}},B_{k_{i,n_{i}}}^{*})p_{j}(B_{k_{j,1}},B_{k_{j,1}}^{*},\dots,B_{k_{j,n_{j}}},B_{k_{j,n_{j}}}^{*})^{*}b_{j},b_{i}\right\rangle$
+$=\sum_{i\in J}\sum_{j\in J}\left\langle p_{j}(B_{k_{j,1}},B_{k_{j,1}}^{*},\dots,B_{k_{j,n_{j}}},B_{k_{j,n_{j}}}^{*})^{*}b_{j},p_{i}(B_{k_{i,1}},B_{k_{i,1}}^{*},\dots,B_{k_{i,n_{i}}},B_{k_{i,n_{i}}}^{*})^{*}b_{i}\right\rangle$
+$=\left\langle\sum_{j\in J}p_{j}(B_{k_{j,1}},B_{k_{j,1}}^{*},\dots,B_{k_{j,n_{j}}},B_{k_{j,n_{j}}}^{*})^{*}b_{j},\sum_{i\in J}p_{i}(B_{k_{i,1}},B_{k_{i,1}}^{*},\dots,B_{k_{i,n_{i}}},B_{k_{i,n_{i}}}^{*})^{*}b_{i}\right\rangle$
+$=\|Wh\|^{2}.$
+Hence, if $h,k\in\mathcal{H}_{0}$ are such that $\|h-k\|=0$, we get $\|Wh-Wk\|=\|W(h-k)\|=\|h-k\|=0$.
 Therefore, $W$ is a well-defined conjugate-linear isometry.
 Moreover, by its construction, $W$ is surjective.
-By density of $H_{0}$ and $K_{0}$ in $L_{N,X}$ and $L_{M,Y}$, respectively, it follows that $W$ can be extended to a conjugate-linear surjective isometry from $L_{N,X}$ to $L_{M,Y}$, which we also denote by $W$.
-Furthermore, it can be seen that $Wx_{i}=y_{i}$ for every $i\in I$.
-Now, fix $p\in\mathbb{C}[x,y]$ and let us check that $Wp(N_{X},N_{X}^{*})=p(M_{Y},M_{Y}^{*})^{*}W$. Let $q\in\mathbb{C}[x,y]$, $i\in I$, and put $h=q(N,N^{*})x_{i}$. Then,
-$Wp(N_{X},N_{X}^{*})h=Wp(N,N^{*})h$ $=Wp(N,N^{*})q(N,N^{*})x_{i}=p(M,M^{*})^{*}q(M,M^{*})^{*}y_{i}$
-$=p(M,M^{*})^{*}Wq(N,N^{*})x_{i}=p(M_{Y},M_{Y}^{*})^{*}Wh.$
-Since $q$ and $i$ were arbitrary, and $Wp(N_{X},N_{X}^{*})$ and $p(M_{Y},M_{Y}^{*})^{*}W$ are both conjugate-linear, we obtain $Wp(N_{X},N_{X}^{*})h=p(M_{Y},M_{Y}^{*})^{*}Wh$ for every $h\in H_{0}$.
-Consequently, by density of $H_{0}$ in $L_{N,X}$, we conclude that $Wp(N_{X},N_{X}^{*})=p(M_{Y},M_{Y}^{*})^{*}W$.
-This completes the proof.
+By density of $\mathcal{H}_{0}$ and $\mathcal{K}_{0}$ in $L_{\mathcal{A},A}$ and $L_{\mathcal{B},B}$, respectively, it follows that $W$ can be extended to a conjugate-linear surjective isometry from $L_{\mathcal{A},A}$ to $L_{\mathcal{B},B}$, which we also denote by $W$.
+Furthermore, it can be seen that $Wa_{i}=b_{i}$ for every $i\in I$.
+Now, let $n\in\mathbb{Z}^{+}$, $p\in\mathcal{P}_{2n}(\mathbb{C})$ and $\{k_{l}\}_{1\leq l\leq n}\subseteq K$, and let us check that
+$T:=Wp(T_{k_{1}},T_{k_{1}}^{*},\dots,T_{k_{n}},T_{k_{n}}^{*})=S:=p(S_{k_{1}},S_{k_{1}}^{*},\dots,S_{k_{n}},S_{k_{n}}^{*})^{*}W.$
+Let $m\in\mathbb{Z}^{+}$, $q\in\mathcal{P}_{2m}(\mathbb{C})$, $\{t_{r}\}_{1\leq r\leq m}\subseteq K$ and $i\in I$, and put $h=q(A_{t_{1}},A_{t_{1}}^{*},\dots,A_{t_{m}},A_{t_{m}}^{*})a_{i}$. Then,
+$Th$ $=Wp(T_{k_{1}},T_{k_{1}}^{*},\dots,T_{k_{n}},T_{k_{n}}^{*})h$
+$=Wp(A_{k_{1}},A_{k_{1}}^{*},\dots,A_{k_{n}},A_{k_{n}}^{*})h$
+$=Wp(A_{k_{1}},A_{k_{1}}^{*},\dots,A_{k_{n}},A_{k_{n}}^{*})q(A_{t_{1}},A_{t_{1}}^{*},\dots,A_{t_{m}},A_{t_{m}}^{*})a_{i}$
+$=q(B_{t_{1}},B_{t_{1}}^{*},\dots,B_{t_{m}},B_{t_{m}}^{*})^{*}p(B_{k_{1}},B_{k_{1}}^{*},\dots,B_{k_{n}},B_{k_{n}}^{*})^{*}b_{i}$
+$=p(B_{k_{1}},B_{k_{1}}^{*},\dots,B_{k_{n}},B_{k_{n}}^{*})^{*}q(B_{t_{1}},B_{t_{1}}^{*},\dots,B_{t_{m}},B_{t_{m}}^{*})^{*}b_{i}$
+$=p(S_{k_{1}},S_{k_{1}}^{*},\dots,S_{k_{n}},S_{k_{n}}^{*})^{*}Wh=Sh.$
+Since $n$, $q$, $\{k_{l}\}_{1\leq l\leq n}$ and $i$ were arbitrary, and taking into account that $T$ and $S$ are both conjugate-linear, we obtain $Th=Sh$ for every $h\in\mathcal{H}_{0}$.
+Consequently, the desired equality follows by density of $\mathcal{H}_{0}$ in $L_{\mathcal{A},A}$.
 ∎
-Theorem 2.10.
-Let $N\in{\mathcal{B}}(H)$ be normal, and let $X=\{x_{i}\}_{i\in I}$ and $Y=\{y_{i}\}_{i\in I}$ be two orthogonal sets of non-zero vectors in $H$.
-Let $M\in\{N,-N\}$ and assume that for every Borel subset $\Delta\subseteq\mathbb{C}$ and all $i,j\in I$, we have
-  1. (a)
-    $\langle E_{N}(\Delta)x_{i},x_{j}\rangle=\langle E_{M}(\Delta)y_{j},y_{i}\rangle$;
-  2. (b)
-    $\langle E_{N}(\Delta)x_{i},y_{j}\rangle=\langle E_{M}(\Delta)x_{j},y_{i}\rangle$.
-Then, there exists a conjugation $C$ on $L_{N,X\cup Y}$ such that
-$CN_{X\cup Y}C=M_{X\cup Y}^{*}\quad\mbox{and}\quad Cx_{i}=y_{i}\;\;\mbox{for every }i\in I.$
-Proof.
-First, note that $L_{N,X\cup Y}=L_{M,X\cup Y}$; hence, $N_{X\cup Y}$ and $M_{X\cup Y}$ are both operators on $L_{N,X\cup Y}$, and so the statement $CN_{X\cup Y}C=M_{X\cup Y}^{*}$ makes sense.
-Additionally, without loss of generality, we may assume that $X$ and $Y$ are orthonormal subsets of $H$.
-In fact, by condition (a),
-$r_{i}:=\|x_{i}\|=\sqrt{\langle E_{N}(\mathbb{C})x_{i},x_{i}\rangle}=\sqrt{\langle E_{M}(\mathbb{C})y_{i},y_{i}\rangle}=\|y_{i}\|$
-for every $i\in I$; thus, the orthonormal subsets $\{r_{i}^{-1}x_{i}\}_{i\in I}$ and $\{r_{i}^{-1}y_{i}\}_{i\in I}$ satisfy conditions (a) and (b).
-Using condition (a), we infer from Proposition 2.9 that there exists a conjugate-linear surjective isometry $W:L_{N,X}\to L_{M,Y}$ such that
-(2.2) $Wx_{i}=y_{i}\;\mbox{and}\;Wp(N_{X},N_{X}^{*})=p(M_{Y},M_{Y}^{*})^{*}W\;\mbox{ for all $i\in I$ and $p\in\mathbb{C}[x,y]$.}$
-In particular,
-(2.3) $WN_{X}=M_{Y}^{*}W\quad\mbox{and}\quad WN_{X}^{*}=M_{Y}W.$
-Choose an arbitrary orthonormal basis $\{e_{j}\}_{j\in J}$ of $L_{N,X}$ that contains $X$ and, for each $j\in J$, put $f_{j}=We_{j}$.
-Then, since $W$ is a conjugate-linear surjective isometry, the set $\{f_{j}\}_{j\in J}$ must be an orthonormal basis of $L_{M,Y}$.
-Note that, for each $j\in J$, there exist sequences of finite subsets $I_{j,n}\subseteq I$ and polynomials $p_{j,n,k}\in\mathbb{C}[x,y]$, for $n\in\mathbb{Z}^{+}$ and $k\in I_{j,n}$, such that
-(2.4) $e_{j}=\lim_{n}\sum_{k\in I_{j,n}}p_{j,n,k}(N,N^{*})x_{k};$
-hence, by (2.2),
-(2.5) $f_{j}=We_{j}=\lim_{n}\sum_{k\in I_{j,n}}p_{j,n,k}(M,M^{*})^{*}y_{k}.$
-Here, $\mathbb{Z}^{+}$ stands for the set of positive integers.
-Claim: For all $i,j\in J$, we have $\langle e_{i},f_{j}\rangle=\langle e_{j},f_{i}\rangle$. First, note that in the light of condition (b), Lemma 2.8 gives
-(2.6) $\langle p(N,N^{*})x_{i},y_{j}\rangle=\langle p(M,M^{*})x_{j},y_{i}\rangle\quad\mbox{for all $i,j\in I$ and $p\in\mathbb{C}[x,y]$}.$
-Now, fix $i,j\in J$ and adopt the notations in (2.4) and (2.5); then
-$\left\langle e_{i},f_{j}\right\rangle$ $=\lim_{n}\left\langle\sum_{k\in I_{i,n}}p_{i,n,k}(N,N^{*})x_{k},\sum_{l\in I_{j,n}}p_{j,n,l}(M,M^{*})^{*}y_{l}\right\rangle$
-$=\lim_{n}\sum_{k\in I_{i,n}}\sum_{l\in I_{j,n}}\left\langle p_{i,n,k}(N,N^{*})x_{k},p_{j,n,l}(M,M^{*})^{*}y_{l}\right\rangle$
-$=\lim_{n}\sum_{k\in I_{i,n}}\sum_{l\in I_{j,n}}\left\langle p_{j,n,l}(M,M^{*})p_{i,n,k}(N,N^{*})x_{k},y_{l}\right\rangle.$
-Hence, using (2.6), if $M=N$, we can see that
-$\left\langle e_{i},f_{j}\right\rangle=\lim_{n}\sum_{k\in I_{i,n}}\sum_{l\in I_{j,n}}\left\langle p_{j,n,l}(N,N^{*})p_{i,n,k}(N,N^{*})x_{l},y_{k}\right\rangle,$
-and if $M=-N$, we obtain
-$\left\langle e_{i},f_{j}\right\rangle=\lim_{n}\sum_{k\in I_{i,n}}\sum_{l\in I_{j,n}}\left\langle p_{j,n,l}(N,N^{*})p_{i,n,k}(-N,-N^{*})x_{l},y_{k}\right\rangle.$
-In both cases, we have
-$\left\langle e_{i},f_{j}\right\rangle=\lim_{n}\sum_{k\in I_{i,n}}\sum_{l\in I_{j,n}}\left\langle p_{j,n,l}(N,N^{*})p_{i,n,k}(M,M^{*})x_{l},y_{k}\right\rangle.$
-Consequently, we obtain
-$\left\langle e_{i},f_{j}\right\rangle$ $=\lim_{n}\sum_{k\in I_{i,n}}\sum_{l\in I_{j,n}}\left\langle p_{j,n,l}(N,N^{*})x_{l},p_{i,n,k}(M,M^{*})^{*}y_{k}\right\rangle$
-$=\lim_{n}\left\langle\sum_{l\in I_{j,n}}p_{j,n,l}(N,N^{*})x_{l},\sum_{k\in I_{i,n}}p_{i,n,k}(M,M^{*})^{*}y_{k}\right\rangle$
-[…]
-This proves the claim.
-Note that $L_{N,X\cup Y}=\vee\left(L_{N,X}+L_{M,Y}\right)$; hence, $\{e_{j}\}_{j\in J}$ and $\{f_{j}\}_{j\in J}$ are both orthonormal subsets of $L_{N,X\cup Y}$.
-By [29, Theorem 2.1], there exists a conjugation $C$ on $L_{N,X\cup Y}$ such that $Ce_{j}=f_{j}$ for every $j\in J$.
-In particular, since $\{e_{j}\}_{j\in J}$ contains $X$, we get from (2.2) that
-$Cx_{i}=Wx_{i}=y_{i}\quad\mbox{for every $i\in I$.}$
-It remains to check that $CN_{X\cup Y}C=M_{X\cup Y}^{*}$.
-Noting that, for every $j\in J$, $Ce_{j}=f_{j}=We_{j}$ and $Cf_{j}=C^{2}e_{j}=e_{j}=W^{-1}f_{j}$, and taking into account that $C$, $W$ and $W^{-1}$ are conjugate-linear, it follows that $W=C_{1}$ and $W^{-1}=C_{2}$ where $C_{1}$ and $C_{2}$ are defined as
-$\begin{array}[]{rcl}C_{1}:L_{N,X}&\to&L_{M,Y}\\ h&\mapsto&Ch\end{array}\quad\mbox{and}\quad\begin{array}[]{rcl}C_{2}:L_{M,Y}&\to&L_{N,X}\\ h&\mapsto&Ch\end{array}.$
-Combining these facts with (2.3), we get that for every $j\in J$,
-$CM_{X\cup Y}Ce_{j}=CM_{X\cup Y}We_{j}=CM_{Y}We_{j}=CWN_{X}^{*}e_{j}=C_{2}WN_{X}^{*}e_{j}=N_{X}^{*}e_{j}=N_{X\cup Y}^{*}e_{j}$
-and
-$CN_{X\cup Y}Cf_{j}=CN_{X\cup Y}e_{j}=CN_{X}e_{j}=C_{1}N_{X}e_{j}=WN_{X}e_{j}=M_{Y}^{*}We_{j}=M_{Y}^{*}f_{j}=M_{X\cup Y}^{*}f_{j}.$
-Suppose that $M=N$; it follows that $CN_{X\cup Y}Ce_{j}=N_{X\cup Y}^{*}e_{j}$ and $CN_{X\cup Y}Cf_{j}=N_{X\cup Y}^{*}f_{j}$ for every $j\in J$.
-Hence, since $L_{N,X\cup Y}=\vee\left(L_{N,X}+L_{M,Y}\right)=\vee\left(\{e_{j},f_{j}\}_{j\in J}\right)$, we obtain $CN_{X\cup Y}C=N_{X\cup Y}^{*}$.
-On the other hand, if $M=-N$, we get $CN_{X\cup Y}Ce_{j}=-N_{X\cup Y}^{*}e_{j}$ and $CN_{X\cup Y}Cf_{j}=-N_{X\cup Y}^{*}f_{j}$ for every $j\in J$.
-Therefore, $CN_{X\cup Y}C=-N_{X\cup Y}^{*}$.
-In both cases, we have the desired result.
-This completes the proof of the theorem.
 The following lemma will be very useful throughout the paper.
 Lemma 2.15.
-Let $N\in\mathcal{B}(H)$ and $M\in\mathcal{B}(K)$ be normal operators, and let $T:H\to K$ be a bounded linear (or conjugate-linear) operator. The following statements hold:
+Let $N\in\mathcal{B}(\mathcal{H})$ and $M\in\mathcal{B}(\mathcal{K})$ be normal operators, and let $T:\mathcal{H}\to\mathcal{K}$ be a bounded linear (or conjugate-linear) operator.
+The following statements hold:
   1. (i)
@@
 (i) Suppose first that $T$ is linear and $TN=MT$, and put
-$L=\begin{bmatrix}N&0\\ 0&M\end{bmatrix}\begin{array}[]{l}H\\ K\end{array}\quad\mbox{and}\quad A=\begin{bmatrix}0&0\\ T&0\end{bmatrix}\begin{array}[]{l}H\\ K\end{array}.$
-Then, $L$ and $A$ are bounded linear operators acting on the Hilbert space $H\oplus K$.
+$L=\begin{bmatrix}N&0\\ 0&M\end{bmatrix}\begin{array}[]{l}\mathcal{H}\\ \mathcal{K}\end{array}\quad\mbox{and}\quad A=\begin{bmatrix}0&0\\ T&0\end{bmatrix}\begin{array}[]{l}\mathcal{H}\\ \mathcal{K}\end{array}.$
+Then, $L$ and $A$ are bounded linear operators acting on the Hilbert space $\mathcal{H}\oplus\mathcal{K}$.
 Furthermore, one can verify that $AL=LA$.
@@
 Fix a Borel subset $\Delta\subseteq\mathbb{C}$, and note that
-$E_{L}(\Delta)=\begin{bmatrix}E_{N}(\Delta)&0\\ 0&E_{M}(\Delta)\end{bmatrix}\begin{array}[]{l}H\\ K\end{array}.$
+$E_{L}(\Delta)=\begin{bmatrix}E_{N}(\Delta)&0\\ 0&E_{M}(\Delta)\end{bmatrix}\begin{array}[]{l}\mathcal{H}\\ \mathcal{K}\end{array}.$
 Therefore, the matrix equality $AE_{L}(\Delta)=E_{L}(\Delta)A$ implies $TE_{N}(\Delta)=E_{M}(\Delta)T$.
 Assume now that $T$ is a conjugate-linear operator such that $TN=M^{*}T$, and let $\Delta\subseteq\mathbb{C}$ be a Borel subset.
-Since $M$ is normal, there exists a conjugation $C$ on $K$ such that $CMC=M^{*}$.
-Hence, $TN=M^{*}T=(CMC)T$, and since $C^{2}=I$, we obtain
-(2.7) $(CT)N=C^{2}MCT=M(CT).$
-Clearly, $CT:H\to K$ is a bounded linear operator. By our previous result, we have
+Since $M$ is normal, there exists a conjugation $C$ on $\mathcal{K}$ such that $CMC=M^{*}$.
+Hence, $TN=M^{*}T=(CMC)T$, and since $C^{-1}=C$, we obtain
+(2.3) $(CT)N=M(CT).$
+Clearly, $CT:\mathcal{H}\to\mathcal{K}$ is a bounded linear operator. By our previous result, we have
 (2.4) $CTE_{N}(\Delta)=E_{M}(\Delta)CT.$
-By [8, Page 287, Lemma 8.7], $E_{M}(\Delta)=\chi_{\Delta}(M)\in W^{*}(M)$.
-Hence, there exists a net $\{p_{i}\}_{i}$ in $\mathbb{C}[x,y]$ such that $p_{i}(M,M^{*})$ converges to $E_{M}(\Delta)$ in the weak operator topology.
-One can easily check that $Cp_{i}(M,M^{*})C=p_{i}(M,M^{*})^{*}$.
+By [9, p.
+287, Lemma 8.7], $E_{M}(\Delta)=\chi_{\Delta}(M)\in W^{*}(M)$, the von Neumann algebra generated by $M$.
+Hence, there exists a net $\{p_{i}\}_{i}$ in $\mathcal{P}_{2}(\mathbb{C})$ such that $p_{i}(M,M^{*})$ converges to $E_{M}(\Delta)$ in the weak operator topology.
+One can easily check that $Cp_{i}(M,M^{*})C=p_{i}(M,M^{*})^{*}$ for any $i$.
 Consequently, taking the limit in the weak operator topology, we get $CE_{M}(\Delta)C=E_{M}(\Delta)^{*}=E_{M}(\Delta).$ Therefore, from (2.4), we obtain
@@
 This completes the proof. ∎
-We are now in a position to prove Theorems 2.1 and 2.3 and 2.4.
+We are now in a position to prove Theorems 2.1 and 2.4 and 2.8.
 Proof of Theorem 2.1.
-The implication (i)$\implies$(ii) follows directly from the properties of $C$ as a conjugate-linear isometry, satisfying $Cx_{i}=y_{i}$, $Cy_{i}=C^{2}x_{i}=x_{i}$ and $CT=CTC^{2}=T^{*}C$ for all $i\in I$ and $T\in\mathcal{M}$.
+(i)$\implies$(ii).
+Adopt the notations in Theorem 2.1 (i).
+Since, for every $k\in K$, $CN_{k}=M_{k}^{*}C$ and $CM_{k}=N_{k}^{*}C$, it follows by Lemma 2.15 that for all $k\in K_{0}$ and $l\in K_{1}$, $CE_{M_{k}}(\Delta_{k})=E_{N_{k}}(\Delta_{k})C$ and $CE_{N_{l}}(D_{l})=E_{M_{l}}(D_{l})C$.
+Hence,
+$\left(\prod_{k\in K_{0}}E_{N_{k}}(\Delta_{k})\right)C=C\left(\prod_{k\in K_{0}}E_{M_{k}}(\Delta_{k})\right)\quad\mbox{and}\quad\left(\prod_{l\in K_{1}}E_{M_{l}}(D_{l})\right)C=C\left(\prod_{l\in K_{1}}E_{N_{l}}(D_{l})\right).$
+Thus, taking into account the equalities $Cx_{i}=y_{i}$ and $Cy_{i}=x_{i}$, we get
+$\left\langle\left(\prod_{k\in K_{0}}E_{N_{k}}(\Delta_{k})\right)x_{i},\left(\prod_{l\in K_{1}}E_{M_{l}}(D_{l})\right)y_{j}\right\rangle$ $=\left\langle\left(\prod_{k\in K_{0}}E_{N_{k}}(\Delta_{k})\right)Cy_{i},\left(\prod_{l\in K_{1}}E_{M_{l}}(D_{l})\right)Cx_{j}\right\rangle$
+$=\left\langle C\left(\prod_{k\in K_{0}}E_{M_{k}}(\Delta_{k})\right)y_{i},C\left(\prod_{l\in K_{1}}E_{N_{l}}(D_{l})\right)x_{j}\right\rangle$
+$=\left\langle\left(\prod_{l\in K_{1}}E_{N_{l}}(D_{l})\right)x_{j},\left(\prod_{k\in K_{0}}E_{M_{k}}(\Delta_{k})\right)y_{i}\right\rangle.$
+The other equality can be obtained in a similar way.
+(ii)$\implies$(i).
+Put $K_{1}=K_{2}=K$ and let $\widetilde{K}$ be the disjoint union of $K_{1}$ and $K_{2}$.
+Let $\mathcal{A}=\{A_{k}\}_{k\in\widetilde{K}}$ and =$\mathcal{B}=\{B_{k}\}_{k\in\widetilde{K}}$ where
+$A_{k}=\begin{cases}N_{k}&\text{ if }k\in K_{1};\\ M_{k}&\text{ if }k\in K_{2},\end{cases}\quad\mbox{and}\quad B_{k}=\begin{cases}M_{k}&\text{ if }k\in K_{1};\\ N_{k}&\text{ if }k\in K_{2}.\end{cases}$
+Now, define $\widetilde{I}$ to be the disjoint union of $I_{1}$ and $I_{2}$ where $I_{1}=I_{2}=I$, and let $A=\{a_{i}\}_{i\in\widetilde{I}}$ and $B=\{b_{i}\}_{i\in\widetilde{I}}$ be such that
+$a_{i}=\begin{cases}x_{i}&\text{ if }i\in I_{1};\\ y_{i}&\text{ if }i\in I_{2},\end{cases}\quad\mbox{and}\quad b_{i}=\begin{cases}y_{i}&\text{ if }i\in I_{1};\\ x_{i}&\text{ if }i\in I_{2}.\end{cases}$
+Since, for all subsets $\{k_{l}\}_{l=1}^{4}\subseteq K$ and Borel subsets $\{\Delta_{k_{l}}\}_{l=1}^{4}$ of $\mathbb{C}$, the orthogonal projections $E_{N_{k}}(\Delta_{k_{1}})$, $E_{N_{k}}(\Delta_{k_{2}})$, $E_{M_{k}}(\Delta_{k_{3}})$ and $E_{M_{k}}(\Delta_{k_{4}})$ are mutually commuting, one can easily see that $\mathcal{A}=\{A_{k}\}_{k\in\widetilde{K}}$, =$\mathcal{B}=\{B_{k}\}_{k\in\widetilde{K}}$, $A$ and $B$ satisfy the conditions in Proposition 2.14.
+Consequently, and since $L_{\mathcal{A},A}=L_{\mathcal{B},B}=L_{\mathcal{N}\cup\mathcal{M},X\cup Y}=\mathcal{H}$, there exists a conjugate-linear surjective isometry $C:\mathcal{H}\to\mathcal{H}$ such that
+(2.5) $Ca_{i}=b_{i}\quad\mbox{and}\quad Cp(A_{k_{1}},A_{k_{1}}^{*},\dots,A_{k_{n}},A_{k_{n}}^{*})=p(B_{k_{1}},B_{k_{1}}^{*},\dots,B_{k_{n}},B_{k_{n}}^{*})^{*}C$
+for all $i\in I$, $n\in\mathbb{Z}^{+}$, $p\in\mathcal{P}_{2n}(\mathbb{C})$ and $\{k_{l}\}_{1\leq l\leq n}\subseteq\widetilde{K}$.
+Furthermore, by definition of $\mathcal{A}=\{A_{k}\}_{k\in\widetilde{K}}$, =$\mathcal{B}=\{B_{k}\}_{k\in\widetilde{K}}$, $A$ and $B$, the roles of $A_{k_{l}}$ (resp.
+$a_{i}$) and $B_{k_{l}}$ (resp.
+$b_{i}$) can be interchanged in (2.5); that is,
+(2.6) $Cb_{i}=a_{i}\quad\mbox{and}\quad Cp(B_{k_{1}},B_{k_{1}}^{*},\dots,B_{k_{n}},B_{k_{n}}^{*})=p(A_{k_{1}},A_{k_{1}}^{*},\dots,A_{k_{n}},A_{k_{n}}^{*})^{*}C.$
+In particular, we have
+$Cx_{i}=y_{i}\quad\mbox{and}\quad CN_{k}=M_{k}^{*}C\quad\mbox{for every $k\in K$}.$
+Now, we only need to show that $C$ is a conjugation.
+Let us first show that $C^{2}=I_{\mathcal{H}}$.
+Since $\mathcal{H}=L_{\mathcal{A},A}$, it suffices to show that $C^{2}h=h$ for every vector $h$ of the form $h=p(A_{k_{1}},A_{k_{1}}^{*},\dots,A_{k_{n}},A_{k_{n}}^{*})a_{i}$ where $i\in I$, $n\in\mathbb{Z}^{+}$, $p\in\mathcal{P}_{2n}(\mathbb{C})$ and $\{k_{l}\}_{1\leq l\leq n}\subseteq\widetilde{K}$.
+Let $h$ be such vector; then, (2.5) and (2.6) imply
+$C^{2}h=C^{2}p(A_{k_{1}},A_{k_{1}}^{*},\dots,A_{k_{n}},A_{k_{n}}^{*})a_{i}$ $=Cp(B_{k_{1}},B_{k_{1}}^{*},\dots,B_{k_{n}},B_{k_{n}}^{*})^{*}b_{i}$
+$=p(A_{k_{1}},A_{k_{1}}^{*},\dots,A_{k_{n}},A_{k_{n}}^{*})a_{i}=h.$
+Finally, the isometric property in Definition 1.1 (ii) can be obtain from (2.7) since $C^{\#}=C^{-1}=C$. ∎
+Proof of Theorem 2.4.
+The implication (i)$\implies$(ii) is obvious.
 (ii)$\implies$(iii).
 Let $V_{0}:{\rm ker}(V)^{\perp}\to{\rm ran}(V)$ be the invertible conjugate-linear isometry given by $V_{0}h=Vh$.
-Define the conjugate-linear map $V^{\#}:H\to H$ by $V^{\#}h=V_{0}^{-1}h$ if $h\in{\rm ran}(V)$ and $V^{\#}h=0$ if $h\in{\rm ran}(V)^{\perp}$.
+Define the conjugate-linear map $V^{\#}:\mathcal{H}\to\mathcal{H}$ by $V^{\#}h=V_{0}^{-1}h$ if $h\in{\rm ran}(V)$ and $V^{\#}h=0$ if $h\in{\rm ran}(V)^{\perp}$.
 Then, $V^{\#}V$ is the orthogonal projection onto ${\rm ker}(V)^{\perp}$.
@@
 Thus,
-(2.9) $\langle Vh,k\rangle=\langle V^{\#}k,h\rangle\quad\mbox{for all }h,k\in H.$
-Now, let $P\in\mathcal{M}$ be an orthogonal projection. Using the above relation, for all $i,j\in I$, we obtain
+(2.7) $\langle Vh,k\rangle=\langle V^{\#}k,h\rangle\quad\mbox{for all }h,k\in\mathcal{H}.$
+Let $K_{0}\subseteq K$ be a finite subset, and let $\{\Delta_{k}\}_{k\in K_{0}}$ be Borel subsets of $\mathbb{C}$.
+In view of Lemma 2.15, for all $k\in K_{0}$, we have $VE_{N_{k}}(\Delta_{k})=E_{N_{k}}(\Delta_{k})V$.
+Hence, $V\left(\prod_{k\in K_{0}}E_{N_{k}}(\Delta_{k})\right)=\left(\prod_{k\in K_{0}}E_{N_{k}}(\Delta_{k})\right)V$.
+Put $P=\prod_{k\in K_{0}}E_{N_{k}}(\Delta_{k})$.
+Thus,
 (2.8) $\begin{split}\langle Py_{j},y_{i}\rangle=\langle PVx_{j},Vx_{i}\rangle&=\langle VPx_{j},Vx_{i}\rangle=\langle V^{\#}Vx_{i},Px_{j}\rangle\\ &=\langle x_{i},Px_{j}\rangle=\langle Px_{i},x_{j}\rangle,\end{split}$
@@
 (2.9) $\begin{split}\langle Px_{j},y_{i}\rangle=\langle PVy_{j},Vx_{i}\rangle&=\langle VPy_{j},Vx_{i}\rangle=\langle V^{\#}Vx_{i},Py_{j}\rangle\\ &=\langle x_{i},Py_{j}\rangle=\langle Px_{i},y_{j}\rangle.\end{split}$
-(iii)$\implies$(ii).
-According to [10, Page 52, Lemma II.2.8], there exists a self-adjoint operator $A\in{\mathcal{B}}(H)$ such that $\mathcal{M}=W^{*}(A)$.
-It suffices to show that there exists a conjugation $C$ on $H$ such that $CAC=A$ and $Cx_{i}=y_{i}$ for every $i\in I$.
-In fact, any polynomial in $A$ is $C$-symmetric; moreover, since the class of $C$-symmetric operators is closed in the weak operator topology, every operator in $W^{*}(A)$, and thus in $\mathcal{M}$, is $C$-symmetric.
-Let $\Delta\subseteq\mathbb{C}$ be a Borel subset.
-Then, by [8, Page 287, Lemma 8.7], the orthogonal projection $E_{A}(\Delta)$ belongs to $\mathcal{M}$.
-Therefore, by hypothesis, for all $i,j\in I$, we have $\langle E_{A}(\Delta)x_{i},x_{j}\rangle=\langle E_{A}(\Delta)y_{j},y_{i}\rangle$ and $\langle E_{A}(\Delta)x_{i},y_{j}\rangle=\langle E_{A}(\Delta)x_{j},y_{i}\rangle$.
-For convenience, put $Z=\{x_{i},y_{i}\}_{i\in I}$, $H_{0}=L_{A,Z}$ and $H_{1}=L_{A,Z}^{\perp}$.
-Then, $H_{0}$ and $H_{1}$ are reducing subspaces of $A$ and $Z\subset H_{0}$.
-According to Theorem 2.10, there exists a conjugation $C_{0}$ on $H_{0}$ such that $C_{0}A_{|H_{0}}C=A_{0}$ and $C_{0}x_{i}=y_{i}$ for every $i\in I$.
-Since $A_{|H_{1}}$ is normal, there exists a conjugation $C_{1}$ on $H_{1}$ such that $C_{1}A_{|H_{1}}C_{1}=A_{|H_{1}}$.
-Clearly, $C=C_{0}\oplus C_{1}$ is a conjugation on $H$; additionally, $CAC=A$ and $Cx_{i}=C_{0}x_{i}=y_{i}$ for every $i\in I$.
-∎
-Proof of Theorem 2.3.
-The implication (i)$\implies$(ii) is obvious, and the implication (ii)$\implies$(iii) follows a similar argument as (i)$\implies$(ii) in Theorem 2.1.
-To show that (iii)$\implies$(iv), first note that by Lemma 2.11, for all Borel subsets $\Delta\subseteq\mathbb{C}$ and $k\in K$, we have $VE_{N_{k}}(\Delta)=E_{N_{k}}(\Delta)V$.
-Hence, for any finite subset $K_{0}\subseteq K$ and Borel subsets $\{\Delta_{k}\}_{k\in K_{0}}$ of $\mathbb{C}$, we have $V\left(\prod_{k\in K_{0}}E_{N_{k}}(\Delta_{k})\right)=\left(\prod_{k\in K_{0}}E_{N_{k}}(\Delta_{k})\right)V$.
-Put $P=\prod_{k\in K_{0}}E_{N_{k}}(\Delta_{k})$; the remainder of the proof follows similarly to (2.10) and (2.11).
-(iv)$\implies$(i).
-For convenience, put $\mathcal{M}=W^{*}(N_{k}:k\in K)$.
-By the Fuglede Theorem, we have $N_{k}N_{l}^{*}=N_{l}^{*}N_{k}$ for any $k,l\in K$.
-Therefore, it is not hard to see that $\mathcal{M}$ is an abelian von Neumann algebra.
-Hence, by Theorem 2.1, we need to show that for all orthogonal projections $P\in\mathcal{M}$ and $i,j\in I$,
-(2.12) $\langle Px_{i},x_{j}\rangle=\langle Py_{j},y_{i}\rangle\quad\mbox{and}\quad\langle Px_{i},y_{j}\rangle=\langle Px_{j},y_{i}\rangle.$
-First, we show that $\mathcal{M}={\rm cl-WOT}\mathcal{M}_{0}$; i.e., the closure in the weak operator topology of
-$\mathcal{M}_{0}={\rm span}\left\{\prod_{k\in K_{0}}E_{N_{k}}(\Delta_{k}):K_{0}\subseteq K\mbox{ is finite}\mbox{ and }\Delta_{k}\mbox{ is a Borel subset of }\mathbb{C}\right\}.$
-By [8, Page 287, Lemma 8.7], any spectral projection $E_{N_{k}}(\Delta)=\chi_{\Delta}(N_{k})\in W^{*}(N_{k})\subseteq\mathcal{M}$.
-Hence, $\mathcal{M}_{0}\subseteq\mathcal{M}$, and consequently ${\rm cl-WOT}\mathcal{M}_{0}\subseteq\mathcal{M}$.
-Let us show the other containment.
-According to spectral theory, for every $k\in K$, we have
-$N_{k}\in\bigvee\big{\{}E_{N_{k}}(\Delta):\Delta\mbox{ is a Borel subset of }\mathbb{C}\big{\}}.$
-This implies that $N_{k}\in{\rm cl-WOT}\mathcal{M}_{0}$ for every $k\in K$.
-It is easy to see that ${\rm cl-WOT}\mathcal{M}_{0}$ is a self-adjoint subalgebra of ${\mathcal{B}}(H)$ that contains the identity.
-Moreover, since, by the Fuglede Theorem, the spectral measures $\{E_{N_{k}}\}_{k\in K}$ mutually commute and $E_{N_{k}}(\Delta)E_{N_{k}}(D)=E_{N_{k}}(\Delta\cap D)$ for any Borel subsetes $\Delta$ and $D$ of $\mathbb{C}$, one can easily show that $\mathcal{M}_{0}$, and thus ${\rm cl-WOT}\mathcal{M}_{0}$, is also multiplicative.
-Hence, ${\rm cl-WOT}\mathcal{M}_{0}$ is a von Neumann algebra containing $\{N_{k}\}_{k\in K}$.
-Therefore, by definition of $\mathcal{M}$, we must have $\mathcal{M}\subseteq{\rm cl-WOT}\mathcal{M}_{0}$.
-Now, by hypothesis, any $T\in\mathcal{M}_{0}$ satisfies the equalities (2.12).
-Then, by taking limits, in the weak operator topology, of operators in $\mathcal{M}_{0}$, we obtain that (2.12) is true for every $T\in\mathcal{M}$; in particular, for orthogonal projections in $\mathcal{M}$.
-∎
+(iii)$\implies$(i).
+Assume that $M_{k}=N_{k}$ for every $k\in K$.
+For convenience, put $\mathcal{L}_{0}=L_{\mathcal{N},X\cup Y}$ and $\mathcal{L}_{1}=\mathcal{L}_{0}^{\perp}$.
+Then, $\mathcal{L}_{0}$ reduces $N_{k}$ for every $k\in K$; additionally, $x_{i},y_{i}\in\mathcal{L}_{0}$ for all $i\in I$.
+It is not hard to see that $\left\{{N_{k}}_{|\mathcal{L}_{0}}\right\}_{k\in K}$, $\left\{{M_{k}}_{|\mathcal{L}_{0}}\right\}_{k\in K}$, $X$ and $Y$ satisfy assertion (ii) of Theorem 2.1.
+Thus, by the same theorem, there exists a conjugation $C_{0}$ on $\mathcal{L}_{0}$ such that $C_{0}{N_{k}}_{|\mathcal{L}_{0}}C_{0}=\left({N_{k}}_{|\mathcal{L}_{0}}\right)^{*}$ and $C_{0}x_{i}=y_{i}$ for all $k\in K$ and $i\in I$.
+On the other hand, since the normal operators ${N_{k}}_{|\mathcal{L}_{1}}$ are mutually commuting, they generate an abelian von Neumann algebra in $\mathcal{B}(\mathcal{L}_{1})$.
+Hence, according to [11, p.
+52, Lemma II.2.8], there exists a self-adjoint operator $A\in\mathcal{B}(\mathcal{L}_{1})$ such that ${N_{k}}_{|\mathcal{L}_{1}}\in W^{*}(A)$ for every $k\in K$.
+By self-adjointness of $A$, there exists a conjugation $C_{1}$ on $\mathcal{L}_{1}$ such that $C_{1}AC_{1}=A$.
+Thus, any polynomial in $A$ is $C_{1}$-symmetric; moreover, since the class of $C_{1}$-symmetric operators is closed in the weak operator topology, we get
+(2.10) $C_{1}{N_{k}}_{|\mathcal{L}_{1}}C_{1}=\left({N_{k}}_{|\mathcal{L}_{1}}\right)^{*}\quad\mbox{for every $k\in K$}.$
+Therefore, one can readily see that $C=C_{0}\oplus C_{1}$ is a conjugation on $\mathcal{H}=\mathcal{L}_{0}\oplus\mathcal{L}_{1}$ that satisfies all desired equalities. ∎
 Proof of Theorem 2.8.
 The implication (ii)$\implies$(iii) is obvious.
-The implication (iv)$\implies$(i) follows from Theorem 2.3 (iv)$\implies$(i).
-In fact, for any finite subset $K_{0}\subseteq K$ and Borel subsets $\{\Delta_{k}\}_{k\in K_{0}}$ of $\mathbb{C}$, the orthogonal projections $\{E_{N_{k}}(\Delta_{k})\}_{k\in K_{0}}$ are mutually commuting.
-Thus, their product $P$ is also an orthogonal projection; which gives
-$\langle Px,x\rangle=\langle P^{2}x,x\rangle=\langle Px,Px\rangle=\|Px\|^{2}=\|Py\|^{2}=\langle Py,y\rangle.$
 (iii)$\implies$(iv).
@@
 $\|Py\|^{2}=\langle Py,y\rangle=\langle PVx,Vx\rangle=\langle Px,V^{*}Vx\rangle=\langle Px,x\rangle=\|Px\|^{2}.$
-(i)$\implies$(v). Clearly, $\|x\|=\|Cx\|=\|y\|$. It is easy to check that $C(x\otimes y)C=(Cx)\otimes(Cy)$; hence, for any $k\in K$ and $\lambda\in\mathbb{C}$, we have
+The implication (iv)$\implies$(i) follows from Theorem 2.4 (iii)$\implies$(i).
+In fact, for any finite subset $K_{0}\subseteq K$ and Borel subsets $\{\Delta_{k}\}_{k\in K_{0}}$ of $\mathbb{C}$, the orthogonal projections $\{E_{N_{k}}(\Delta_{k})\}_{k\in K_{0}}$ are mutually commuting.
+Thus, their product $P$ is also an orthogonal projection; which gives
+$\langle Px,x\rangle=\langle P^{2}x,x\rangle=\langle Px,Px\rangle=\|Px\|^{2}=\|Py\|^{2}=\langle Py,y\rangle.$
+(i)$\implies$(v). In view of (1.1), for any $k\in K$ and $\lambda\in\mathbb{C}$, we have
 $C\left(N_{k}+\lambda x\otimes y\right)C=CN_{k}C+\overline{\lambda}y\otimes x=N_{k}^{*}+\overline{\lambda}(x\otimes y)^{*}=\left(N_{k}+\lambda x\otimes y\right)^{*};$
@@
 (v)$\implies$(ii).
-Since the class of $C$-symmetric operators forms a subspace of ${\mathcal{B}}(H)$, we obtain, for every $k\in K$, $CN_{k}C=N_{k}^{*}$ and $C(x\otimes y)C=(x\otimes y)^{*}=y\otimes x$.
+Since the class of $C$-symmetric operators forms a subspace of ${\mathcal{B}}(\mathcal{H})$, we obtain, for every $k\in K$, $CN_{k}C=N_{k}^{*}$ and $C(x\otimes y)C=(x\otimes y)^{*}=y\otimes x$.
 Given that $x$ and $y$ have the same norm, the latter equality implies the existence of a unimodular $\alpha\in\mathbb{C}$ such that $Cx=\alpha y$.
-Applying Theorem 2.3 (iv)$\implies$(ii) to $\{x_{i}\}_{i}=\{y_{i}\}_{i}=\{x\}$, we obtain a conjugation $J$ on $H$ such that $JN_{k}J=N_{k}^{*}$ for every $k\in K$ and $Jx=x$.
-Put $U=\overline{\alpha}CJ$; then, since $C$ and $J$ are both surjective isometries and $\alpha$ is unimodular, the linear operator $U:H\to H$ must be unitary.
+Applying Theorem 2.4 (iii)$\implies$(i) to the particular case $\{x_{i}\}_{i}=\{y_{i}\}_{i}=\{x\}$, we obtain a conjugation $J$ on $\mathcal{H}$ such that $JN_{k}J=N_{k}^{*}$ for every $k\in K$ and $Jx=x$.
+Put $U=\overline{\alpha}CJ$; then, since $C$ and $J$ are both surjective isometries and $\alpha$ is unimodular, the linear operator $U:\mathcal{H}\to\mathcal{H}$ must be unitary.
 Moreover, we have
@@
 3. Interpolation theorems for conjugations : C-skew symmetric versions
-A conjugate-linear operator $V$ on $H$ is said to be anti-unitary if it is a surjective isometry. Clearly, every conjugation on $H$ is anti-unitary.
+A conjugate-linear operator $V$ on $\mathcal{H}$ is said to be anti-unitary if it is a surjective isometry. Clearly, every conjugation on $\mathcal{H}$ is anti-unitary.
 For two operators $A$ and $B$ on Hilbert spaces, we denote $A\cong B$ if there exists an isometric isomorphism $U$ between the underlying Hilbert spaces such that $UAU^{-1}=B$; or equivalently, $UA=BU$.
 It is worth mentioning that both complex symmetry and skew symmetry are preserved under the equivalence relation $\cong$.
-For $k\in\{1,2,\dots,\infty\}$, where $\infty=\aleph_{0}$, $H^{(k)}$ denotes the direct sum of $H$ with itself $k$ times.
-If $T\in{\mathcal{B}}(H)$, then $T^{(k)}\in\mathcal{B}(H^{(k)})$ is the direct sum of $T$ with itself $k$ times.
+For $k\in\{1,2,\dots,\infty\}$, where $\infty=\aleph_{0}$, let $\mathcal{H}^{(k)}$ denote the direct sum of $\mathcal{H}$ with itself $k$ times.
+Accordingly, we denote by $T^{(k)}\in\mathcal{B}(\mathcal{H}^{(k)})$ the direct sum of $T\in{\mathcal{B}}(\mathcal{H})$.
 3.1. Main results
 Theorem 3.1.
-Let $N\in{\mathcal{B}}(H)$ be normal, and let $Z=X\cup Y$ where $X=\{x_{i}\}_{i\in I}$ and $Y=\{y_{i}\}_{i\in I}$ are orthogonal sets of non-zero vectors in $H$.
-Then, the following statements are equivalent:
+The following statements are equivalent:
   1. (i)
-    There exists a conjugation $C$ on $H$ such that $CNC=-N^{*}$ and $Cx_{i}=y_{i}$ for every $i\in I$.
+    There exists a conjugation $C$ on $\mathcal{H}$ such that $CN_{k}C=-N_{k}^{*}$ and $Cx_{i}=y_{i}$ for all $k\in K$ and $i\in I$.
   2. (ii)
-    There exists an anti-unitary operator $V$ on $H$ such that $VNV^{-1}=-N^{*}$, $Vx_{i}=y_{i}$ and $Vy_{i}=x_{i}$ for every $i\in I$.
+    There exists an anti-unitary operator $V$ on $\mathcal{H}$ such that $VN_{k}V^{-1}=-N_{k}^{*}$, $Vx_{i}=y_{i}$ and $Vy_{i}=x_{i}$ for all $k\in K$ and $i\in I$.
   3. (iii)
-    $N_{|L_{N,Z}^{\perp}}$ is skew-symmetric and, for all Borel subsets $\Delta\subseteq\mathbb{C}$ and $i,j\in I$, we have
-    $\langle E_{N}(\Delta)x_{i},x_{j}\rangle=\langle E_{N}(-\Delta)y_{j},y_{i}\rangle\quad\mbox{and}\quad\langle E_{N}(\Delta)x_{i},y_{j}\rangle=\langle E_{N}(-\Delta)x_{j},y_{i}\rangle.$
+    There exists an anti-unitary operator $V_{1}$ on $\mathcal{L}_{1}:=\left(L_{\mathcal{N},X\cup Y}\right)^{\perp}$ such that
+    $V_{1}{N_{k}}_{|\mathcal{L}_{1}}V_{1}^{-1}=-\left({N_{k}}_{|\mathcal{L}_{1}}\right)^{*}\quad\mbox{for every $k\in K$},$
+    and for all finite subsets $K_{0}\subseteq K$, Borel subsets $\{\Delta_{k}\}_{k\in K_{0}}$ of $\mathbb{C}$, and $i,j\in I$ we have
+    $\left\langle\left(\prod_{k\in K_{0}}E_{N_{k}}(\Delta_{k})\right)x_{i},x_{j}\right\rangle=\left\langle\left(\prod_{k\in K_{0}}E_{N_{k}}(-\Delta_{k})\right)y_{j},y_{i}\right\rangle$
+    and
+    $\left\langle\left(\prod_{k\in K_{0}}E_{N_{k}}(\Delta_{k})\right)x_{i},y_{j}\right\rangle=\left\langle\left(\prod_{k\in K_{0}}E_{N_{k}}(-\Delta_{k})\right)x_{j},y_{i}\right\rangle.$
 Before stating the second main result of this section, we recall some facts about normal operators on separable complex Hilbert spaces.
 Given a finite measure with compact support $\mu$ on $\mathbb{C}$, denote by $M_{\mu}$ the operator defined on $L^{2}(\mu)$ by $[M_{\mu}f](z)=zf(z)$ for all $f\in L^{2}(\mu)$.
-According to spectral multiplicity theory (see, for instance, [8, Page 298, Theorem 10.16]), each normal operator $N$ acting on $H$ is unitarily equivalent to
+According to spectral multiplicity theory (see, for instance, [9, p.
+298, Theorem 10.16]), each normal operator $N$ acting on $\mathcal{H}$ is unitarily equivalent to
 (3.1) $M_{\mu_{\infty}}^{(\infty)}\oplus\bigoplus_{1\leq i<\infty}M_{\mu_{i}}^{(i)}$
@@
 (3.2) $\mu_{i}(\Delta)=0\;\;\Longleftrightarrow\;\;\nu_{i}(\Delta)=0\quad\mbox{for every Borel subset }\Delta\subseteq\mathbb{C}.$
-In [17, Theorem 1.11], Li et al. proved that $N$ is skew-symmetric if and only if the sequence $\{\mu_{i}\}_{1\leq i\leq\infty}$ can be chosen so that, for every $i\in\{1,2,\dots,\infty\}$,
+In [18, Theorem 1.11], Li et al. proved that $N$ is skew-symmetric if and only if the sequence $\{\mu_{i}\}_{1\leq i\leq\infty}$ can be chosen so that, for every $i\in\{1,2,\dots,\infty\}$,
 (3.3) $\mu_{i}(\Delta)=\mu_{i}(-\Delta)\quad\mbox{for every Borel subset }\Delta\subseteq\mathbb{C}.$
-Given a normal operator $N\in{\mathcal{B}}(H)$, we write $m(N)<\infty$ if the measure $\mu_{\infty}$ in (3.1) is zero; in other words, $M_{\mu_{\infty}}^{(\infty)}$ can be omitted from the decomposition (3.1).
+Given a normal operator $N\in{\mathcal{B}}(\mathcal{H})$, we write $m(N)<\infty$ if the measure $\mu_{\infty}$ in (3.1) is zero; in other words, $M_{\mu_{\infty}}^{(\infty)}$ can be omitted from the decomposition (3.1).
 This definition makes sense due to the uniqueness property described in (3.2).
 Theorem 3.2.
-Let $N\in{\mathcal{B}}(H)$ be a skew-symmetric normal operator, and let $Z$ be the set in Theorem 3.1 such that $m\left({N_{Z}}_{|{\rm ker}(N_{Z})^{\perp}}\right)<\infty$.
-Then, the following statements are equivalent:
+Let $N\in{\mathcal{B}}(\mathcal{H})$ be a skew-symmetric normal operator, $N_{0}$ be the restriction of $N$ to $L_{\{N\},X\cup Y}$, and assume that $m\left({N_{0}}_{|{\rm ker}(N_{0})^{\perp}}\right)<\infty$.
+The following statements are equivalent:
   1. (i)
-    There exists a conjugation $C$ on $H$ such that $CNC=-N^{*}$ and $Cx_{i}=y_{i}$ for every $i\in I$.
+    There exists a conjugation $C$ on $\mathcal{H}$ such that $CNC=-N^{*}$ and $Cx_{i}=y_{i}$ for every $i\in I$.
   2. (ii)
@@
     $\langle E_{N}(\Delta)x_{i},x_{j}\rangle=\langle E_{N}(-\Delta)y_{j},y_{i}\rangle\quad\mbox{and}\quad\langle E_{N}(\Delta)x_{i},y_{j}\rangle=\langle E_{N}(-\Delta)x_{j},y_{i}\rangle.$
-In particular, if $I$ is finite, then $m\left({N_{Z}}_{|{\rm ker}(N_{Z})^{\perp}}\right)<\infty$ and the previous statements are equivalent.
+In particular, if $I$ is finite, then $m\left({N_{0}}_{|{\rm ker}(N_{0})^{\perp}}\right)<\infty$ and the previous statements are equivalent.
+The proofs of Theorems 3.1 and 3.2 will be given at the end of this section.
 The reader can see from Theorem 3.1 that the implication (i)$\implies$(ii), in the previous theorem, is true without the assumption that $m\left({N_{Z}}_{|{\rm ker}(N_{Z})^{\perp}}\right)<\infty$.
@@
 Example 3.3.
-Suppose that $H=K_{1}\oplus K_{2}$, where $K_{1}$ and $K_{2}$ are infinite-dimensional separable complex Hilbert spaces, and let $\{e_{i}\}_{i\geq 1}$ and $\{f_{i}\}_{i\geq 1}$ be orthonormal bases of $K_{1}$ and $K_{2}$, respectively.
-For each $i\geq 1$, put $x_{i}=e_{i}\oplus 0$ and $y_{i}=0\oplus f_{i+1}$.
-Let $N$ be the normal operator on $H$ given by $N=I_{K_{1}}\oplus-I_{K_{2}}$.
+Suppose that $\mathcal{H}=\mathcal{K}_{1}\oplus\mathcal{K}_{2}$, where $\mathcal{K}_{1}$ and $\mathcal{K}_{2}$ are infinite-dimensional separable complex Hilbert spaces, and let $\{e_{i}\}_{i\geq 1}$ and $\{f_{i}\}_{i\geq 1}$ be orthonormal bases of $\mathcal{K}_{1}$ and $\mathcal{K}_{2}$, respectively.
+For each $i\geq 1$, put $x_{i}=e_{i}\oplus 0$ $y_{i}=0\oplus f_{i+1}$.
+Let $N$ be the normal operator on $\mathcal{H}$ given by $N=I_{\mathcal{K}_{1}}\oplus-I_{\mathcal{K}_{2}}$.
 Clearly, $N\cong-N$; hence, by Lemma 3.7, $N$ is skew-symmetric.
-For convenience, put $H_{1}=\vee\{x_{i}:i\geq 1\}$, $H_{2}=\vee\{y_{i}:i\geq 1\}$ and $H_{3}=\vee\{f_{1}\}$. Then, $H=H_{1}\oplus H_{2}\oplus H_{3}$ and $L_{N,Z}=H_{1}\oplus H_{2}$. Furthermore,
-${N_{Z}}_{|{\rm ker}(N_{Z})^{\perp}}={N_{Z}}=\begin{bmatrix}I_{H_{1}}&0\\ 0&-I_{H_{2}}\end{bmatrix}\begin{array}[]{l}H_{1}\\ H_{2}\end{array}\cong M_{\delta_{1}}^{(\infty)}\oplus M_{\delta_{-1}}^{(\infty)}\cong\left(M_{\delta_{1}}\oplus M_{\delta_{-1}}\right)^{(\infty)}\cong M_{\mu}^{(\infty)}$
+For convenience, put $\mathcal{H}_{1}=\vee\{x_{i}:i\geq 1\}$, $\mathcal{H}_{2}=\vee\{y_{i}:i\geq 1\}$ and $\mathcal{H}_{3}=\vee\{0\oplus f_{1}\}$.
+Then, $\mathcal{H}=\mathcal{H}_{1}\oplus\mathcal{H}_{2}\oplus\mathcal{H}_{3}$ and $L_{\{N\},X\cup Y}=\mathcal{H}_{1}\oplus\mathcal{H}_{2}$.
+Furthermore,
+${N_{0}}_{|{\rm ker}(N_{0})^{\perp}}={N_{0}}=\begin{bmatrix}I_{\mathcal{H}_{1}}&0\\ 0&-I_{\mathcal{H}_{2}}\end{bmatrix}\begin{array}[]{l}\mathcal{H}_{1}\\ \mathcal{H}_{2}\end{array}\cong M_{\delta_{1}}^{(\infty)}\oplus M_{\delta_{-1}}^{(\infty)}\cong\left(M_{\delta_{1}}\oplus M_{\delta_{-1}}\right)^{(\infty)}\cong M_{\mu}^{(\infty)}$
 where $\mu=\delta_{1}+\delta_{-1}$ and $\delta_{a}$ denotes the Dirac measure, at $a\in\mathbb{C}$, on Borel subsets of $\mathbb{C}$.
-Thus, $m\left({N_{Z}}_{|{\rm ker}(N_{Z})^{\perp}}\right)\nless\infty$.
+Thus, $m\left({N_{0}}_{|{\rm ker}(N_{0})^{\perp}}\right)\nless\infty$.
 Let us check that $N$ satisfies the conditions in Theorem 3.2 (ii).
@@
   2. (b)
-    $E_{N}(\Delta)=E_{N}(-\Delta)=I_{H}$;
+    $E_{N}(\Delta)=E_{N}(-\Delta)=I_{\mathcal{H}}$;
   3. (c)
-    $E_{N}(\Delta)=I_{K_{1}}\oplus 0$ and $E_{N}(-\Delta)=0\oplus I_{K_{2}}$;
+    $E_{N}(\Delta)=I_{\mathcal{K}_{1}}\oplus 0$ and $E_{N}(-\Delta)=0\oplus I_{\mathcal{K}_{2}}$;
   4. (d)
-    $E_{N}(\Delta)=0\oplus I_{K_{2}}$ and $E_{N}(-\Delta)=I_{K_{1}}\oplus 0$.
+    $E_{N}(\Delta)=0\oplus I_{\mathcal{K}_{2}}$ and $E_{N}(-\Delta)=I_{\mathcal{K}_{1}}\oplus 0$.
 In all cases, we have
 $\langle E_{N}(\Delta)x_{i},x_{j}\rangle=\langle E_{N}(-\Delta)y_{j},y_{i}\rangle\quad\mbox{and}\quad\langle E_{N}(\Delta)x_{i},y_{j}\rangle=\langle E_{N}(-\Delta)x_{j},y_{i}\rangle.$
-Now, let us show that there exists no conjugation $C$ on $H$ satisfying the first assertion of Theorem 3.2. Suppose for contradiction that such a conjugation exists. Then,
-$CH_{1}=C\vee\{x_{i}:i\geq 1\}=\vee\{Cx_{i}:i\geq 1\}=\vee\{y_{i}:i\geq 1\}=H_{2},$
-and since $C^{2}=I_{H}$, we also have $CH_{2}=H_{1}$. As $C$ is a surjective isometry, it follows that $CH_{3}=C(H_{1}+H_{2})^{\perp}=(H_{1}+H_{2})^{\perp}=H_{3}$. Therefore, $C$ takes the form
-$C=\begin{bmatrix}0&*&0\\ *&0&0\\ 0&0&C_{3}\end{bmatrix}\begin{array}[]{l}H_{1}\\ H_{2}\\ H_{3}\end{array}$
-where $C_{3}$ is a conjugation on $H_{3}$.
-Finally, since $H_{3}$ reduces $N$ and $CNC=-N^{*}$, we obtain $-I_{H_{3}}=C_{3}(-I_{H_{3}})C_{3}=C_{3}N_{|H_{3}}C_{3}=-N_{|H_{3}}^{*}=I_{H_{3}}$; a contradiction.
-When $X=\{x\}$, $Y=\{y\}$, and $\{x,y\}$ are linearly independent, we can add the following equivalent perturbation-type statement to Theorem 3.1.
+Now, let us show that there exists no conjugation $C$ on $\mathcal{H}$ satisfying the first assertion of Theorem 3.2. Suppose for contradiction that such a conjugation exists. Then,
+$C\mathcal{H}_{1}=C\vee\{x_{i}:i\geq 1\}=\vee\{Cx_{i}:i\geq 1\}=\vee\{y_{i}:i\geq 1\}=\mathcal{H}_{2},$
+and since $C^{-1}=C$, we also have $C\mathcal{H}_{2}=\mathcal{H}_{1}$, and hence $C(\mathcal{H}_{1}+\mathcal{H}_{2})=\mathcal{H}_{1}+\mathcal{H}_{2}$.
+As $C$ is a surjective isometry, it follows that $C\mathcal{H}_{3}=C(\mathcal{H}_{1}+\mathcal{H}_{2})^{\perp}=(\mathcal{H}_{1}+\mathcal{H}_{2})^{\perp}=\mathcal{H}_{3}$.
+Therefore, $C$ takes the form
+$C=\begin{bmatrix}0&*&0\\ *&0&0\\ 0&0&C_{3}\end{bmatrix}\begin{array}[]{l}\mathcal{H}_{1}\\ \mathcal{H}_{2}\\ \mathcal{H}_{3}\end{array}$
+where $C_{3}$ is a conjugation on $\mathcal{H}_{3}$.
+Finally, since $\mathcal{H}_{3}$ reduces $N$ and $CNC=-N^{*}$, we obtain $-I_{\mathcal{H}_{3}}=C_{3}(-I_{\mathcal{H}_{3}})C_{3}=C_{3}N_{|\mathcal{H}_{3}}C_{3}=-N_{|\mathcal{H}_{3}}^{*}=I_{\mathcal{H}_{3}}$; a contradiction.
+When $X=\{x\}$, $Y=\{y\}$, and $\{x,y\}$ are linearly independent, we can add the following statement to Theorem 3.1.
 Theorem 3.4.
-Let $N\in{\mathcal{B}}(H)$ be normal, and let $\{x,y\}$ be linearly independent vectors in $H$. The following statements are equivalent:
+Let $\{x,y\}$ be linearly independent vectors in $\mathcal{H}$. The following statements are equivalent:
   1. (i)
-    There exists a conjugation $C$ on $H$ satisfying $CNC=-N^{*}$ and $Cx=y$.
+    There exists a conjugation $C$ on $\mathcal{H}$ satisfying $CN_{k}C=-N_{k}^{*}$, for every $k\in K$, and $Cx=y$.
   2. (ii)
-    There exists a conjugation $C$ on $H$ such that $N+\lambda(x\otimes x-y\otimes y)$ is $C$-skew symmetric for every $\lambda\in\mathbb{C}$.
+    There exists a conjugation $C$ on $\mathcal{H}$ such that $N_{k}+\lambda(x\otimes x-y\otimes y)$ is $C$-skew symmetric for all $k\in K$ and $\lambda\in\mathbb{C}$.
 The proof of the previous theorem does not require preliminary results and can be presented here.
 Proof of Theorem 3.4.
-(i)$\implies$(ii). Let $C$ be a conjugation satisfying (i). Then, for every $\lambda\in\mathbb{C}$, we have
-$C\big{(}N+\lambda(x\otimes x-y\otimes y)\big{)}C$ $=CNC+\overline{\lambda}\big{(}(Cx)\otimes(Cx)-(Cy)\otimes(Cy)\big{)}$
-$=-N^{*}+\overline{\lambda}(y\otimes y-x\otimes x)$
-$=-\left(N+\lambda(x\otimes x-y\otimes y)\right)^{*}.$
+(i)$\implies$(ii). Let $C$ be a conjugation satisfying (i). Then, for all $k\in K$ and $\lambda\in\mathbb{C}$, we have
+$C\big{(}N_{k}+\lambda(x\otimes x-y\otimes y)\big{)}C$ $=CN_{k}C+\overline{\lambda}\big{(}(Cx)\otimes(Cx)-(Cy)\otimes(Cy)\big{)}$
+$=-N_{k}^{*}+\overline{\lambda}(y\otimes y-x\otimes x)$
+$=-\left(N_{k}+\lambda(x\otimes x-y\otimes y)\right)^{*}.$
 (ii)$\implies$(i).
 For convenience, put $T=x\otimes x-y\otimes y$.
-By the fact that the class of $C$-skew symmetric operators is a subspace of ${\mathcal{B}}(H)$, we clearly have $CNC=-N^{*}$ and $CTC=-T^{*}=-T$.
+By the fact that the class of $C$-skew symmetric operators is a subspace of ${\mathcal{B}}(\mathcal{H})$, we clearly have $CN_{k}C=-N_{k}^{*}$, for every $k\in K$, and $CTC=-T^{*}=-T$.
 Noting that $T$ is skew-symmetric, it must have a skew-symmetric matrix with respect to some orthonormal basis.
 In particular, since $T$ is a rank-two operator, its trace is well-defined and should be zero.
-Combining this with the fact that $T$ is self-adjoint, it follows that there exists orthonormal vectors $e_{1},e_{2}\in H$ and $\alpha\in\mathbb{R}$ such that
-${\rm ran}(T)={\rm ker}(T-\alpha I)\oplus{\rm ker}(T+\alpha I),$
-${\rm ker}(T-\alpha I)=\vee\{e_{1}\}\quad\mbox{and}\quad{\rm ker}(T+\alpha I)=\vee\{e_{2}\}.$
-As $CTC=-T$, we get $C(T+\alpha I)C=-(T-\alpha I)$; hence
-(3.4) $C(\vee\{e_{1}\})=C{\rm ker}(T-\alpha I)={\rm ker}(T+\alpha I)=\vee\{e_{2}\}.$
+Combining this with the fact that $T$ is self-adjoint, it follows that there exists orthonormal vectors $e_{1},e_{2}\in\mathcal{H}$ and $\alpha\in\mathbb{R}$ such that
+${\rm ran}(T)={\rm ker}(T-\alpha I_{\mathcal{H}})\oplus{\rm ker}(T+\alpha I_{\mathcal{H}}),$
+${\rm ker}(T-\alpha I_{\mathcal{H}})=\vee\{e_{1}\}\quad\mbox{and}\quad{\rm ker}(T+\alpha I_{\mathcal{H}})=\vee\{e_{2}\}.$
+As $CTC=-T$, we get $C(T+\alpha I_{\mathcal{H}})C=-(T-\alpha I_{\mathcal{H}})$; hence
+(3.4) $C(\vee\{e_{1}\})=C{\rm ker}(T-\alpha I_{\mathcal{H}})={\rm ker}(T+\alpha I_{\mathcal{H}})=\vee\{e_{2}\}.$
 Note that $x$ and $y$ have the same norm.
 In fact, $\langle x,x\rangle-\langle y,y\rangle$ is the trace of $T$.
-Hence, using [29, Theorem 2.1], we obtain a conjugation $J$ on $H$ satisfying $Jx=y$; one can check that $JTJ=-T$.
+Hence, using [30, Theorem 2.1], we obtain a conjugation $J$ on $\mathcal{H}$ satisfying $Jx=y$; one can check that $JTJ=-T$.
 The argument in (3.4) shows that $J(\vee\{e_{1}\})=\vee\{e_{2}\}$.
 Then, since $C$ and $J$ are isometries, there exist unimodular scalars $a,b\in\mathbb{C}$ such that $Ce_{1}=ae_{2}$ and $Je_{1}=be_{2}$.
-Furthermore, since $C^{2}=J^{2}=I$, we get $Ce_{2}=ae_{1}$ and $Je_{2}=be_{1}$.
+Furthermore, since $C^{-1}=C$ and $J^{-1}=J$, we get $Ce_{2}=ae_{1}$ and $Je_{2}=be_{1}$.
 Noting that $C$ and $J$ are both conjugate-linear, we get $Ch=a\overline{b}Jh$ for every $h\in\vee\{e_{1},e_{2}\}$.
-In particular, since $x,y\in{\rm ran}(T)=\vee\{e_{1},e_{2}\}$, we obtain $Cx=a\overline{b}Jx=a\overline{b}y$.
-Put $\widetilde{C}=\overline{a}bC$; one can easily check that $\widetilde{C}$ is a conjugation that satisfies $\widetilde{C}N\widetilde{C}=-N^{*}$ and $\widetilde{C}x=y$.
+In particular, since $x\in{\rm ran}(T)=\vee\{e_{1},e_{2}\}$, we obtain $Cx=a\overline{b}Jx=a\overline{b}y$.
+Put $\widetilde{C}=\overline{a}bC$; one can easily check that $\widetilde{C}$ is a conjugation that satisfies $\widetilde{C}N_{k}\widetilde{C}=-N_{k}^{*}$, for every $k\in K$, and $\widetilde{C}x=y$.
 This ends the proof of the theorem.
 ∎
+Next, we shall apply the main results of this section to characterize skew symmetry of operators of the form
+(3.5) $\sum_{j\in J}r_{j}(e_{j}\otimes e_{j}-f_{j}\otimes f_{j})+{\rm i}B$
+where the set $\{e_{j},f_{j}:j\in J\}\subset\mathcal{H}$ is orthonormal, $\{r_{j}\}_{j\in J}$ are distinct positive scalars and $B\in{\mathcal{B}}(\mathcal{H})$ is self-adjoint.
+Before proceeding, it is worth noting that skew symmetric operators having the previous form are dense in ${\rm SSO}$.
+Let us prove this claim.
+Let $T\in{\mathcal{B}}(\mathcal{H})$ be a $C$-skew symmetric operator for some conjugation $C$ on $\mathcal{H}$.
+A quick examination of the cartesian decomposition of $T$ shows that we can write $T=A+{\rm i}B$ where $A$ and $B$ are $C$-skew symmetric self-adjoint operators.
+Using [5, Lemma 3.2], we obtain a sequence of self-adjoint diagonalizable $C$-skew symmetric operators $A_{n}$ that converges to $A$, with $\dim{\rm ker}(A_{n}-rI_{\mathcal{H}})=1$ whenever $r$ is an eigenvalue of $A_{n}$.
+Now, by skew-symmetry of $A_{n}$, we have $\dim{\rm ker}(A_{n}-rI_{\mathcal{H}})=\dim{\rm ker}(A_{n}+rI_{\mathcal{H}})$ for all real scalars $r$.
+Consequently, $T$ can be approximated by operators of the form (3.5).
+Corollary 3.5.
+Let $T\in{\mathcal{B}}(\mathcal{H})$ be an operator of the form (3.5) and denote by $B_{1}$ the restriction of $B$ to $\left(L_{\{B\},\{e_{j},f_{j}:j\in J\}}\right)^{\perp}$.
+Then, $T$ is skew-symmetric if and only if $B_{1}\cong-B_{1}$ and there exists unimodular scalars $\lambda_{j}$, for $j\in J$, such that
+(3.6) $\lambda_{i}\left\langle E_{B}(\Delta)e_{i},e_{j}\right\rangle=\lambda_{j}\left\langle E_{B}(-\Delta)f_{j},f_{i}\right\rangle\;\mbox{and}\;\lambda_{i}\left\langle E_{B}(\Delta)e_{i},f_{j}\right\rangle=\lambda_{j}\left\langle E_{B}(-\Delta)e_{j},f_{i}\right\rangle$
+for every Borel subset $\Delta$ of $\mathbb{C}$.
+Proof.
+Assume that (3.6) holds for every Borel subset $\Delta$ of $\mathbb{C}$.
+Since $B_{1}\cong-B_{1}$, Lemma 3.7 implies that $VB_{1}V^{-1}=-B_{1}^{*}$ for some anti-unitary operator $V$.
+Then, applying Theorem 3.1 to the self-adjoint operator $B$ and the orthonormal subsets $\{\lambda_{j}e_{j}\}_{j\in J}$ and $\{f_{j}\}_{j\in J}$, we obtain a conjugation $C$ on $\mathcal{H}$ such that $CBC=-B$ and $C(\lambda e_{j})=f_{j}$ for every $j\in J$.
+One can check that $C(e_{j}\otimes e_{j}-f_{j}\otimes f_{j})C=-(e_{j}\otimes e_{j}-f_{j}\otimes f_{j})=-(e_{j}\otimes e_{j}-f_{j}\otimes f_{j})^{*}$ for every $j\in J$.
+Thus, $T$ is $C$-skew symmetric as the sum of two $C$-skew symmetric ones.
+This proves the sufficiency.
+Suppose that $T$ is $C$-skew symmetric for some conjugation $C$ on $\mathcal{H}$, and denote $A=\sum_{j\in J}r_{j}(e_{j}\otimes e_{j}-f_{j}\otimes f_{j})$.
+Then, $A$ and $B$ are both $C$-skew symmetric.
+Moreover, for every $j\in J$, ${\rm ker}(A-r_{j}I_{\mathcal{H}})=\vee\{e_{j}\}$ and ${\rm ker}(A+r_{j}I_{\mathcal{H}})=\vee\{f_{i}\}$.
+As in (3.4), we have that $Ce_{j}\in{\rm span}\{f_{j}\}$ for every $j\in J$.
+Hence, there exist unimodular scalars $\lambda_{j}$, for $j\in J$, so that $C(\lambda_{j}e_{j})=f_{j}$ for every $j\in J$.
+The necessity follows now by combining Theorem 3.1 and Lemma 3.7.
+∎
+Remark 3.6.
+  1. (i)
+    If the operator $T$ has rather the form
+    $T=B+{\rm i}\sum_{j\in J}r_{j}(e_{j}\otimes e_{j}-f_{j}\otimes f_{j})$
+    with $\{e_{j},f_{j}\}_{j\in J}$, $\{r_{j}\}_{j\in J}$ and $B$ being as described in (3.5), then, taking into account that skew symmetry is invariant under scalar multiplication, the equivalence in the previous corollary still holds true.
+  2. (ii)
+    If $J$ is finite, then using Theorem 3.2, one can make minor changes in the proof of Corollary 3.5 to show that the condition “$B_{1}\cong-B_{1}$” can be replaced with the simpler one “$B\cong-B$”.
 The remainder of this section is devoted to proving Theorems 3.1 and 3.2.
-3.2. Proof of Theorem 3.1 and 3.2
-First, we begin with the following characterization of skew-symmetric normal operators.
+3.2. Proof of Theorems 3.1 and 3.2
+The following lemma provides conditions under which a family of mutually commuting normal operators have a common orthonormal basis in which their corresponding matrices are skew-symmetric.
+This result is inspired by [18, Theorem 1.10].
 Lemma 3.7.
-Let $N\in{\mathcal{B}}(H)$ be a normal operator. The following statements are equivalent:
+Let $\{N_{k}\}_{k\in K}\subset{\mathcal{B}}(\mathcal{H})$ be mutually commuting normal operators. The following statements are equivalent:
   1. (i)
-    There exists a unitary operator $U$ on $H$ such that $UNU^{*}=-N$.
+    There exists a conjugation $C$ on $\mathcal{H}$ such that
+    $CN_{k}C=-N_{k}^{*}\quad\mbox{for every $k\in K$}.$
   2. (ii)
-    There exists a conjugation $C$ on $H$ such that $CNC=-N^{*}$.
+    There exists an anti-unitary operator $V$ on $\mathcal{H}$ such that
+    $VN_{k}V^{-1}=-N_{k}^{*}\quad\mbox{for every $k\in K$}.$
   3. (iii)
-    There exists an anti-unitary operator $V$ on $H$ such that $VNV^{-1}=-N^{*}$.
+    There exists a unitary operator $U$ on $\mathcal{H}$ such that
+    $UN_{k}U^{*}=-N_{k}\quad\mbox{for every $k\in K$}.$
 Proof.
-The implication (ii)$\implies$(iii) is obvious.
+The implication (i)$\implies$(ii) is obvious.
+(ii)$\implies$(iii).
+Let $V$ be an anti-unitary operator on $\mathcal{H}$ such that $VN_{k}V^{-1}=-N_{k}^{*}$ for every $k\in K$.
+As in (2.10), there exists a conjugation $J$ on $\mathcal{H}$ that satisfies $JN_{k}J=N_{k}^{*}$ for every $k\in K$.
+Hence, we have $VN_{k}V^{-1}=-JN_{k}J$, and so, since $J^{-1}=J$, we get
+(3.7) $(JV)N_{k}(V^{-1}J)=-J^{2}N_{k}J^{2}=-N_{k}\quad\mbox{for every $k\in K$}.$
+Put $U=JV$; then $U\in{\mathcal{B}}(\mathcal{H})$ is unitary and $U^{*}=U^{-1}=(JV)^{-1}=V^{-1}J^{-1}=V^{-1}J$. Therefore, for any $k\in K$, $UN_{k}U^{*}=-N_{k}$.
 (iii)$\implies$(i).
-Let $V$ be an anti-unitary operator on $H$ such that $VNV^{-1}=-N^{*}$.
-By normality of $N$, there exists a conjugation $J$ on $H$ that satisfies $JNJ=N^{*}$.
-Hence, we have $VNV^{-1}=-JNJ$, and so, since $J^{2}=I$, we get
-$(JV)N(V^{-1}J)=-J^{2}NJ^{2}=-N.$
-Put $U=JV$; then $U\in{\mathcal{B}}(H)$ is unitary and $U^{*}=U^{-1}=(JV)^{-1}=V^{-1}J^{-1}=V^{-1}J$. Therefore, $UNU^{*}=-N$.
-(i)$\implies$(ii).
-Let $U\in{\mathcal{B}}(H)$ be a unitary operator satisfying $UN=-NU$.
+It is easy to see that we may assume without loss of generality that $K$ is infinite.
+Let $U\in{\mathcal{B}}(\mathcal{H})$ be a unitary operator satisfying $UN_{k}=-N_{k}U$ for every $k\in K$.
+First, let us reduce the problem to the case where $\bigcap_{k\in K}{\rm ker}(N_{k})=\{0\}$.
+Denote this subspace by $\mathcal{L}_{\infty}$.
+One can easily see that $\mathcal{L}_{\infty}$ reduces $U$, and that each $N_{k}$ has the form $N_{k}=M_{k}\oplus 0$ with respect to the decomposition $\mathcal{H}=\mathcal{L}_{\infty}^{\perp}\oplus\mathcal{L}_{\infty}$.
+Clearly, $\bigcap_{k\in K}{\rm ker}(M_{k})=\{0\}$; hence, if a conjugation $C$ on $\mathcal{L}_{\infty}^{\perp}$ exists such that $CM_{k}C=-M_{k}^{*}$ for each $k\in K$, then one can see that the map $C\oplus C_{\infty}$, where $C_{\infty}$ is an arbitrary conjugation on $\mathcal{L}_{\infty}$, is a conjugation on $\mathcal{H}$ that satisfies all the desired equalities.
+Note that we can further assume that there exists a countable subset $\{N_{k_{p}}\}_{p\in\mathbb{Z}^{+}}$ of $\{N_{k}\}_{k\in K}$ such that $\bigcap_{p\in\mathbb{Z}^{+}}{\rm ker}(N_{k_{p}})=\{0\}$.
+In fact, since $\mathcal{H}=\bigvee\left(\bigcup_{k\in K}{\rm ker}(N_{k})^{\perp}\right)$, one can use the separability of $\mathcal{H}$ to show that there exists a countable subset $\{k_{p}\}_{p\in\mathbb{Z}^{+}}\subseteq K$ so that $\mathcal{H}=\bigvee\left(\bigcup_{p\in\mathbb{Z}^{+}}{\rm ker}(N_{k_{p}})^{\perp}\right)$.
 Fix a Borel subset $\Delta_{0}\subset\mathbb{C}$ such that $\mathbb{C}\setminus\Delta_{0}=-\Delta_{0}\cup\{0\}$.
-By Lemma 2.11, we have $UE_{N}(\Delta_{0})=E_{N}(-\Delta_{0})U$ and $UE_{N}(-\Delta_{0})=E_{N}(\Delta_{0})U$.
-Consequently,
-$U{\rm ran}(E_{N}(\Delta_{0}))={\rm ran}(E_{N}(-\Delta_{0}))\quad\mbox{and}\quad U{\rm ran}(E_{N}(-\Delta_{0}))={\rm ran}(E_{N}(\Delta_{0})).$
-Given that $\mathbb{C}=\Delta_{0}\cup(-\Delta)\cup\{0\}$ is a partition, it follows from the above equalities that $U$ has the form
-$U=\begin{bmatrix}0&*&0\\ W&0&0\\ 0&0&*\end{bmatrix}\begin{array}[]{l}{\rm ran}(E_{N}(\Delta_{0}))\\ {\rm ran}(E_{N}(-\Delta_{0}))\\ {\rm ran}(E_{N}(\{0\}))\end{array}$
-Where $W$ is an isometric isomorphism between Hilbert spaces. Noting that $N$ can also be written as
-$N=\begin{bmatrix}N_{1}&0&0\\ 0&N_{2}&0\\ 0&0&0\end{bmatrix}\begin{array}[]{l}{\rm ran}(E_{N}(\Delta_{0}))\\ {\rm ran}(E_{N}(-\Delta_{0}))\\ {\rm ran}(E_{N}(\{0\}))\end{array},$
-and since $UN=-NU$, simple matrix calculations show that $WN_{1}=-N_{2}W$.
-This implies that $N_{1}\cong-N_{2}$, and hence $N_{|{\rm ker}(N)^{\perp}}=N_{1}\oplus N_{2}\cong N_{1}\oplus-N_{1}$.
-Therefore, $N$ is skew symmetric by [17, Theorem 1.10].
+By Lemma 2.15, we have $UE_{N_{k_{1}}}(\Delta_{0})=E_{N_{k_{1}}}(-\Delta_{0})U$ and $UE_{N_{k_{1}}}(-\Delta_{0})=E_{N_{k_{1}}}(\Delta_{0})U.$ Consequently,
+(3.8) $U{\rm ran}(E_{N_{k_{1}}}(\Delta_{0}))={\rm ran}(E_{N_{k_{1}}}(-\Delta_{0}))\;\;\mbox{and}\;\;U{\rm ran}(E_{N_{k_{1}}}(-\Delta_{0}))={\rm ran}(E_{N_{k_{1}}}(\Delta_{0})).$
+For convenience, put $\mathcal{H}_{+}={\rm ran}(E_{N_{k_{1}}}(\Delta_{0}))$, $\mathcal{H}_{-}={\rm ran}(E_{N_{k_{1}}}(-\Delta_{0}))$ and $\mathcal{L}_{1}={\rm ran}(E_{N_{k_{1}}}(\{0\}))={\rm ker}(N_{k_{1}})$.
+Given that $\mathbb{C}=\Delta_{0}\cup(-\Delta_{0})\cup\{0\}$ is a partition, it follows from (3.8) that $U$ has the form
+$U=\begin{bmatrix}0&*&0\\ W&0&0\\ 0&0&U_{1}\end{bmatrix}\begin{array}[]{l}\mathcal{H}_{+}\\ \mathcal{H}_{-}\\ \mathcal{L}_{1}\end{array}$
+where $W$ and $U_{1}$ are isometric isomorphisms between Hilbert spaces. Additionally, since for every $k\in K$,
+$N_{k}E_{N_{k_{1}}}(\Delta_{0})=E_{N_{k_{1}}}(\Delta_{0})N_{k}\quad\mbox{and}\quad N_{k}E_{N_{k_{1}}}(-\Delta_{0})=E_{N_{k_{1}}}(-\Delta_{0})N_{k},$
+we can also write
+$N_{k}=\begin{bmatrix}N_{k,+}&0&0\\ 0&N_{k,-}&0\\ 0&0&R_{k,1}\end{bmatrix}\begin{array}[]{l}\mathcal{H}_{+}\\ \mathcal{H}_{-}\\ \mathcal{L}_{1}\end{array}$
+where $N_{k,+}$, $N_{k,-}$ and $R_{k,1}$, $k\in K$, are normal operators such that the $R_{k,1}$’s are mutually commuting.
+Since $UN_{k}=-N_{k}U$, simple matrix calculations show that
+(3.9) $WN_{k,+}=-N_{k,-}W\quad\mbox{and}\quad U_{1}R_{k,1}=-R_{k,1}U_{1}\quad\mbox{for every $k\in K$}.$
+Put $\mathcal{M}_{1}=\mathcal{H}_{+}\oplus\mathcal{H}_{-}$ and let $V:\mathcal{M}_{1}\to\mathcal{H}_{-}\oplus\mathcal{H}_{-}$ be the unitary operator given by
+$\begin{array}[]{rcl}V:\mathcal{H}_{+}\oplus\mathcal{H}_{-}&\to&\mathcal{H}_{-}\oplus\mathcal{H}_{-}\\ (x,y)&\mapsto&(Wx,y).\end{array}$
+Then, it is easy to see that $V{N_{k}}_{|\mathcal{M}_{1}}V^{-1}=(-N_{k,-})\oplus N_{k,-}$ for every $k\in K$.
+Since the normal operators $\{N_{k,-}\}_{k\in K}$ are mutually commuting, as in (2.10), there exists a conjugation $J$ on $\mathcal{H}_{-}$ such that $JN_{k,-}J=\left(N_{k,-}\right)^{*}$ for every $k\in K$.
+Put $J_{1}=\begin{bmatrix}0&J\\ J&0\\ \end{bmatrix}\begin{array}[]{l}\mathcal{H}_{-}\\ \mathcal{H}_{-}\end{array}.$ It is easy to see that $J_{1}$ is a conjugation on $\mathcal{H}_{-}\oplus\mathcal{H}_{-}$ satisfying
+$J_{1}\left((-N_{k,-})\oplus N_{k,-}\right)J_{1}=-\left((-N_{k,-})\oplus N_{k,-}\right)^{*}\quad\mbox{for every $k\in K$}.$
+Hence, $C_{1}:=V^{-1}J_{1}V$ is a conjugation on $\mathcal{M}_{1}$ that satisfies $C_{1}{N_{k}}_{|\mathcal{M}_{1}}C_{1}=-\left({N_{k}}_{|\mathcal{M}_{1}}\right)^{*}$ for every $k\in K$.
+Since $\{R_{k,1}\}_{k\in K}$ are mutually commuting and $U_{1}R_{k,1}U_{1}^{*}=-R_{k,1}$ for every $k\in K$, by applying the preceding argument to $R_{k_{2},1}\in\mathcal{B}(\mathcal{L}_{1})$, we obtain subspaces $\mathcal{M}_{2}$ and $\mathcal{L}_{2}$ of $\mathcal{L}_{1}$, with $\mathcal{L}_{1}=\mathcal{M}_{2}\oplus\mathcal{L}_{2}$, a conjugation $C_{2}$ on $\mathcal{M}_{2}$ such that
+$C_{2}{N_{k}}_{|\mathcal{M}_{2}}C_{2}=C_{2}{R_{k,1}}_{|\mathcal{M}_{2}}C_{2}=-\left({R_{k,1}}_{|\mathcal{M}_{2}}\right)^{*}=-\left({N_{k}}_{|\mathcal{M}_{2}}\right)^{*}\quad\mbox{for every $k\in K$},$
+a unitary operator $U_{2}$ on $\mathcal{L}_{2}$, and a set of mutually commuting normal operators $\{R_{k,2}\}_{k\in K}$ on $\mathcal{L}_{2}$ such that ${N_{k}}_{|\mathcal{L}_{2}}=R_{k,2}$ and $U_{2}R_{k,2}U_{2}^{*}=-\left(R_{k,2}\right)^{*}$.
+This process produces a sequence $\mathcal{M}_{i}$, $i\geq 1$, of reducing subspaces of $N_{k}$, for $k\in K$, and conjugations $C_{i}$ on $\mathcal{M}_{i}$ such that ${N_{k}}_{|\mathcal{M}_{i}}$ is $C_{i}$-skew symmetric for all $i\in\mathbb{Z}^{+}$ and $k\in K$.
+For each $i\geq 1$, it is easy to see that $\left(\sum_{j=1}^{i}\mathcal{M}_{j}\right)^{\perp}=\mathcal{L}_{i}=\bigcap_{j=1}^{i}{\rm ker}(N_{k_{j}}).$ Hence,
+$\left(\bigoplus_{i=1}^{\infty}\mathcal{M}_{i}\right)^{\perp}=\bigcap_{i=1}^{\infty}\left(\sum_{j=1}^{i}\mathcal{M}_{j}\right)^{\perp}=\bigcap_{i=1}^{\infty}{\rm ker}(N_{k_{i}})=\{0\}.$
+Consequently, $\mathcal{H}=\bigoplus_{i=1}^{\infty}\mathcal{M}_{i}$; moreover, one can check that $C:=\bigoplus_{i=1}^{\infty}C_{i}$ is a conjugation on $\mathcal{H}$ satisfying $CN_{k}C=-N_{k}^{*}$ for every $k\in K$.
 ∎
-Note that every bounded linear $T$ acting on a Hilbert space can be written as $T=N\oplus A$ where $N$ is normal and $A$ is an operator that has no reducing subspace for which the restriction of $A$ is normal (of course, either $N$ or $A$ could be absent).
+Note that every bounded linear operator $T$ acting on a Hilbert space can be written as $T=N\oplus A$ where $N$ is normal and $A$ is an operator that has no reducing subspace for which the restriction of $A$ is normal (of course, either $N$ or $A$ could be absent).
 In fact, $N$ is the restriction of $T$ to the reducing subspace
-(3.5) $H_{\rm nor}=\bigcap_{1\leq n,m<\infty}{\rm ker}({T^{*}}^{n}T^{m}-T^{m}{T^{*}}^{n}).$
+(3.10) $\mathcal{H}_{\rm nor}=\bigcap_{1\leq n,m<\infty}{\rm ker}({T^{*}}^{n}T^{m}-T^{m}{T^{*}}^{n}).$
 The operator $N$ is called the normal part of $T$, while $A$ is called the abnormal part of $T$. In this case, we denote $T_{\rm nor}=N$ and $T_{\rm abnor}=A$.
@@
 (3.11) $(M\oplus T)_{\rm nor}=M\oplus T_{\rm nor}\quad\mbox{and}\quad(M\oplus T)_{\rm abnor}=T_{\rm abnor}.$
-The following lemma is an immediate consequence of [28, Proposition 3.10] and (3.6).
+The following lemma is an immediate consequence of [29, Proposition 3.10] and (3.11).
 Lemma 3.8.
-Let $M\in{\mathcal{B}}(H)$ be normal and $T\in{\mathcal{B}}(K)$ such that $M\oplus T$ is skew-symmetric. Then, $M\oplus T_{\rm nor}$ and $T_{\rm abnor}$ are skew-symmetric.
+Let $M\in{\mathcal{B}}(\mathcal{H})$ be normal and $T\in{\mathcal{B}}(\mathcal{K})$ such that $M\oplus T$ is skew-symmetric. Then, $M\oplus T_{\rm nor}$ and $T_{\rm abnor}$ are skew-symmetric.
 Lemma 3.9.
-Let $N\in{\mathcal{B}}(H)$ be a skew-symmetric normal operator and $\Delta\subseteq\mathbb{C}$ a Borel subset such $\Delta=-\Delta$.
+Let $N\in{\mathcal{B}}(\mathcal{H})$ be a skew-symmetric normal operator and $\Delta\subseteq\mathbb{C}$ a Borel subset such $\Delta=-\Delta$.
 Then, the restriction of $N$ to the reducing subspace ${\rm ran}(E_{N}(\Delta))$ is skew-symmetric.
 Proof.
-Let $C$ be a conjugation on $H$ such that $CNC=-N^{*}$.
+Let $C$ be a conjugation on $\mathcal{H}$ such that $CNC=-N^{*}$.
 By Lemma 2.15, we have $CE_{N}(\Delta)=E_{N}(-\Delta)C=E_{N}(\Delta)C$; hence, $C{\rm ran}(E_{N}(\Delta))={\rm ran}(E_{N}(\Delta))$.
@@
 ∎
-In [16, Lemma 3.2] and [2, Corollary 3.4], the authors proved that if $N$ is normal, then a direct sum $N\oplus T$ is complex symmetric (resp.
+In [17, Lemma 3.2] and [2, Corollary 3.4], the authors proved that if $N$ is normal, then a direct sum $N\oplus T$ is complex symmetric (resp.
 conjugate-normal) if and only if $T$ is complex-symmetric (resp.
@@
 Theorem 3.10.
-Let $N\in{\mathcal{B}}(H)$ be normal. The followings statements are equivalents:
+Let $N\in{\mathcal{B}}(\mathcal{H})$ be normal. The followings statements are equivalent:
   1. (i)
-    For every separable complex Hilbert space $K$ and every $T\in\mathcal{B}(K)$,
+    For every separable complex Hilbert space $\mathcal{K}$ and every $T\in\mathcal{B}(\mathcal{K})$,
 […]
@@
 Suppose statement (i) holds.
-Taking $T=0$, the zero operator on any Hilbert space $K$, we obtain that $N\oplus 0$ is skew-symmetric.
-By [17, Lemma 2.1], it follows that $N_{|({\rm ker}N)^{\perp}}$ is skew-symmetric.
+Taking $T=0$, the zero operator on any Hilbert space $\mathcal{K}$, we obtain that $N\oplus 0$ is skew-symmetric.
+By [18, Lemma 2.1], it follows that $N_{|({\rm ker}N)^{\perp}}$ is skew-symmetric.
 Therefore, $N=N_{|({\rm ker}N)^{\perp}}\oplus 0$ is skew-symmetric.
@@
 Fix a bounded linear operator $T$, assume that $N\oplus T$ is skew-symmetric, and let us show that $T$ is skew-symmetric.
-In view of Lemma 3.6 and [17, Lemma 2.1], we can assume without loss of generality that $T$ is normal and $N=N_{|{\rm ker}(N)^{\perp}}$. Moreover, by (3.1) and (3.3), we may directly assume that
-$N=\bigoplus_{1\leq i<\infty}M_{\mu_{i}}^{(i)}\quad\mbox{and}\quad H=\bigoplus_{1\leq i<\infty}L^{2}(\mu_{i})^{(i)}$
+In view of Lemma 3.8 and [18, Lemma 2.1], we can assume without loss of generality that $T$ is normal and $N=N_{|{\rm ker}(N)^{\perp}}$. Moreover, by (3.1) and (3.3), we may directly assume that
+$N=\bigoplus_{1\leq i<\infty}M_{\mu_{i}}^{(i)}\quad\mbox{and}\quad\mathcal{H}=\bigoplus_{1\leq i<\infty}L^{2}(\mu_{i})^{(i)}$
 where the $\mu_{i}$’s are mutually singular and satisfy $\mu_{i}(\Delta)=\mu_{i}(-\Delta)$ for every Borel subset $\Delta\subseteq\mathbb{C}$.
@@
 Let $\Delta_{0}=\mathbb{C}\setminus\cup_{i\in\mathbb{Z}^{+}}\Delta_{i}$; clearly, $\Delta_{0}=-\Delta_{0}$.
-For convenience, put $N_{i}=M_{\mu_{i}}^{(i)}$, $H_{i}=L^{2}(\mu_{i})^{(i)}$ for $i\in\mathbb{Z}^{+}$, and let $T=T_{0}\oplus\bigoplus_{i\in\mathbb{Z}^{+}}T_{i}$ be the decomposition of $T$ with respect to the orthogonal decomposition $K={\rm ran}(E_{T}(\Delta_{0}))\oplus\bigoplus_{i\in\mathbb{Z}^{+}}{\rm ran}(E_{T}(\Delta_{i}))$.
-For $0\leq i<\infty$, put $K_{i}={\rm ran}(E_{T}(\Delta_{i}))$.
+For convenience, put $N_{i}=M_{\mu_{i}}^{(i)}$, $\mathcal{H}_{i}=L^{2}(\mu_{i})^{(i)}$ for $i\in\mathbb{Z}^{+}$, and let $T=T_{0}\oplus\bigoplus_{i\in\mathbb{Z}^{+}}T_{i}$ be the decomposition of $T$ with respect to the orthogonal decomposition $\mathcal{K}={\rm ran}(E_{T}(\Delta_{0}))\oplus\bigoplus_{i\in\mathbb{Z}^{+}}{\rm ran}(E_{T}(\Delta_{i}))$.
+For $0\leq i<\infty$, put $\mathcal{K}_{i}={\rm ran}(E_{T}(\Delta_{i}))$.
 Since by (3.12), $\mu_{i}(\Delta_{0})=\mu_{i}(\Delta_{j})=0$ all $i,j\in\mathbb{Z}^{+}$ with $i\neq j$, it follows that $E_{N_{i}}(\Delta_{0})=E_{N_{i}}(\Delta_{j})=0$. Therefore,
-${\rm ran}(E_{N\oplus T}(\Delta_{0}))=K_{0}\quad\mbox{and}\quad{\rm ran}(E_{N\oplus T}(\Delta_{i}))=H_{i}\oplus K_{i}\;\;\mbox{for every }i\in\mathbb{Z}^{+}.$
+${\rm ran}(E_{N\oplus T}(\Delta_{0}))=\mathcal{K}_{0}\quad\mbox{and}\quad{\rm ran}(E_{N\oplus T}(\Delta_{i}))=\mathcal{H}_{i}\oplus\mathcal{K}_{i}\;\;\mbox{forevery }i\in\mathbb{Z}^{+}.$
 It follows that
@@
 Fix $i\in\mathbb{Z}^{+}$, and let us prove that $T_{i}$ is skew-symmetric.
-Given $\mu_{i}(\Delta)=\mu_{i}(-\Delta)$ for every Borel subset $\Delta\subseteq\mathbb{C}$, [17, Theorem 1.11] implies that $N_{i}$ is skew-symmetric; hence, $N_{i}\cong-N_{i}$ by Lemma 3.5.
+Given $\mu_{i}(\Delta)=\mu_{i}(-\Delta)$ for every Borel subset $\Delta\subseteq\mathbb{C}$, [18, Theorem 1.11] implies that $N_{i}$ is skew-symmetric; hence, $N_{i}\cong-N_{i}$ by Lemma 3.7.
 Combining this with (3.13), we obtain $N_{i}\oplus T_{i}\cong N_{i}\oplus-T_{i}$.
-Since $M_{\mu_{i}}$ is a star-cyclic normal operator ([8, Page 269, Theorem 3.4]), applying [8, Page 295, Proposition 10.6] $i$ times to $M_{\mu_{i}}$, we conclude that $T_{i}\cong-T_{i}$, and so $T_{i}$ is skew-symmetric by Lemma 3.5.
+Since $M_{\mu_{i}}$ is a star-cyclic normal operator ([9, p.
+269, Theorem 3.4]), applying [9, p.
+295, Proposition 10.6] $i$ times to $M_{\mu_{i}}$, we conclude that $T_{i}\cong-T_{i}$, and so $T_{i}$ is skew-symmetric by Lemma 3.7.
 Thus, $T=T_{0}\oplus\bigoplus_{i\in\mathbb{Z}^{+}}T_{i}$ is skew-symmetric as a direct sum of skew-symmetric operators. ∎
 Lemma 3.11.
-Let $N\in{\mathcal{B}}(H)$ be normal, and suppose that there exist reducing subspaces $\{H_{i}\}_{1\leq i\leq r}$, $r<\infty$, of $N$ such that $H=\vee(\sum_{i=1}^{r}H_{i})$ and $N_{|H_{i}}$ is star-cyclic for each $i\in\{1,2,\dots,r\}$.
+Let $N\in{\mathcal{B}}(\mathcal{H})$ be normal, and suppose that there exist reducing subspaces $\{\mathcal{H}_{i}\}_{1\leq i\leq r}$, $r<\infty$, of $N$ such that $\mathcal{H}=\vee(\sum_{i=1}^{r}\mathcal{H}_{i})$ and $N_{|\mathcal{H}_{i}}$ is star-cyclic for each $i\in\{1,2,\dots,r\}$.
 Then, for every reducing subspace $L$ of $N$, we have $m(N_{|L})<\infty$.
@@
 Assume the lemma holds for some integer $r\geq 1$.
-Let $H_{1},H_{2},\dots H_{r+1}$ be $r+1$ reducing subspaces of $N$ such that $H=\vee(\sum_{i=1}^{r+1}H_{i})$ and $N_{|H_{i}}$ is star cyclic for $i\in\{1,2,\dots,r+1\}$.
-Let $K=\vee(\sum_{i=1}^{r}H_{i})$ and $P\in{\mathcal{B}}(H)$ be the orthogonal projection onto $K$.
-Then, $K$ reduces $N$; moreover, by the induction hypothesis, $N_{|K}$ is a finite direct sum of star-cyclic operators.
-By star-cyclicity of $N_{|H_{r+1}}$, there exists $h\in H$ such that $H_{r+1}=L_{N,h}$.
-To complete the proof, it is clear that it suffices to show that $H=K\oplus L_{N,(I-P)h}$.
-Note that $K$ and $L_{N,(I-P)h}$ are orthogonal.
-In fact, since $K$ reduces $N$, its orthogonal projection $P$, and hence $(I-P)$, commutes with $N$ and $N^{*}$.
+Let $\mathcal{H}_{1},\mathcal{H}_{2},\dots\mathcal{H}_{r+1}$ be $r+1$ reducing subspaces of $N$ such that $\mathcal{H}=\vee(\sum_{i=1}^{r+1}\mathcal{H}_{i})$ and $N_{|\mathcal{H}_{i}}$ is star cyclic for $i\in\{1,2,\dots,r+1\}$.
+Let $\mathcal{K}=\vee(\sum_{i=1}^{r}\mathcal{H}_{i})$ and $P\in{\mathcal{B}}(\mathcal{H})$ be the orthogonal projection onto $\mathcal{K}$.
+Then, $\mathcal{K}$ reduces $N$; moreover, by the induction hypothesis, $N_{|\mathcal{K}}$ is a finite direct sum of star-cyclic operators.
+By star-cyclicity of $N_{|\mathcal{H}_{r+1}}$, there exists $h\in\mathcal{H}$ such that $\mathcal{H}_{r+1}=L_{\{N\},\{h\}}$.
+To complete the proof, it is clear that it suffices to show that $\mathcal{H}=\mathcal{K}\oplus L_{\{N\},\{(I-P)h\}}$.
+Note that $\mathcal{K}$ and $L_{\{N\},\{(I-P)h\}}$ are orthogonal.
+In fact, since $\mathcal{K}$ reduces $N$, its orthogonal projection $P$, and hence $(I-P)$, commutes with $N$ and $N^{*}$.
 Consequently,
-$L_{N,(I-P)h}$ =$=\bigvee\big{\{}p(N,N^{*})(I-P)h:p\in\mathbb{C}[x,y]\big{\}}$
-$=(I-P)\bigvee\big{\{}p(N,N^{*})h:p\in\mathbb{C}[x,y]\big{\}}\subseteq K^{\perp}.$
-It remains to prove that $H=K+L_{N,(I-P)h}$; we need only to show that $H_{r+1}\subseteq K+L_{N,(I-P)h}$.
-For $p\in\mathbb{C}[x,y]$, we have $p(N,N^{*})h=p(N,N^{*})Ph+p(N,N^{*})(I-P)h$.
-Since $Ph\in K$ and $K$ reduces $p(N,N^{*})$, we have $p(N,N^{*})Ph\in K$.
-Hence, $p(N,N^{*})h\in K+L_{N,(I-P)h}$, and thus $H_{r+1}\subseteq K+L_{N,(I-P)h}$.
+$L_{\{N\},\{(I-P)h\}}$ =$=\bigvee\big{\{}p(N,N^{*})(I-P)h:p\in\mathcal{P}_{2}(\mathbb{C})\big{\}}$
+$=(I-P)\bigvee\big{\{}p(N,N^{*})h:p\in\mathcal{P}_{2}(\mathbb{C})\big{\}}\subseteq\mathcal{K}^{\perp}.$
+It remains to prove that $\mathcal{H}=\mathcal{K}+L_{\{N\},\{(I-P)h\}}$; we need only to show that $\mathcal{H}_{r+1}\subseteq\mathcal{K}+L_{\{N\},\{(I-P)h\}}$.
+For $p\in\mathcal{P}_{2}(\mathbb{C})$, we have
+$p(N,N^{*})h=p(N,N^{*})Ph+p(N,N^{*})(I-P)h.$
+Since $Ph\in\mathcal{K}$ and $\mathcal{K}$ reduces $p(N,N^{*})$, we have $p(N,N^{*})Ph\in\mathcal{K}$.
+Hence, $p(N,N^{*})h\in\mathcal{K}+L_{\{N\},\{(I-P)h\}}$, and thus $\mathcal{H}_{r+1}\subseteq\mathcal{K}+L_{\{N\},\{(I-P)h\}}$.
 Now, let $L$ be a reducing subspace of $N$ and assume, for contradiction, that $m(N_{|L})\nless\infty$.
@@
 $N\oplus M_{\mu}^{(2)}\cong N_{0}\oplus M_{\mu}^{(\infty)}\oplus M_{\mu}^{(2)}\cong N_{0}\oplus M_{\mu}^{(\infty)}\oplus M_{\mu}\cong N\oplus M_{\mu}.$
-Since $N$ is a direct sum of $n$ star-cyclic normal operators $M_{1},M_{2},\dots,M_{n}$ with $n<\infty$, applying [8, Page 295, Proposition 10.6] to each $M_{i}$, we conclude $M_{\mu}\cong M_{\mu}^{(2)}$; which contradicts (3.2).
+Since $N$ is a direct sum of $n$ star-cyclic normal operators $M_{1},M_{2},\dots,M_{n}$ with $n<\infty$, applying [9, p.
+295, Proposition 10.6] to each $M_{i}$, we conclude $M_{\mu}\cong M_{\mu}^{(2)}$; which contradicts (3.2).
 ∎
@@
 (ii)$\implies$(iii).
-First, let us show that $N_{|L_{N,Z}^{\perp}}$ is skew-symmetric.
-By the argument in (2.7), there exists a conjugation $J$ on $H$ so that $(JV)N=-N(JV)$ and $JN^{*}J=N$.
-By the Fuglede-Putnam Theorem, it follows that $(JV)N^{*}=-N^{*}(JV)$; hence, $VN^{*}=J^{2}VN^{*}=-JN^{*}JV=-NV$.
-Consequently, for any polynomial $p\in\mathbb{C}[x,y]$, one can easily see that $Vp(N,N^{*})=p(-N,-N^{*})^{*}V$.
+First, let us show that $V_{1}$ exists.
+Since $VN_{k}V^{-1}=-N_{k}^{*}$, one can use (2.7) to show that $VN_{k}^{*}V^{-1}=-N_{k}$ for every $k\in K$.
+Consequently, for any $n\in\mathbb{Z}^{+}$, $p\in\mathcal{P}_{2n}(\mathbb{C})$ and $\{k_{l}\}_{l=1}^{n}\subseteq K$, one can easily see that
+$Vp(N_{k_{1}},N_{k_{1}}^{*},\dots,N_{k_{n}},N_{k_{n}}^{*})=p(-N_{k_{1}},-N_{k_{1}}^{*},\dots,-N_{k_{n}},-N_{k_{n}}^{*})^{*}V.$
 Therefore, for every $i\in I$,
-$Vp(N,N^{*})x_{i}=p(-N,-N^{*})^{*}Vx_{i}=p(-N,-N^{*})^{*}y_{i}$
+$Vp(N_{k_{1}},N_{k_{1}}^{*},\dots,N_{k_{n}},N_{k_{n}}^{*})x_{i}$ $=p(-N_{k_{1}},-N_{k_{1}}^{*},\dots,-N_{k_{n}},-N_{k_{n}}^{*})^{*}Vx_{i}$
+$=p(-N_{k_{1}},-N_{k_{1}}^{*},\dots,-N_{k_{n}},-N_{k_{n}}^{*})^{*}y_{i}$
 and
-$Vp(N,N^{*})y_{i}=p(-N,-N^{*})^{*}Vy_{i}=p(-N,-N^{*})^{*}x_{i}.$
-This implies that if $L_{0}={\rm span}\big{\{}p(N,N^{*})x_{i}+q(N,N^{*})y_{i}:p,q\in\mathbb{C}[x,y],i\in I\big{\}}$, then $VL_{0}=L_{0}$.
-Taking the closure of $L_{0}$ in $H$, we obtain $VL_{N,Z}=L_{N,Z}$; moreover, since $V$ is anti-unitary, we get $VL_{N,Z}^{\perp}=L_{N,Z}^{\perp}$.
-Let $V_{1}$ and $N_{1}$ denote the restrictions of $V$ and $N$ to $L_{N,Z}^{\perp}$, respectively.
-Then, $V_{1}$ is anti-unitary on $L_{N,Z}^{\perp}$, and the relation $VN=-N^{*}V$ implies $V_{1}N_{1}=-N_{1}^{*}V_{1}$.
-Consequently, $N_{1}$ is skew-symmetric by Lemma 3.5.
-Now, by Lemma 2.11, for every Borel subset $\Delta\subseteq\mathbb{C}$, we have $VE_{N}(-\Delta)=E_{N}(\Delta)V$. Hence, using (2.9), for all $i,j\in I$,
-$\langle E_{N}(\Delta)x_{i},y_{j}\rangle$ $=\langle E_{N}(\Delta)Vy_{i},Vx_{j}\rangle=\langle VE_{N}(-\Delta)y_{i},Vx_{j}\rangle$
-$=\langle V^{-1}Vx_{j},E_{N}(-\Delta)y_{i}\rangle=\langle E_{N}(-\Delta)x_{j},y_{i}\rangle,$
-and similarly, $\langle E_{N}(\Delta)x_{i},x_{j}\rangle=\langle E_{N}(-\Delta)y_{j},y_{i}\rangle$.
-(iii)$\implies$(i). Let $M=-N$. For every Borel subset $\Delta\subseteq\mathbb{C}$ and for all $i,j\in I$,
-$\langle E_{N}(\Delta)x_{i},y_{j}\rangle=\langle E_{N}(-\Delta)x_{j},y_{i}\rangle=\langle E_{-N}(\Delta)x_{j},y_{i}\rangle=\langle E_{M}(\Delta)x_{j},y_{i}\rangle,$
-and similarly, $\langle E_{N}(\Delta)x_{i},x_{j}\rangle=\langle E_{M}(\Delta)y_{j},y_{i}\rangle$.
-For convenience, put $H_{0}=L_{N,Z}$, $H_{1}=H_{0}^{\perp}$, $N_{0}=N_{|H_{0}}$ and $N_{1}=N_{|H_{1}}$.
-By Theorem 2.10, there exists a conjugation $C_{0}$ on $H_{0}$ such that $C_{0}N_{0}C_{0}=-N_{0}^{*}$ and $C_{0}x_{i}=y_{i}$ for every $i\in I$.
-Since $N_{1}$ is skew-symmetric, there exists a conjugation $C_{1}$ on $H_{1}$ such that $C_{1}N_{1}C_{1}=-N_{1}^{*}$, and thus the conjugation $C=C_{0}\oplus C_{1}$ satisfies all required properties.
+$Vp(N_{k_{1}},N_{k_{1}}^{*},\dots,N_{k_{n}},N_{k_{n}}^{*})y_{i}$ $=p(-N_{k_{1}},-N_{k_{1}}^{*},\dots,-N_{k_{n}},-N_{k_{n}}^{*})^{*}Vy_{i}$
+$=p(-N_{k_{1}},-N_{k_{1}}^{*},\dots,-N_{k_{n}},-N_{k_{n}}^{*})^{*}x_{i}.$
+Put $\mathcal{L}_{0}=L_{\mathcal{N},X\cup Y}$.
+Then, the above equalities imply that $V\mathcal{L}_{0}=\mathcal{L}_{0}$.
+Since $V$ is anti-unitary, we get $V\mathcal{L}_{1}=V\mathcal{L}_{0}^{\perp}=\mathcal{L}_{0}^{\perp}=\mathcal{L}_{1}$.
+Let $V_{1}$ denote the restriction of $V$ to $\mathcal{L}_{1}$.
+Then, $V_{1}$ is anti-unitary on $\mathcal{L}_{1}$, and the relations $VN_{k}V^{-1}=-N_{k}^{*}$, for every $k\in K$, imply $V_{1}N_{k|\mathcal{L}_{1}}V_{1}^{-1}=-\left(N_{k|\mathcal{L}_{1}}\right)^{*}$.
+This proves the first part.
+Now, by Lemma 2.15, for all Borel subsets $\Delta\subseteq\mathbb{C}$ and $k\in K$, we have $VE_{N_{k}}(-\Delta)=E_{N_{k}}(\Delta)V$. The remainder of the proof is now similar to (2.8) and (2.9).
+(iii)$\implies$(i).
+For each $k\in K$, put $M_{k}=-N_{k}$.
+Let $K_{0},K_{1}\subseteq K$ be finite subsets, and let $\{\Delta_{k}\}_{k\in K_{0}}$ and $\{D_{k}\}_{k\in K_{1}}$ be Borel subsets of $\mathbb{C}$.
+Then, for all $i,j\in I$, we have
+$\left\langle\left(\prod_{k\in K_{0}}E_{N_{k}}(\Delta_{k})\right)x_{i},\left(\prod_{l\in K_{1}}E_{M_{l}}(D_{l})\right)x_{j}\right\rangle=\left\langle\left(\prod_{k\in K_{0}}E_{N_{k}}(\Delta_{k})\right)x_{i},\left(\prod_{l\in K_{1}}E_{-N_{l}}(D_{l})\right)x_{j}\right\rangle$
+$=\left\langle\left(\prod_{k\in K_{0}}E_{N_{k}}(\Delta_{k})\right)x_{i},\left(\prod_{l\in K_{1}}E_{N_{l}}(-D_{l})\right)x_{j}\right\rangle$
+$=\left\langle\left(\prod_{k\in K_{0}}E_{N_{k}}(-\Delta_{k})\right)x_{j},\left(\prod_{l\in K_{1}}E_{N_{l}}(D_{l})\right)x_{i}\right\rangle$
+$=\left\langle\left(\prod_{k\in K_{0}}E_{M_{k}}(\Delta_{k})\right)x_{j},\left(\prod_{l\in K_{1}}E_{N_{l}}(D_{l})\right)x_{i}\right\rangle.$
+Similarly,
+$\left\langle\left(\prod_{k\in K_{0}}E_{N_{k}}(\Delta_{k})\right)x_{i},\left(\prod_{l\in K_{1}}E_{M_{l}}(D_{l})\right)y_{j}\right\rangle=\left\langle\left(\prod_{k\in K_{0}}E_{M_{k}}(\Delta_{k})\right)x_{j},\left(\prod_{l\in K_{1}}E_{N_{l}}(D_{l})\right)y_{i}\right\rangle.$
+It follows by Theorem 2.1 that there exists a conjugation $C_{0}$ on $\mathcal{L}_{0}:=L_{\mathcal{N},X\cup Y}$ such that $C_{0}{N_{k}}_{|\mathcal{L}_{0}}C_{0}=-\left({N_{k}}_{|\mathcal{L}_{0}}\right)^{*}$ and $C_{0}x_{i}=y_{i}$ for all $k\in K$ and $i\in I$.
+Now, since by hypothesis $V_{1}{N_{k}}_{|\mathcal{L}_{1}}V_{1}^{-1}=-\left({N_{k}}_{|\mathcal{L}_{1}}\right)^{*}$ for every $k\in K$, Lemma 3.7 implies that there exists a conjugation $C_{1}$ on $\mathcal{L}_{1}$ so that $C_{1}{N_{k}}_{|\mathcal{L}_{1}}C_{1}=-\left({N_{k}}_{|\mathcal{L}_{1}}\right)^{*}$ for every $k\in K$.
+Therefore, $C=C_{0}\oplus C_{1}$ is a conjugation on $\mathcal{H}=\mathcal{L}_{0}\oplus\mathcal{L}_{1}$ that satisfies all required equalities.
 ∎
@@
 (ii)$\implies$(i).
-Let $N_{0}$ and $N_{1}$ be the normal operators in the proof of Theorem 3.2.
-In the light of the same theorem, we only need to show that $N_{1}$ is skew-symmetric.
+Let $\mathcal{H}_{0}=L_{\{N\},X\cup Y}$ and $\mathcal{H}_{1}=\mathcal{H}_{0}^{\perp}$, and denote by $N_{0}$ and $N_{1}$ the restrictions of $N$ to $\mathcal{H}_{0}$ and $\mathcal{H}_{1}$, respectively.
+In the light of Theorem 3.2 and Lemma 3.7, we only need to show that $N_{1}$ is skew-symmetric.
 This follows immediately from Theorem 3.10 and the facts that $m\left({N_{0}}_{|{\rm ker}(N_{0})^{\perp}}\right)<\infty$ and $N=N_{0}\oplus N_{1}$ is skew-symmetric.
 Let us now prove the second part of the theorem. Assume that $I$ is finite. Since
-$H_{0}=\bigvee\left(\sum_{i\in I}L_{N,\{x_{i}\}}+L_{N,\{y_{i}\}}\right)=\bigvee\left(\sum_{i\in I}L_{N_{0},\{x_{i}\}}+L_{N_{0},\{y_{i}\}}\right)$
-and the restrictions of $N_{0}$ to $L_{N_{0},\{x_{i}\}}$ and $L_{N_{0},\{y_{i}\}}$ are star-cyclic, Lemma 3.9 implies that $m({N_{0}}_{|{\rm ker}(N_{0})^{\perp}})<\infty$. This concludes the proof. ∎
-4. Applications to some equations in $L^{\infty}(\mu)$
-Throughout this section, let $\mu$ denote a finite measure on Borel subsets of $\mathbb{C}$ with compact support, and let $n$ be a positive integer.
-All functions are considered defined on $\mathbb{C}$.
-For a given function $f$, the notation $\widetilde{f}$ stands for the function defined by $\widetilde{f}(z)=\overline{f(-z)}$ for $z\in\mathbb{C}$.
-Let $[\phi_{i,j}]_{1\leq i,j\leq n}$ be a matrix of functions in $L^{\infty}(\mu)$.
-We denote $[\phi_{i,j}]_{1\leq i,j\leq n}\in\mathcal{U}(\mu,n)$ if the matrix $[\phi_{i,j}(z)]_{1\leq i,j\leq n}$ is unitary for $\mu$-almost every $z\in\mathbb{C}$.
-If, in addition, the functions $\phi_{i,j}$ satisfy the condition $\phi_{i,j}(z)=\phi_{j,i}(-z)$ for $\mu$-almost every $z\in\mathbb{C}$, then we write $[\phi_{i,j}]_{1\leq i,j\leq n}\in\mathcal{SU}(\mu,n)$.
-For a function $\phi\in L^{\infty}(\mu)$, $M_{\phi}$ denotes the multiplication operator by $\phi$ on $L^{2}(\mu)$; that is, $[M_{\phi}f](z)=\phi(z)f(z)$ for every $f\in L^{2}(\mu)$.
-Let $\{f_{i}\}_{1\leq i\leq n}$ and $\{g_{i}\}_{1\leq i\leq n}$ be functions in $L^{2}(\mu)$ so that $\sum_{k=1}^{n}|f_{k}|^{2}=\sum_{k=1}^{n}|g_{k}|^{2}$ $\mu$-almost everywhere ($\mu$-a.e.).
-It is clear that we can find functions $\phi_{i,j}$ so that, for $\mu$-almost every $z\in\mathbb{C}$, the matrix $[\phi_{i,j}(z)]_{1\leq i,j\leq n}$ is unitary and
-$\left[\phi_{i,j}(z)\right]_{1\leq i,j\leq n}\begin{bmatrix}f_{1}(z),f_{2}(z),\cdots,f_{n}(z)\end{bmatrix}^{\rm tr}=\begin{bmatrix}g_{1}(z),g_{2}(z),\cdots,g_{n}(z)\end{bmatrix}^{\rm tr}.$
-However, it is not clear whether the functions $\phi_{i,j}$ can be chosen to be Borel measurable. The next theorem provides a positive answer.
-Theorem 4.1.
-Let $\{f_{i}\}_{1\leq i\leq n}$ and $\{g_{i}\}_{1\leq i\leq n}$ be two families of functions in $L^{2}(\mu)$. Then, the following equation
-[…]
-has a solution $[\phi_{i,j}]_{1\leq i,j\leq n}\in\mathcal{U}(\mu,n)$ if and only if $\sum_{k=1}^{n}|f_{k}|^{2}=\sum_{k=1}^{n}|g_{k}|^{2}$ $\mu$-a.e.
-Lemma 4.2.
-Let $U\in\mathcal{B}\left(L^{2}(\mu)^{(n)}\right)$.
-Then, $U$ is unitary and commutes with $M_{\mu}^{(n)}$ if and only if $U=\left[M_{\phi_{i,j}}\right]_{1\leq i,j\leq n}$ for some $[\phi_{i,j}]_{1\leq i,j\leq n}\in\mathcal{U}(\mu,n)$.
-Proof.
-Suppose $U=\left[M_{\phi_{i,j}}\right]_{1\leq i,j\leq n}$ for some $[\phi_{i,j}]_{1\leq i,j\leq n}\subset L^{\infty}(\mu)$. Then,
-$UU^{*}$ $=\left[\sum_{k=1}^{n}M_{\phi_{i,k}}M_{\phi_{j,k}}^{*}\right]_{1\leq i,j\leq n}=\left[\sum_{k=1}^{n}M_{\phi_{i,k}}M_{\overline{\phi_{j,k}}}\right]_{1\leq i,j\leq n}$
-$=\left[\sum_{k=1}^{n}M_{\phi_{i,k}\overline{\phi_{j,k}}}\right]_{1\leq i,j\leq n}=\left[M_{\sum_{k=1}^{n}\phi_{i,k}\overline{\phi_{j,k}}}\right]_{1\leq i,j\leq n}.$
-Similarly, $U^{*}U=\left[M_{\sum_{k=1}^{n}\overline{\phi_{k,i}}\phi_{k,j}}\right]_{1\leq i,j\leq n}$. Hence, the operator $U$ is unitary if and only if, for all $1\leq i,j\leq n$,
-[…]
-where $\delta_{ij}$ denotes the Kronecker delta.
-That is, $[\phi_{i,j}]_{1\leq i,j\leq n}\in\mathcal{U}(\mu,n)$.
-The desired equivalence follows now directly from [8, Page 276, Proposition 6.1 (b)] and [8, Page 279, Corollary 6.9].
+$\mathcal{H}_{0}=\bigvee\left(\sum_{i\in I}L_{\{N\},\{x_{i}\}}+L_{\{N\},\{y_{i}\}}\right)=\bigvee\left(\sum_{i\in I}L_{\{N_{0}\},\{x_{i}\}}+L_{\{N_{0}\},\{y_{i}\}}\right)$
+and the restrictions of $N_{0}$ to $L_{\{N_{0}\},\{x_{i}\}}$ and $L_{\{N_{0}\},\{y_{i}\}}$ are star-cyclic, Lemma 3.11 implies that $m\left({N_{0}}_{|{\rm ker}(N_{0})^{\perp}}\right)<\infty$.
+This concludes the proof.
 ∎
-Proof of Theorem 4.1.
-In the light of the previous lemma, the equation (4.1) has a solution in $\mathcal{U}(\mu,n)$ if and only if there exists a unitary operator $U$ commuting with $M_{\mu}^{(n)}$ such that $UF=G$, where
-$F=\begin{bmatrix}f_{1},f_{2},\cdots,f_{n}\end{bmatrix}^{\rm tr}\quad\mbox{and}\quad G=\begin{bmatrix}g_{1},g_{2},\cdots,g_{n}\end{bmatrix}^{\rm tr}.$
-By Theorem 2.4, this is true if and only if $\left\|E_{M_{\mu}^{(n)}}(\Delta)F\right\|_{L^{2}(\mu)^{(n)}}=\left\|E_{M_{\mu}^{(n)}}(\Delta)G\right\|_{L^{2}(\mu)^{(n)}}$ for every Borel subset $\Delta\subseteq\mathbb{C}$.
-Note that
-$E_{M_{\mu}^{(n)}}(\Delta)=\bigoplus_{k=1}^{n}E_{M_{\mu}}(\Delta)=\bigoplus_{k=1}^{n}M_{\chi_{\Delta}}$
-(see, for instance, [9, Page 49, Proposition 10.4 (a)]). Consequently, (4.1) has a solution in $\mathcal{U}(\mu,n)$ if and only if, for every Borel subset $\Delta\subseteq\mathbb{C}$, we have
-$\int\sum_{k=1}^{n}|f_{k}|^{2}\chi_{\Delta}\,{\rm d}\mu$ $=\sum_{k=1}^{n}\int|f_{k}|^{2}\chi_{\Delta}\,{\rm d}\mu=\sum_{k=1}^{n}\left\|E_{M_{\mu}}(\Delta)f_{k}\right\|_{L^{2}(\mu)}^{2}=\left\|E_{M_{\mu}^{(n)}}(\Delta)F\right\|_{L^{2}(\mu)^{(n)}}^{2}$
-$=\left\|E_{M_{\mu}^{(n)}}(\Delta)G\right\|_{L^{2}(\mu)^{(n)}}^{2}=\int\sum_{k=1}^{n}|g_{k}|^{2}\chi_{\Delta}\,{\rm d}\mu.$
-That is, $\sum_{k=1}^{n}|f_{k}|^{2}=\sum_{k=1}^{n}|g_{k}|^{2}$ $\mu$-a.e. ∎
-The measure $\mu$ is called a symmetric measure if $\mu(\Delta)=\mu(-\Delta)$ for every Borel subset $\Delta\subseteq\mathbb{C}$.
-Theorem 4.3.
-Assume that $\mu$ is symmetric, and let $\{f_{i}\}_{1\leq i\leq n}$ and $\{g_{i}\}_{1\leq i\leq n}$ be two families of functions in $L^{2}(\mu)$. Then, the following equation
-[…]
-has a solution $[\phi_{i,j}]_{1\leq i,j\leq n}\in\mathcal{SU}(\mu,n)$ if and only if $\sum_{k=1}^{n}|f_{k}(z)|^{2}=\sum_{k=1}^{n}|g_{k}(z)|^{2}$ and $\sum_{k=1}^{n}f_{k}(z)g_{k}(-z)=\sum_{k=1}^{n}f_{k}(-z)g_{k}(z)$ for $\mu$-almost every $z\in\mathbb{C}$.
-Lemma 4.4.
-Let $U\in{\mathcal{B}}(H)$ be a unitary operator and $J$ a conjugation on $H$. Then, $U=JC$ for some conjugation $C$ on $H$ if and only if $JUJ=U^{*}$.
-Proof.
-In the light of the facts that $J^{2}=I$ and $\langle JUh,JUk\rangle=\langle Uk,Uh\rangle=\langle k,h\rangle$ for all $h,k\in H$, it is easy to see that
-[…]
-$\iff(JU)=(JU)^{-1}=U^{*}J^{-1}$
-$\iff JUJ=U^{*}.$
-Proof of Theorem 4.3.
-First, note that the map $J$ defined by $Jf(z)=\overline{f(-z)}$ is a conjugation on $L^{2}(\mu)$.
-In fact, $J$ is a well-defined conjugate-linear map with $J^{2}=I_{L^{2}(\mu)}$.
-Additionally, since $\mu$ is symmetric, we have
-$\langle Jf,Jg\rangle_{L^{2}(\mu)}=\int\overline{f(-z)}g(-z)\,{\rm d}\mu=\int\overline{f(z)}g(z)\,{\rm d}\mu=\langle g,f\rangle_{L^{2}(\mu)}$
-for all $f,g\in L^{2}(\mu)$.
-Also, for every $\phi\in L^{\infty}(\mu)$, we have
-$JM_{\phi}Jf(z)=JM_{\phi}\overline{f(-z)}=J\phi(z)\overline{f(-z)}=\overline{\phi(-z)}f(z)$
-for every $f\in L^{2}(\mu)$. Hence,
-(4.3) $JM_{\phi}J=M_{\widetilde{\phi}}.$
-In particular, $J^{(n)}M_{\mu}^{(n)}J^{(n)}=-M_{\mu}^{(n)*}$, and $M_{\mu}^{(n)}$ is skew-symmetric since $J^{(n)}$ is a conjugation on $L^{2}(\mu)^{(n)}$.
-Claim: An operator $U\in\mathcal{B}\left(L^{2}(\mu)^{(n)}\right)$ is unitary, commutes with $M_{\mu}^{(n)}$ and satisfies $J^{(n)}UJ^{(n)}=U^{*}$ if and only if $U=\left[M_{\phi_{i,j}}\right]_{1\leq i,j\leq n}$ for some $[\phi_{i,j}]_{1\leq i,j\leq n}\in\mathcal{SU}(\mu,n)$.
-By Lemma 4.2, it suffices to show that if $U=\left[M_{\phi_{i,j}}\right]_{1\leq i,j\leq n}$, then
-$J^{(n)}UJ^{(n)}=U^{*}\quad\Longleftrightarrow\quad\phi_{i,j}(z)=\phi_{j,i}(-z)\;\mu\mbox{-a.e. for all }1\leq i,j\leq n.$
-In fact, simple matrix calculations show that
-$J^{(n)}UJ^{(n)}=U^{*}$ $\quad\Longleftrightarrow\quad\left(JM_{\phi_{i,j}}J\right)_{1\leq i,j\leq n}=\left(M_{\phi_{j,i}}^{*}\right)_{1\leq i,j\leq n}$
-$\quad\Longleftrightarrow\quad\left(JM_{\phi_{i,j}}J\right)_{1\leq i,j\leq n}=\left(M_{\overline{\phi_{j,i}}}\right)_{1\leq i,j\leq n}$
-[…]
-$\quad\Longleftrightarrow\quad\phi_{i,j}(z)=\phi_{j,i}(-z)\;\;\mu\mbox{-a.e. for all }1\leq i,j\leq n.$
-This proves the claim.
-Now, one can see that the equation (4.2) has a solution in $\mathcal{SU}(\mu,n)$ if and only if, there exists a unitary operator $U$ on $L^{2}(\mu)^{(n)}$ commuting with $M_{\mu}^{(n)}$ such that $J^{(n)}UJ^{(n)}=U^{*}$ and
-$U\begin{bmatrix}f_{1},f_{2},\cdots,f_{n}\end{bmatrix}^{\rm tr}=\begin{bmatrix}g_{1},g_{2},\cdots,g_{n}\end{bmatrix}^{\rm tr}.$
-By Lemma 4.4, this true if and only if there exists a conjugation $C$ on $L^{2}(\mu)^{(n)}$ such that $J^{(n)}C$ commutes with $M_{\mu}^{(n)}$ and
-(4.4) $J^{(n)}C\begin{bmatrix}f_{1},f_{2},\cdots,f_{n}\end{bmatrix}^{\rm tr}=\begin{bmatrix}g_{1},g_{2},\cdots,g_{n}\end{bmatrix}^{\rm tr}.$
-Note that
-$J^{(n)}CM_{\mu}^{(n)}=M_{\mu}^{(n)}J^{(n)}C$ $\quad\Longleftrightarrow\quad CM_{\mu}^{(n)}C=J^{(n)}M_{\mu}^{(n)}J^{(n)}$
-[…]
-Hence, taking into account (4.4), the equation (4.2) has a solution in $\mathcal{SU}(\mu,n)$ if and only if there exists a conjugation $C$ on $L^{2}(\mu)^{(n)}$ such that $CM_{\mu}^{(n)}C=-M_{\mu}^{(n)*}$ and
-$C\begin{bmatrix}f_{1},f_{2},\cdots,f_{n}\end{bmatrix}^{\rm tr}=J^{(n)}\begin{bmatrix}g_{1},g_{2},\cdots,g_{n}\end{bmatrix}^{\rm tr}=\begin{bmatrix}\widetilde{g_{1}},\widetilde{g_{2}},\cdots,\widetilde{g_{n}}\end{bmatrix}^{\rm tr}.$
-For convenience, put $F=\begin{bmatrix}f_{1},f_{2},\cdots,f_{n}\end{bmatrix}^{\rm tr}$ and $\widetilde{G}=\begin{bmatrix}\widetilde{g_{1}},\widetilde{g_{2}},\cdots,\widetilde{g_{n}}\end{bmatrix}^{\rm tr}$.
-Then, by Theorem 3.2, the equation (4.2) has a solution in $\mathcal{SU}(\mu,n)$ if and only if
-$\left\langle E_{M_{\mu}^{(n)}}(\Delta)F,\widetilde{G}\right\rangle_{L^{2}(\mu)^{(n)}}=\left\langle E_{M_{\mu}^{(n)}}(-\Delta)F,\widetilde{G}\right\rangle_{L^{2}(\mu)^{(n)}}$
-and
-$\left\langle E_{M_{\mu}^{(n)}}(\Delta)F,F\right\rangle_{L^{2}(\mu)^{(n)}}=\left\langle E_{M_{\mu}^{(n)}}(-\Delta)\widetilde{G},\widetilde{G}\right\rangle_{L^{2}(\mu)^{(n)}}$
-for every Borel subset $\Delta\subseteq\mathbb{C}$. That is,
-$\left\langle\left(\bigoplus_{i=1}^{n}M_{\chi_{\Delta}}\right)F,\widetilde{G}\right\rangle_{L^{2}(\mu)^{(n)}}=\left\langle\left(\bigoplus_{i=1}^{n}M_{\chi_{-\Delta}}\right)F,\widetilde{G}\right\rangle_{L^{2}(\mu)^{(n)}}$
-and
-$\left\langle\left(\bigoplus_{i=1}^{n}M_{\chi_{\Delta}}\right)F,F\right\rangle_{L^{2}(\mu)^{(n)}}=\left\langle\left(\bigoplus_{i=1}^{n}M_{\chi_{-\Delta}}\right)\widetilde{G},\widetilde{G}\right\rangle_{L^{2}(\mu)^{(n)}}$
-for every Borel subset $\Delta\subseteq\mathbb{C}$; which means that
-$\int\sum_{i=1}^{n}f_{i}(z)g_{i}(-z)\chi_{\Delta}\,{\rm d}\mu=\int\sum_{i=1}^{n}f_{i}(z)g_{i}(-z)\chi_{-\Delta}\,{\rm d}\mu$
-and
-$\int\sum_{i=1}^{n}|f_{i}(z)|^{2}\chi_{\Delta}\,{\rm d}\mu=\int\sum_{i=1}^{n}|g_{i}(-z)|^{2}\chi_{-\Delta}\,{\rm d}\mu$
-for every Borel subset $\Delta\subseteq\mathbb{C}$. Therefore, by symmetry of $\mu$, for every Borel subset $\Delta\subseteq\mathbb{C}$,
-$\int\sum_{i=1}^{n}f_{i}(z)g_{i}(-z)\chi_{\Delta}\,{\rm d}\mu=\int\sum_{i=1}^{n}f_{i}(-z)g_{i}(z)\chi_{\Delta}\,{\rm d}\mu$
-and
-$\int\sum_{i=1}^{n}|f_{i}(z)|^{2}\chi_{\Delta}\,{\rm d}\mu=\int\sum_{i=1}^{n}|g_{i}(z)|^{2}\chi_{\Delta}\,{\rm d}\mu;$
-which lead to the desired conditions. ∎